TABLE OF CONTENTS



Cruise Location Map (Fig.1)................................................................................................ 3

Errata ................................................................................................................................... 4

NeMO '99 Scientific Party................................................................................................... 5

1 CRUISE OVERVIEW........................................................................................................6
Dives on Axial (Fig. 2)......................................................................................................... 6b

SE Caldera SRZ, Northern Vents and Dive Tracks (Fig. 3)................................................. 7

SE Caldera SRZ, Southern Vents and Dive Tracks (Fig. 4)................................................. 8

ASHES Vent Field, Vent Names and Locations (Fig. 5)...................................................... 9

2 DISCIPLINE SUMMARIES..............................................................................................10

2.1 VOLCANOLOGY...............................................................................................................10

2.2 CHEMISTRY...................................................................................................................... 11

2.2a Hydrothermal Fluid Sampling.............................................................................................. 11

2.2b OsmoSampler and OsmoAnalyzer Operations..................................................................... 12

2.2c Gas Sampling........................................................................................................................ 13

2.2c Studies of Dissolved Gases from Hydrothermal Vent Systems............................................ 14

2.3 MICROBIOLOGY..............................................................................................................14

2.3a Microbiological Sampling for Molecular Microbial Ecology Analysis............................... 14

2.3b Hydrothermal Fluid Microbiology....................................................................................... 16

2.3c Microbial Food Webs........................................................................................................... 17

2.4 MACROBIOLOGY............................................................................................................ 18

2.4a Biology of Low Temperature Sites....................................................................................... 18

2.5 HYDROTHERMAL DEPOSITS.......................................................................................20

2.6 ROCK SAMPLING AND PETROLOGIC STUDIES ....................................................20

3.0 Non-ROPOS OPERATIONS............................................................................................. 22

3.1 Mooring Deployments /CTD's /XRF Analysis.....................................................................22

3.2 NeMO'99 Website and Public Outreach.............................................................................. 22

4 NAVIGATION................................................................................................................... 23

4.1 Navigation Overview........................................................................................................... 23

4.2 Final Calibrated Transponder Positions............................................................................... 24

4.3 Vents and Markers Location Table...................................................................................... 24

5 NeMO OPERATIONS....................................................................................................... 27

5.1 1998 Dive Dates and Locations - ROPOS Dives R460 -R480......................................... 27

5.2 1999 Dive Dates, Locations and Tasks - ROPOS Dives R482 - R503............................ 28

5.3 Experiments Deployed and Recovered - 1998 and 1999................................................. 30

5.4 1999 ROPOS Samples - Dives R483 - R503..................................................................... 32

5.5 Dive Map Nomenclature..................................................................................................... 50

5.6 1999 ROPOS Dive Logs - Dives R483 - R503 ................................................................. 51

(dive maps inserted in dive log, where appropriate)

R483 Dive Log..................................................................................................................... 51

R484 Dive Log..................................................................................................................... 54

R485 Dive Log..................................................................................................................... 58

R486 Dive Log..................................................................................................................... 60

R487Dive Log..................................................................................................................... 61

R488 Dive Log..................................................................................................................... 63

R489 Dive Log ................................................................................................................. 78

R490 Dive Log ................................................................................................................. 79

R491 Dive Log..................................................................................................................... 80

R492 Dive Log..................................................................................................................... 96

R493 Dive Log..................................................................................................................... 108

R494 Dive Log..................................................................................................................... 112

R495 Dive Log..................................................................................................................... 130

R496 Dive Log..................................................................................................................... 149

R497 Dive Log..................................................................................................................... 155

R498 Dive Log..................................................................................................................... 165

R499 Dive Log..................................................................................................................... 168

R500 Dive Log..................................................................................................................... 171

R501 Dive Log..................................................................................................................... 179

R502 Dive Log..................................................................................................................... 192

R503 Dive Log..................................................................................................................... 203



Cover: Image of high resolution pencil beam bathymetry (Imagenex). Bathymetry grid cell size is 2 meters. Depths range from 1500 - 1575 meters. Orange/red colors shallowest, blues/purples deepest. Imagenex data were collected by the NOAA Vents Program using ROPOS during NeMO'98 and NeMO'99. Data were supplemented with additional bathymetry collected on the Cleft'99 cruise in September 1999. Bathymetry processing and cartography by William W. Chadwick Jr., with assistance from Robert W. Embley and Susan G. Merle.



********************************************************************************************



Errata



Dive Logs:

ROPOS dive logs were saved in .html or Excel .csv (comma delimited) files. The .html files were not an option for editing, as they can only be saved as text files, and lose all their rows and columns. The .csv files were a little easier to work with so they were the files edited for the dive logs. (As you may recall, we could not edit the logs at sea.) There was a problem with the .csv dive log files as well. Any double letters or double numbers were absent. I tried to find those errors and correct them, but be aware, especially regarding numbers, as those errors were particularly hard to spot.



Navigation in general:

1999 navigation not as robust as 1998. 1999 nav fixes were often to the west or southwest of '98 fixes. Unfortunately this is not always the case. The farther south we went, the worse the acoustic nav. That is because the southern transponder net could not be calibrated with the ASHES net. (See section 4 for more details.) Be aware of that and use the dive plots as a guide.



Dive Plots:

R492 Missing acoustic nav at end of dive, after JD183 2130. We weren't getting any fixes, and there were no good fixes recorded in the nav logbook to supplement the acoustic fixes. The dive continued for a couple more hours, ending about 250 meters southeast of Bag City. Will fill in last two hours of nav with adjusted ship nav, eventually.



R494 Video went out at JD185 1810, therefore there was no bottom image from that time until ROPOS came up. 1810 official end of dive.



R495 There are gaps in the nav from JD186 1454 - 1640 and 2236 - JD187-0000. We were getting out of the range of the ASHES net transponders. The dive plots just show a straight line connecting the times. Acoustic nav ends on JD187 0115, but the dive continued until 0158. All the nav gaps will be filled with adjusted ship nav, eventually.



R501 No acoustic nav from JD191 2251 - JD192 0521. ROPOS on the deck or in the water column from 0230 - 0530. Need to fill in JD191 2251 - JD192 0230 with adjusted ship nav, eventually. The dive plot shows the ROPOS track from JD192 1411 - 1520, but ROPOS was transiting off the bottom at the time and there was no video.



R494 and R497 Plots state that the nav fixes for the ship are at the stern. That is not accurate. We presume the GPS fixes were referenced to the GPS antennae, about 37 meters fore of the stern. To get actual position at stern one would have to apply a 37 meter lay-back to the ship navigation. (To confuse the matter more: Navigation fixes transcribed in the nav logbook are positions at stern, because we put the pointer in that position on the nav screen and wrote down that fix. More information than you want to know?)

Scientific Participants, Title and Affiliation:



Name Title Affiliation

Robert Embley Chief Scientist/Geologist PMEL

Dave Butterfield Chemist JISAO/PMEL

Bill Chadwick Geologist CIMRS/OSU/PMEL

Kim Juniper Biologist UQ

Michael Perfit Geologist UF

Craig Moyer Microbiologist WWU

Steve Scott Geologist UT

Thomas Chapin Chemist MBARI

Leigh Evans Chemist CIMRS/OSU/PMEL

Jim Gendron Chemist PMEL

Susan Merle Navigation/Geologist CIMRS/OSU/PMEL

Julie Huber Microbiologist/Grad Student UW

Marlene le Bel Biologist/Student UQ

Catherine Charpentier Biologist/Grad Student UQ

Christian Levesque Biologist/Grad Student UQ

Sheryl Roadruck Microbiologist UW

Andy Graham Chemist UW

Jean Marcus Navigation/Biologist/Grad Student UV

Maia Tsurumi Biologist/Grad Student UV

Paul Johnson Navigation/Geologist U Hawaii

Karen Lynch Geologist WWU

John Chadwick Geologist/Grad Student UF

Kevin Roe Chemist JISAO/PMEL

Nicole Nasby NeMO Web/Grad Student OSU

Naaznin Pastakia Geologist/Grad Student UT

Greg Pillette Teacher Beaverton High School

Keith Shepherd ROPOS Team Leader CSSF

Bob Holland Engineer CSSF

Keith Tamburri Engineer CSSF

Kim Wallace Engineer CSSF

Mike Dempsey Engineer CSSF

Craig Elder Engineer CSSF

Ian Murdock Engineer CSSF



Participating Organizations:

PMEL Pacific Marine Environmental Laboratory

UW University of Washington

OSU Oregon State University

CIMRS Cooperative Institute for Marine Resources Studies (NOAA/OSU)

JISAO Joint Institute for the Study of Atmosphere and Oceans (NOAA/UW)

WWU Western Washington State University

CSSF Canadian Scientific Submersible Facility

UF University of Florida

UQ University of Quebec

UT University of Toronto

UV University of Victoria

MBARI Monterey Bay Aquarium Research Institute

1 OVERVIEW OF NeMO'99 - Bob Embley



During the NeMO99 cruise, ROPOS recovered more than 260 geologic, biologic and chemical samples, retrieved 29 experiments and and instruments, and deployed 25 experiments. ROPOS made 18 successful dives totaling 204 hours of bottom time in 21 days on site. Despite major problems with the hydraulic systems on the ROPOS, which caused us considerable downtime, the hardworking ROPOS team and the efforts of the science party were able to recover to make this year's cruise rival the success of NeMO98.



The chemical and biological effects of the 1998 diking event are still apparent on the eastern side of the caldera where the lava flow was erupted. Vigorous hydrothermal venting continues in many places (e.g. Marker 33, Cloud, and Magnesia), whereas there has been reduction or cessation in others. Although the amount of visually apparent microbial activity (such as the white floc in the water column) decreased, "snowblower" activity continued at some vents (e.g. Magnesia Vent) and hyperthermophilic microbes were again found in abundance at most of the diffuse (low temperature) sites, (whose temperature is well below their optimal growth rate). These observations show that there is still enhanced subsurface microbial growth more than a year and a half after the eruption. Of particular interest was the discovery of a hyperthemophile growing optimally at 102C.



Colonization of the new vents continued. Tubeworms had arrived at Marker 33 and were present in greater densities at sites that only contained small patches in 1998 (e.g., Nascent Vent). Analyses of the water samples, biological grabs, and video observation should tell us much about the evolution of the chemical and biological systems on the summit of Axial.

Mapping of the new lava flow and associated hydrothermal system continued in 1999. We located the (probable) southern end of the lava flow about 6 kilometers south of the caldera along a spectacular fissure system that was followed for hundreds of meters. The eruption was entirely contained in the 1 to 2 meter wide fissure until it overflowed its fissure and erupted in the large mound of lava that was found in 1998 at 4552'N. Farther north towards the caldera, the trace of the 1998 eruption became more confused as the slightly older eruptions mixed with the 1998 lava. Several new vents (Bag City, Crevice, Coquille) were found south of Marker 113 (the southernmost point investigated in 1998). The vent fluid sampler recovered 113 water and filter samples from all the (still active) major vents.

The ASHES vent field was only visited on two dives in 1999 because of the early downtime of ROPOS, and part of one dive was again spent at the CASM vent field in the northern part of the caldera.



Active hydrothermal venting was only found on the summit of Axial, probably where the underlying zone of molten lava permanently resides. The summit area contains hydrothermal systems formed during several different volcanic or tectonic (earthquake generated) events, providing a natural laboratory for studying colonization and evolution of these unique chemosynthetic ecosystems.









2 DISCIPLINE SUMMARIES



2.1 VOLCANOLOGY - Robert W. Embley and William W. Chadwick

Volcanology studies during NeMO'99 focused on, 1) continuing our instrumental monitoring efforts, 2) expanding our Imagenex sonar bathymetry of the 1998 eruption site, and 3) extending our mapping of the 1998 lava flow south to the area of the southern SeaBeam depth anomaly at 45 52' (confirmed as a new lava flow during dive R465 last year).



The extensometers that were recovered from Axial's north rift zone during NeMO'98 yielded an interesting data set. The instruments had been in place during the 1998 eruption and recorded a 4 cm contraction at the time of the eruption. This represents deflation of the entire summit area of Axial as magma intruded into the south rift zone, and is consistent with the rumbleometer instruments that recorded dramatic subsidence in Axial caldera. The extensometer and rumbleometer data can be modeled together to estimate the depth of the magma reservoir beneath the caldera and the volume of magma removed during the eruption/intrusion. These results were published in the December 1999 issue of Geophysical Research Letters. After the data was downloaded from the extensometers in 1998, they were re-deployed in the same area, and we recovered these again in 1999. Unfortunately, all 4 instruments failed during the 1999 deployment due to connector leaks and failure of components on the circuit boards. The status of these instruments for future deployments is currently being assessed.



The high-resolution bathymetry that can be collected with the Imagenex sonar was one of the highlights of NeMO'98. We were able to double the area of Imagenex coverage in 1999, mainly to the west of the area mapped in 1998. This helped define the western contact of the 1998 lava flow, and also showed many drain-out channels in the older lava to the west, suggesting that the previous eruptions from this area were more voluminous that the 1998 eruption. The Imagenex map fills a critical observational gap between multibeam bathymetry and bottom observations, and helps us understand what we see from ROPOS. It helps us map out the 1998 lava flow in unprecedented detail, including collapse areas and flow morphologies, puts the locations of vents sites and samples in a meaningful spatial context, and helps us interpret the sequence of events during the eruption. This is the first high-resolution map of a mid-ocean ridge eruption site and we will continue to expand the Imagenex coverage in future years.



During NeMO'98, our geologic mapping efforts were concentrated in the 3 km north of Marker 113. However, we know that the 1998 eruption extended at least 7 km south of Marker 113, so during NeMO'99, we extended this mapping to the south. Observations during dive R495 showed that even the large pillow mound that erupted at 45 52' was fed from a narrow fissure. Within a few hundred meters south of the pillow mound, 1998 lava can be seen filling, but not overflowing, a fissure that is 1-1.5 m wide. This extraordinary observation is essentially the surface outcrop of the 1998 dike that fed the eruption at this location, and the width of the fissure we saw is the minimum amount of seafloor spreading that occurred during this event. Between the pillow mound at 45 52' and Marker 113, the story is more complicated because there are other very young lava flows that are difficult to distinguish from the 1998 lava, except that locally they host clearly pre-1998 vent communities (mainly large, dead tubeworms). We were able to map out the distribution of lavas of at least 4 different ages during dives R492, R493, R494, R495, and R501. However, these dives only covered limited parts of this large area, and additional mapping will be necessary in future years to complete this mapping effort. In collaboration with M. Perfit and his students at the University of Florida, a careful collection of basalt samples was made during these mapping dives to determine if there are distinctive chemical differences between the 1998 lava flow and the surrounding older lavas. These results might help in the mapping effort and have implications for the evolution Axial's magmatic system.



The geologic results from this project also aid in putting into context the remarkable changes observed in the chemical and biologic systems at Axial Volcano since the 1998 eruption. The mapping provides the geologic context for the along-strike variations in the chemical and biologic characteristics of the hydrothermal system. For example, we were able to compare the general pattern of hydrothermal flow from the new eruption with that of the pre-eruptive system (mapped by previous camera tows and dives), suggesting that the conduit system was not dramatically changed by the dike injection and eruption. The north-south pattern appears to be similar, but there was apparently a small East-West displacement between the pre- and post-eruptive hydrothermal systems. The center of maximum hydrothermal flow appears to have remained in roughly the same place. This is also apparently where the maximum diversity in the microbial population is. The "snow blower" vents occur primarily at the northern portion of the system, where the lowest flow rates are located.



2.2 CHEMISTRY



2.2a Hydrothermal Fluid Sampling - David Butterfield

The main goals for fluid sampling for this cruise were 1) re-sample vent fluids from sites sampled last year and expand the sampling to cover a broader range of the eruption area in the SE caldera and rift zone, 2) re-sample diffuse, warm vents and hot vents in the ASHES field, and 3) sample vent fluids from CASM. The vent fluid studies are intended to use chemistry to understand the connections between volcanic activity, hydrothermal heat loss, and microbial activity in hydrothermal systems. Fluid chemistry is also an important measure of habitat conditions for studies of vent fauna colonization and species distribution.



Very few studies to date have focused intensively on the variation in diffuse fluid chemistry within a region. This work may yield new insights into the variability of sub-seafloor processes in a hydrothermal system.



Hot Fluid Sampler

We used the Hot Fluid Sampler (HFS) for the second year during this cruise, after several improvements were made over the proto-type model of last year. The sampler has a lower profile in front of the ROV so that visibility and ability to sample are not severely impaired. The plumbing layout was improved to allow more filters on the instrument. This year we were configured to take 6 piston samples (3 modified with high-vacuum valves for gas sampling), 8 collapsible bag samples with optional in-line filters, and 10 additional filters without water collection. One of the bag samplers was used as a flushing line for two titanium gas-tight samplers so that fluids of known temperature could be sampled by different types of samplers for comparison. On one dive, we can collect 10 discrete water samples for fluid chemistry, 5 water samples for gas analysis, and up to 17 filters for a variety of analytical purposes.



Several types of filters were used. Membrane filters of .45 or .2 micron pore size were used to collect particulate material for: xrf chemical analysis and SEM, particulate elemental sulfur, particulate organic carbon, lipids, and fluorescent in-situ hybridization (FISH) analysis. In addition, "Sterivex" high through-put cartridge filters were used to collect particles for DNA extraction. All of the analytical work done on the filters is designed to help us understand the chemical and microbiological variation among different types of vents. The in-situ filtration capability of the fluid sampler is an important advance over all previous vent fluid sampling instrumentation in that it allows us to concentrate particulate material from large volumes of water for multiple analyses at a given vent site, and it also allows us to separate the dissolved components from the particulate components at the time of venting, thereby eliminating uncertainty about post-sampling particle formation within the water sample containers. In many instances, we collect both filtered and unfiltered water from the same site for comparison. Overall, we have an excellent set of samples for fluid chemistry, gas chemistry, particle chemistry, and microbiology.



On-board fluid analysis:

Kevin Roe was the principal fluid analyst on this cruise, and carried out analysis for pH, alkalinity, hydrogen sulfide, dissolved silica, and ammonia. Most of our chemical analysis occurs on shore. We measured refractive index of the hotter vent fluids to estimate their salinity or chlorinity. Dave Butterfield analyzed particulate elemental sulfur by colorimetry on a subset of samples.



We were fortunate to have Andy Graham on board. Andy received cuts of most of the fluid samples for methane and hydrogen analysis by gas chromatography. In many cases, vent fluid samples arrived on deck containing a gas phase and a liquid phase. When this occurred we measured the volume of the liquid and gas portions and Andy analyzed both for gas content. In this way, the fluid sampler appears to be superior to the traditional titanium major samplers because it has much better gas retention capability, and the samplers are transparent, allowing us to separate the gas portion from the liquid. We cannot rule out some gas loss through the check valves, but comparison of the gas results from different samplers will help to evaluate this.



Initial results:

HFS was deployed on 2 short aborted dives and 3 complete dives in the 1998 eruption area, plus one complete dive at ASHES. HFS proved to be quite efficient. For example, on the first HFS dive, we had only 30 minutes of bottom time at marker 33 and took 5 samples (2 water chemistry, one gas chemistry, and two filters). For the entire cruise, we collected 43 vent fluid samples for water chemistry with HFS, 13 gas piston samples, and over 50 filters. In addition, titanium gas-tight samples were collected on many dives (see Leigh Evans report) and these will be analyzed for major element chemistry.



We have not yet processed the raw data from the last two dives, but the initial results appear similar to last year. Relative to the CoAxial eruption area, where there were large changes in composition and a decay in heat output one year after the eruption, we saw little change on the seafloor in the eruption area. Diffuse venting continued in many areas. The maximum measured temperature at marker 33 is higher than last year (78 versus 55 degrees). Looking at the ratio of hydrogen sulfide to silica (the ratio is only slightly sensitive to the amount of entrained seawater during sampling) it appears that the samples from throughout the eruption area this year are just slightly lower in sulfide relative to silica. Methane appears to be slightly higher this year than last year. We cannot say anything about the changes in salinity until we do high-precision analysis on shore.



2.2b OsmoSampler and OsmoAnalyzer Operations - Thomas Chapin

Changes in the chemical composition of hydrothermal effluent after a tectonic-volcanic event have been documented (e.g., Baker et al., 1987, 1998; Butterfield and Massoth, 1994; Von Damm et al, 1995; Massoth et al., 1995; Massoth et al., in press; Wheat et al., to be submitted) and a conceptual model has been developed that theorizes the chemical evolution of venting fluids (Butterfield et al., 1997). However, the timing of these changes is uncertain. To date observations of temporal variability in the chemical composition of hydrothermal fluids has relied on repeated submersible operations and the collection of discrete samples. While this technique provides some temporal constraints, a continuous water sampler or analyzer allows one to collect more samples with limited need for costly submersible operations. Our goal for this cruise was to deploy two short-term (two weeks) and two long-term (one year) continuous sampling systems to provide temporal constraints for observing hourly to daily and weekly to monthly chemical cycles in the hydrothermal effluent. Data from these samplers and their comparison to samples collected using traditional discrete sampling techniques will allow us to determine the temporal scale of chemical change in the hydrothermal effluent as the hydrothermal system evolves and may provide constraints for understanding the physical and chemical conditions at depth and the path for fluid circulation.

Two sampling systems were deployed, OsmoSamplers and OsmoAnalyzers. OsmoSamplers are continuous water samplers that use the osmotic pressure that is created across a semi-permeable membrane by solutions of differing salinity (Theeuwes and Yum, 1976; Jannasch et al., submitted). This pressure drives water across the membrane at a speed that is dependent on the surface area of the membrane, type of membrane, salt gradient, and temperature. An excess of salt is maintained on one side of the membrane, thus only temperature affects the flow of water in the sampler. Pumps in an OsmoSampler are used to continuously draw sample through a small bore (0.8 mm id) tubing that is attached to a 40-cm-long T-handle. An additional pump was used to add acid to the sample stream in most of the OsmoSamplers. A 1.5-m-long section of tubing separates the sample intake from the pump to allow the pump to be placed in an area void of hydrothermal influence and thus minimizes temperature (pump rate) fluctuations. A temperature recorder with a resolution of 0.0018C is attached to the T-handle to monitor the same water that is being collected by the OsmoSampler. Chemical data are obtained by retrieving the sampler, cutting the sample tubing into sections, extracting the seawater, and analyzing the seawater for chemical species of interest. Time-stamps for individual samples are determined assuming a uniform temperature at the pump that translates into a uniform rate of pumping.



OsmoAnalyzers, in contrast to OsmoSamplers, use osmotic pumps to deliver reagents into a sample stream for in situ analysis (Jannasch et al 1994). An iron OsmoAnalyzer was deployed at Marker 33 to continuously measure Fe at 15-minute intervals over the next 6-9 months.



Two long-term acid addition OsmoSamplers, deployed from the NeMO 1998 September cruise were recovered on this cruise. These samplers, one at Marker 33 and one at Milky vent, continuously collected sample for 9 months providing 163 0.5-mL samples. Analysis for major elements and trace metals will be performed later in our laboratory. Milky vent started at 9.5 C but dropped to 3.6 C by the time the OsmoSampler was recovered. There was no visible floc venting from Milky and the low temperatures at the end of the deployment indicate that the diffuse vent had died out. Marker 33, on the other hand, continued to vent hydrothermal fluids up to 80 C and was quite vigorous. The NeMO-98 long term acid addition OsmoSamplers deployments appear to have been a success and will provide one of the first long-term continuous records of the chemical signature of hydrothermal fluids.



During NeMO-99, four long-term OsmoSamplers were deployed. Three OsmoSamplers were deployed at the Marker 33, in the hottest section with temperatures up to 80 C. The OsmoSamplers consisted of one regular acid addition, one bio-OsmoSampler which has a sodium azide biocide to prevent bacterial growth, and a long term Cu-OsmoSampler which will collect samples for gas analysis. Another long-term acid addition OsmoSampler was deployed in 70 C water at Magnesia vent, just north of Milky vent. Tremendous clouds of white flocculent material were coming out of Magnesia and it looked like a snowstorm.



A 2-week record high temperature record from the Hell vent was recovered. Unfortunately the deployment of a long-term acid addition OsmoSampler and temperature probe was not successful.



References:

Baker, E. T., G. J. Massoth, and R. A. Feely. 1987. Cataclysmic hydrothermal venting on the Juan de Fuca Ridge. Nature, 329, 149-151.

Baker, E. T., G. J. Massoth, R. A. Feely, G. A. Cannon, and R. E. Thomson. 1998. The rise and fall of the CoAxial hydrothermal site, 1993-1996. J. Geophys. Res., 103, 9791-9806.

Butterfield, D.A., and G. J. Massoth. 1994. Geochemistry of north Cleft segment vent fluids: Temporal changes in chlorinity and their possible relation to recent volcanism. J. Geophys. Res., 99, 4951-4968.

Butterfield, D. A., I. R. Jonasson, G. J. Massoth, R. A. Feely, K. K. Roe, R. E. Embley, J. F. Holden, R. E. McDuff, M. D. Lilley, and J. R. Delaney. 1997. Seafloor eruptions and evolution of hydrothermal fluid chemistry. Phil. Trans. R. Soc. Lond. A, 355, 369-386.

Jannasch, H. W., K. S. Johnson and C. M. Sakamoto. 1994. Submersible, osmotically pumped analyzers for continuous determination of nitrate in situ. Anal. Chem. 66, 3352-3361.

Jannasch, H. W., C. G. Wheat, M. Kastner, and D. Stakes. 1998. Long-term in situ osmotically pumped water samplers. Deep Sea Res., submitted.

Massoth, G. J., E. T. Baker, R. A. Feely, D. A. Butterfield, R. E. Embley, J. E. Lupton, R. E. Thomson, and G. A. Cannon. 1995. Observations of manganese and iron at the CoAxial seafloor eruption site, Juan de Fuca Ridge. Geophys. Res. Lett., 22, 151-154.

Massoth, G. J., E. T. Baker, R. A. Feely, J. E. Lupton, R. W. Collier, J. F. Gendron, K. K. Roe, S. M. Maenner, and J. A. Resing. 1998. Manganese and iron in hydrothermal plumes resulting from the 1996 Gorda Ridge Event. Deep Sea Res., in press.

Theeuwes, F., and S. I. Yum. 1976. Principles of the design and operation of generic osmotic pumps for the delivery of semisolid or liquid drug formulations. Ann. Biomed. Eng., 4, 343-353.

Von Damm, K. L., S. E. Oosting, R. Kozlowski, L. G. Buttermore, D. C. Colodner, H. N. Edmonds, J. M. Edmond, and J. M. Grebmeier. 1995. Evolution of East Pacific Rise hydrothermal fluids following an oceanic eruption. Nature, 375, 47-50.

Wheat, C. G., H. W. Jannasch, F. J. Sansone, J. N. Plant, and C. L. Moyer. 1998. Hydrothermal Fluids From Loihi Seamount After the 1996 Event: A Year of Change Monitored With a Continuous Water Sampler. Earth Planet. Sci. Lett., to be submitted.



2.2c Gas Sampling - Lee Evans

The primary goal of gas sampling during the NeMO '99 expedition was direct sampling of vent fluids by way of Titanium Gastight bottles and modified pistons from the PMEL Hot Fluid Sampler (HFS). Approximately 30 useful samples were gathered and their available gas contents extracted and sealed in glass ampoules for chemical analysis. Analyses include helium isotopes, hydrogen and methane.



As with 1998's samples, the geographic coverage of sampling included the east side of the caldera along the region of the 1998 lava flow, Ashes vent field on the west side and Casm vent field to the north. Time series measurements will be possible at about 5 vent sites. The coverage of diffuse vent samples was extended southward on the east side in the direction of the vestige of the eruptive fissure.



This year's method modifications present a significant improvement as compared with those used to gather samples in 1998. At least some of 1998's collection were a bit more dilute than what is desirable. Both the plumbing scheme for titanium gastight bottles and the sample integrity of the gas piston samplers (HFS) were improved.



2.2d Studies of Dissolved Gases from Hydrothermal Vent Systems - Andy Graham

The main focus of our lab is the study of dissolved gases from hydrothermal vent systems. For this cruise I brought a gas chromatograph on board and analyzed fluid samples from the hot fluid sampler, suction sampler and a niskin bottle mounted on ROPOS. The main gases that I analyzed were dissolved hydrogen and dissolved methane. These gases are important in the hydrothermal vent community because certain microbes can oxidize these gases and use them as an energy source. Over 50 samples were analyzed ranging from 300 oC Inferno vent to 4 oC Magnesia vent. From an initial glance the data vent such as Marker 33 and Virgin Mound still appear to contain high concentrations of both hydrogen and methane. Further analysis and a comparison to last year's data will occur.



2.3 MICROBIOLOGY



2.3a Microbiological Sampling for Molecular Microbial Ecology Analysis

Western Washington University, Biology Department: Craig L. Moyer & Karen Lynch.



Introduction

One of the greatest challenges in microbial ecology is the accurate identification and description of microbial populations within their respective communities. This information is central to determining the extent of global microbial diversity, which remains the least understood of all the biological size classes. To address this challenge, molecular biological techniques using small-subunit ribosomal RNA (SSU rRNA) gene sequences have been applied to describe the structure and diversity of different microbial communities. The current endeavor is to examine specific habitats with known biogeochemical characteristics (e.g., S, Fe, Mn) to learn more about the dominant microorganisms residing therein. The focus of this study at Axial Volcano is to estimate the microbial community structure and diversity to assess the degree of commonality and uniqueness among local hydrothermal vent habitats, (i.e., vent-associated sediments, free-living microbial mats, microbes associated with subsurface floc-ejecta), and to also compare these results with distal hydrothermal vent habitats. This study will also allow for the enhanced development of a comprehensive global perspective regarding the diversity of deep-sea microbial communities.



Selective enrichment culture has severe limitations as an approach to the cultivation of naturally-occurring microorganisms. The majority (typically >90-99%) of microbes in nature have not yet been cultivated using traditional techniques. Consequently, it is very unlikely that collections of microbial isolates are representative of in situ diversity and community structure. Furthermore, because relatively nutrient-rich media are generally used for isolations, "weedy" or opportunistic microorganisms may be selected rather than those dominant in the natural community. The approach, herein, is to ascertain a microbial community's primary members through molecular (i.e., cell component) means and then to attempt to further characterize their respective phylogeny or natural history. Obtaining a better representation of microbial community structure and diversity is crucial to aspects of microbial ecology where Bacteria and Archaea interact with one another and with their environment, e.g., global biogeochemical cycling of matter, succession and disturbance responses, predator-prey relationships, and trophic-level interactions. These lessons can then be used to focus enrichment culture techniques towards ecologically significant taxa. This approach has been successfully used to isolate the dominant iron-oxidizer bacterial taxon found within the microbial community at hydrothermal systems located at Loihi Seamount, North Gorda Ridge, and other habitats (Emerson and Moyer, 1997; unpublished results).



Cell component analyses provide a culture-independent means of investigating microorganisms as they occur at hydrothermal vent systems (Moyer et al., 1994;1995; 1998). While several types of cell components have been analyzed, the SSU rRNA molecule offers an amount and type of information that makes it one of the best culture-independent descriptors or biomarkers of microorganisms. In recent years a detailed theory of evolutionary relationships among the domains Bacteria, Archaea and Eucarya has emerged from comparisons of SSU rRNA "signature" sequences. For example, each SSU rRNA gene contains highly conserved regions found among all living organisms as well as diagnostic variable regions unique to particular organisms or closely related groups. Additionally, each SSU rRNA gene contains about 1,500 nucleotides of sequence information that can be obtained and utilized to differentiate among closely-related and distantly-related groups of microorganisms. This type of molecular approach allows the autecology of microorganisms to be studied whether or not they can be been cultivated (Moyer et al., 1996). In addition, the phylogenetically described taxa or "phylotypes" can be placed in a synecology context through the examination of SSU rRNA clone libraries generated from a microbial community and habitat diversity can be analyzed through rarefaction (Moyer et al., 1998). These features make SSU rRNAs particularly useful for studies of molecular microbial ecology, where a broad and unknown range diversity of microorganisms is likely to exist. Currently, over 10,000 SSU rRNA sequences from both cultured isolates and environmental phylotypes have been made available for study through the Ribosomal Database Project at NSF's Center for Microbial Ecology at Michigan State University.



Experimental Design and Methods

Shipboard Processing and Storage of Samples

A dual approach was used for microbial sampling. First, a "slurp" gun suction device was be used in combination with a rotating rosette of sample bottles to "vacuum" and capture free-living microbial mats from the surface of various hydrothermal vent habitats. Slurp gun samples were successfully obtained from the East-Side of Axial at (1) Marker #33 Vent, (2) Markers N6 & N4 aka Cloud Vent, and (3) Magnesia aka Whiteout Vent. We also began to investigate the phenomena of the "bag creature" this year with slurp samples collected on the East side of Axial at both Axial Gardens (Marker #113) and from a new site entitled Joystick Vent (Marker #42). This is a characteristic jelly-like residue, which looks to be composed of complex polysaccharide globules that form in and around low temperature diffuse flowing vents in conjunction with microbial mats. No suction samples used for microbiology were obtained from the vicinity of the ASHES vent area this year.



Second, the deployment and recovery of microbial traps using glass wool as a substrate for microbial growth. Microbial traps were constructed using a cluster of three 3" sections of 4"o.d. plexiglass tubing, surrounded top and bottom by a 202 m nylon mesh (Nytex) to exclude macrofauna and meiofauna grazing. These were placed directly into diffuse vents and were used to collect colonizing microorganisms in an effort to examine community succession. These were deployed with the idea of attempting a time-series with both short-term (days) and long-term (annual) time scales. This objective was successfully achieved with long-term recoveries from last year's NeMO98 made at both Marker #33 and Cloud Vent (Marker N4), short-term recoveries from deployments made this year were also made at Marker #33, and Cloud Vent (Marker N4) on the East-side of Axial Volcano. New long-term deployments again were made at both of these two East-side sites. Unfortunately, only a single successful recovery from the ASHES Vent Field was made. This occurred at Gollum Vent, where two long-term deployed microbial traps were heroically recovered (in spite of the onset of foul weather conditions) and a fresh trap was deployed. Short-term recoveries from the ASHES area remains illusive, but may again be attempted next year, in addition to attempting continued long-term recoveries from each of the following locations where microbial traps have been previously deployed: Gollum, ROPOS, Hillock, Mushroom Vents.



Microbial samples collected were each independently processed. Microbial biomass preservation was achieved by quick-freezing in liquid nitrogen and storing on dry ice until return to the laboratory. These samples will be used for the direct extraction of nucleic acids. A series of sub-samples were also (i) cryo-preserved (again using liquid nitrogen quick-freezing) with 40% glycerol, and (ii) aliquots were stored at 4C, both for enrichment culture selection. Another series of sub-samples was fixed with 2.5% EM grade glutaraldehyde for examination with SEM and epifluorescence microscopy.



Laboratory Processing and Molecular Biological Analysis

Initially, all samples will be examined by epifluorescence microscopy in an effort to ascertain biomass estimates and examine morphological diversity. A subset of these will also be examined through SEM and an analysis of extractable lipids, which provides an estimate of microbial biomass and initial clues into community structure. The overall molecular biological strategy used will be essentially that of Moyer et al. (1994, 1995; 1998) with a few technical and logistical improvements. The first step will be the efficient and direct extraction of high molecular weight nucleic acids from quick-frozen samples. This will be followed by PCR amplification of SSU rDNAs using previously defined conditions to maximize the equal representation from each population contained within a respective community. The concept is to proportionally amplify or make several copies using the total genomic DNA from a natural community serving as the template for oligonucleotide primers that are complementary to universally conserved SSU rDNA sequence positions. Representative SSU rDNA amplification products are cloned generating a clone library. Clone libraries will then examined through the use of Amplified Ribosomal DNA Restriction Analysis or ARDRA and by using rarefaction as a metric for organismal diversity (Moyer et al., 1998). This approach, using tetrameric restriction enzymes, has been shown to detect >99% of the taxa (i.e., phylotypes) present within a model dataset with maximized diversity (Moyer et al., 1996). SSU rDNA sequences will also be subjected to phylogenetic analysis (using distance matrix and maximum likelihood algorithms) to estimate the affiliated ancestral lineage for each dominant community member thereby yielding clues as to their respective evolutionary history and potential physiology.



References:

Emerson, D., and C. L. Moyer. 1997. Isolation and characterization of novel iron-oxidizing bacteria that grow at circumneutral pH. Appl. Environ. Microbiol. 63:4784-4792.

Moyer, C. L., F. C. Dobbs, and D. M. Karl. 1994. Estimation of diversity and community structure through restriction fragment length polymorphism distribution analysis of bacterial 16S rRNA genes from a microbial mat at an active, hydrothermal vent system, Loihi Seamount, Hawaii. Appl. Environ. Microbiol. 60:871-879.

Moyer, C. L., F. C. Dobbs, and D. M. Karl. 1995. Phylogenetic diversity of the bacterial community from a microbial mat at an active, hydrothermal vent system, Loihi Seamount, Hawaii. Appl. Environ. Microbiol. 61:1555-1562.

Moyer, C. L., J. M. Tiedje, F. C. Dobbs, and D. M. Karl. 1996. A computer-simulated restriction fragment length polymorphism analysis of bacterial SSU rRNA genes: effacacy of selected tetrameric restriction enzymes. Appl. Environ. Microbiol. 62:2501-2507.

Moyer, C. L., J. M. Tiedje, F. C. Dobbs, and D. M. Karl. 1998. Diversity of deep-sea hydrothermal vent Archaea. Deep-Sea Res. II. 45:303-317.



2.3b Hydrothermal Fluid Microbiology - Julie Huber and Sheryl BoltonWe returned to sea this summer to continue our research on microbial communities in diffuse fluids on Axial Seamount. Along with Dave Butterfield, we are trying to quantify the diversity of microbes and their metabolisms in relation to the chemistry of the diffuse fluids over time.



We have successfully cultured mesophiles, thermophiles, and hyperthermophiles from all diffuse fluids collected during this cruise. We have over 100 positive enrichments (which require confirmation on land) at temperatures ranging from 23 C to 110 C in a wide variety of media, mostly anaerobic. One very exciting find was the culturing of a microbe at 110 C from Marker 33. This hyperthermophile (or group of hyperthermophiles) was successfully transferred six times while at sea. Currently, the highest known upper temperature limit for growth of a living organism is 113 C by a sulfur-dependent hyperthermophilic Archaea called Pyrolobus. Our microbe, growing anaerobically with elemental sulfur, was isolated from ~70 C fluid at Marker 33. We also found that it is producing large amounts of hydrogen and some carbon dioxide, as analyzed on the gas chromatograph by Andy Graham. The fact that we have found a microbe (or group of microbes) growing at such a high temperature, yet isolated from fluids much below its temperature of growth, strongly suggests that there is a hotter subsurface environment that these microorganisms are growing and thriving in. By studying this exciting microbe, we hope to learn more about the metabolic wonders of life at high temperature and the limits on life.

Additionally, quantitative enrichments (MPNs, Most-Probable-Number technique) were performed at a variety of temperatures from several sites. The table below contains the 95% confidence interval for the abundance of microorganisms that grow in the given media, given in microbes/liter. These data are preliminary and must be confirmed by microscopy on land.

  Sample Site Incubation

Temp (C)

Fluid

Temp (C)

Media Type Microbes/L
  Marker 33   90   78.0   Anaerobic, high organics   3000-96,000
  Marker 33   90   78.0   Anaerobic, low organics   600-8800
  Magnesia   55   5.6   Anaerobic, low organics   80-2400
  Bag City   90   23.4   Anaerobic, high organics   300-7600
  Bag City   90   23.4   Anaerobic, low organics   80-2400
  15m N of Nascent   90   25.8   Anaerobic, high organics   60-880
  Cloud N6   23   20.0   Aerobic, low organics   >48,000
  Gollum   90   22.3   Anaerobic, high organics   In Progress
  Marshmallow   90   71.8   Anaerobic, low organics   In Progress



We also obtained a number of discrete filtered fluid samples for molecular analysis using the hot fluid sampler. These filtered samples will be used in total community DNA analysis to determine microbial diversity and phylogeny, FISH (Fluorescence In-Situ Hybridization) to quantify and track certain microbes, and lipid analysis to quantify and determine the physiological state of microbes. We will also perform epifluorescent counts on all preserved fluid samples for microbial enumeration.



This combination of culturing, microscopic, and molecular techniques will help us determine how any changes in the fluid chemistry over the past year may be reflected in the microbial community structure. Additionally, we will continue to explore the microbial ecology of new diffuse sites and high temperature microbes found here at Axial Seamount.



2.3c Microbial Food Webs - UQAM Disciplinary Summary, Kim Juniper

Emphasis this year was placed on the consolidation and expansion of a study of the dynamics of microbial food webs. The overall goal of the study is to understand how the structure of vent food webs add structure and complexity as new vents and vent fields are colonized by increasing numbers of species. Particularly relevant are questions of the importance of vestimentiferan tube worms as a keystone species that create habitat for other organisms, and the relative importance of free-living and symbiotic microbial production as food sources for animals. Suites of faunal samples were collected from new vents in the East Rift Zone as well as from mature and senescent vent sites on adjacent older lava flows. Tissue samples from these organisms will be analyzed for stable carbon and nitrogen isotopes and for lipid biomarkers. Samples of microbial mat, biofilms and particulate organic material collected from each site will be similarly analyzed to permit matching of deposit and suspension feeding animals to their food sources, as well as the identification of trophic levels. The microbial and particulate samples will also be analyzed for ATP and total lipid content, as estimates of microbial biomass available for consumer organisms. Samples were also preserved for a molecular study of the diversity of microorganisms that constitute the animal food supply.



Samples collected came almost entirely from basalt-hosted diffuse vents, with the exception of samples from T&S vent at CASM and 2 collections from the sulfide worm habitat at ASHES. Planned collections of sulfide edifice tube worm communities in the ASHES field were missed because of the weather-related early termination of the dive program.



A vent-field scale study of the development of food webs was initiated at the Cloud site in the East Rift Zone. This study will examine spatial relationships between geological features, the location and intensity of venting and the colonization of individual vents, much as we have previously done for sulfide edifices. This is very much a mapping based project that seeks to discover spatial and temporal patterns that lead to the development of testable hypotheses. Video imagery and photographs collected during two dives are being used to develop a map of the distribution of individual organisms in relation to geological features and venting. This map will be incorporated into a small-scale GIS later this summer, to facilitate quantification of organism distribution and habitat. A short Imagenex survey was also perform during the video transects. We will attempt to integrate the resulting data into the map, despite problems with imprecision of navigation along and between transects. Mapping will be repeated in subsequent years to document changes in venting and community composition.



2.4 MACROBIOLOGY



2.4a Biology of Low Temperature Sites - Maia Tsurumi and Jean Marcus

1. Introduction

This biology program focused on two major themes: 1. succession (continued colonization through senescence) of the South Rift Zone vents, and 2. The regional distribution of species and populations of Axial Seamount. An important aspect of our sampling is to couple our collections with vent fluid chemistry. We were very successful with this approach last year, and during NeMO 1999 we were fortunate to extend this coordinated sampling. Of all our samples, there were only three which had no chemistry data.



2. Succession

a) Continued colonization.

We were intrigued last year to find three types of initial colonization at new vents on the new lava. These included: i) Small vestimentiferan recruits with three or four other known species, ii) Dense snails and iii) A mix of scale worm species and other polychaetes. This year we anticipated more homogeneous communities with tube worm recruits at all new vents. This prediction mostly held true. Tube worms occurred at all vents with the exception of the northern area where venting may have subsided, based on the general lack of vent fauna. For example, we found no fauna at Milky vent, although it supported scale worms and other polychaetes last year. Although tube worms were present and of a morphotype typical of new venting, they were not found in the abundance that we anticipated. In fact, this scarcity of tube worms forced us to suction sample at many locations because grabs were not possible.



We were able to repeat sample at all locations where we had samples last year. Further, we were able to sample two new vents on the southern portion of the South Rift Zone. This sampling revealed a species of nemertean worm seen after the 1993 CoAxial eruption, but not observed on the Axial 1998 flow last year. Although the lava in this area is of uncertain age, the tube worm morphology indicates that this is new venting. Preliminary sorting of samples shows that new species have continued the colonization sequence. There also is an observed shift in the relative abundance and biomass of some of the species. Next year should be interesting...



b) Senescence

Since venting is ephemeral, we have the opportunity to see what the vent community looks like at all stages of its "lifecycle". This is next to impossible to do in many other ecosystems because of the time scales involved. Until recently, studies of vent succession have focused on the initial and intermediate stages of the communities. For the last couple of years our lab has been working on understanding the entire sequence of succession from initiation to death. This requires opportunistic sampling of dying or dead vents. We obtained video, a suction sample, and one tube worm grab at a senescent vent and anticipate that this will help us to sort out the final sequence of the vent cycle.



3. Regional Character

Axial Volcano is one of the few places on the Ridge that allows us to study discrete, well-separated communities. A current question in vent ecology is how populations interchange among sites. We are working on better describing species distributions in a regional setting. Some vent species are very patchy and we are attempting to understand in an ecological framework why this is. To understand vent community dynamics within the caldera, it is crucial to sample as many intra- and inter-vent field assemblages as possible. To this end, we re-sampled, CASM and most of the South Rift Zone (North and South). In addition, we have selected one species of polychaete for a population genetics study.



4. A Final Comment

We would like to thank the NeMO 1999 science party for their encouragement, help, and interest. Also a big thank you to all the crew of the Thomas G. Thompson and the C.S.S.F. ROPOS gang.



Macrobiological Sample List from Low Temperature Sites

SOUTH RIFT ZONE

Suction samples

R483-1: Mkr 33, Suction for larvae over tube worms of R483-6

R488-5: Mkr 33, Suction for C. Levesque and J. Marcus

R488-6: Mkr 33, Suction for mat by C. Moyer, picked out large fauna

R488-7: Mkr N8, Snail, Suction for fauna (some animals to C.Levesque)

R488-14: Mkr N4 (Cloud), Suction for fauna

R488-15: Mkr N4 (Cloud), Suction for mat by C. Moyer, picked out large fauna

R488-16: Mkr N4 (Cloud), Suction for tube worms into flushing bottle

R491-5: Mkr 33, Suction for gastropods ~ 3 m from crack (C.Levesque has most animals)

R491-13: Nascent , Suction over tube worm bush for larvae

R491-14: Nascent, Suction where tube worm grab was taken (C.Levesque has most animals)

R492-1: Joystick, Suction because too few tube worms to grab

R492-2/3: Joystick, Suction for mat and bag creatures by C. Moyer, picked out large fauna

R495-16/17: Mkr 113, Suction for mat by C. Moyer, large fauna picked out

R495-4: Non-vent, Suction on new lava for diatom mat

R495-35: Crevice, Suction sample of old worms



Tube worm grabs

R483-6: Mkr 33, area in front of Time Lapse Camera

R491-16: Nascent, from spot where MTR (#4108) was and 3 other spots close by (some animals to C.Levesque)

R491-18: Mkr N41, from spot where MTR (#4126) was

R491-20: Old Flow, old worms on old lava

R491-24: Oldworms, old worms on old lava

R492-6: Coquilles, old worms on old lava

R492-10: Bag City, large worms (up to 1 m), new vent?

R496-3: Mkr 113, from top of pillar, where Moyer's microbial traps #20&21 were

R501-17: Crevice, older worms from intermediate (?) lava



Animals from C. Moyer's microbial traps

R483-3/7: Mkr 33, Animals from C. Moyer's microbial traps (#9&12)

R496-1/2: Mkr 113, Animals from Moyer's microbial traps (#20&21)



CASM

R497-9: base of T & S, where Vemco was recovered and MTR deployed



ASHES

R502-15: Mkr I, suction sample for meiofauna and water chemistry



Water Chemistry

Water samples from all sites listed above except Snail vent (mkr N8), Mkr N41, Old Flow and Crevice.



Time Lapse Camera

TLC was retrieved and redeployed at Mkr 33. Tube worm grab from Marker 33 sampled the photographed area.

2.5 HYDROTHERMAL DEPOSITS - Steve Scott



There are two types of hydrothermal deposits in the caldera of Axial Volcano: Fe-oxyhydroxides and sulfide-sulfate. Both were investigated and sampled during NeMO 99.



Fe-oxyhydroxide deposits

Analyses of Fe-oxyhydroxide samples taken from Steve Mound, Gollum Vent and south of ASHES in 1998 show the material to be silica poor ferrihydrite (nominally Fe5HO8.4H2O) together with amorphous silica. Ferrihydrite is commonly found in soils, oxidized mine wastes and other iron-rich environments, and now also on the deep seafloor (Boyd and Scott, 1999, Can. Mineralog.). The material is essentially amorphous although it can have short-range crystallographic order. It is of biogenic origin, clearly coating bacteria as seen in SEM and STEM images.



During NeMO 99, ferrihydrite deposits were sampled at Oxide Vent (inactive); at Naaz, south of ASHES (inactive); and, at the west wall of the caldera near ASHES (active). Red iron oxides and bacterial fluff were also observed at a few other places, most notably on the new lava in the south rift zone. The Naaz site is a new discovery of clusters of 10-15 cm diameter x 20-60 cm high conical structures that look like termite mounds made of ferrihydrite and of anhydrite coated by ferrihydrite. Naaz covers an area of about 5m (east-west) x 15m (north-south) and lies about 5 m south of Crack Vent. The distribution of the mounds within a cluster and in the peripheral regions where mounds are widely dispersed appears to be controlled by fractures in the otherwise relatively smooth sheet flow surface. Both vent fluids (6oC maximum temperature) and consanguineous oxide-coated bacteria were sampled at the west caldera wall. This will enable a determination of partitioning of elements between fluid and solids, an important step in the study of this biomineralization process.



Sulfide-sulfate spires

The CASM T & S spires, discovered during the 1998 NeMO expedition, were sampled. The five samples include different mineralogical types and both actively venting and inactive areas. A short distance north of T & S, on the east wall of the CASM rift, there is a large pile of oxidizing sulfide talus whose source was not thoroughly investigated. There is also evidence (red staining) of hydrothermal activity in some of the talus blocks at the foot of the north wall of the caldera. One sample each from Castle and Flat Top were lost. The Castle sample, taken from the top of the structure, appeared to be oxidized.



Two recovered HOBOs (Hell Vent and CASM T & S Vent) had ~1 mm thick sulfide deposits on their probe sheath. By determining the mass of precipitated material/ unit area/time the HOBO was deployed, the precipitation rate can be calculated, assuming it is linear with time. Furthermore, this rate can be known as a function of fluid chemisty (the vent fluids were sampled) and temperature (from the HOBO record).



Suction sampling within the Marker 33 vent revealed the presence of sulfides (pyrite + ?chalcopyrite) and anhydrite coating a cm-size piece of basalt and impregnating the basalt=s vesicles. The anhydrite and possible chalcopyrite indicate much higher temperatures than are now observed at Marker 33. It is postulated that the vent system was much hotter just after the January 1998 eruption than it is now.



2.6 ROCK SAMPLING AND PETROLOGIC STUDIES - M. Perfit and J. Chadwick

Introduction:

The objectives of the rock sampling program were four-fold: 1. complete detailed sampling and mapping of the 1998 lava flow to aid in the identification of its boundaries and extent, 2. determine the compositional and petrologic heterogeneity of the 1998 flow and spatial variability, 3. compare the composition of the 1998 flow to that of the surrounding older flows to evaluate temporal changes in magma source, 4. recover samples from Axial's north and south rift zones by rock corer to investigate temporal and spatial variations in magma genesis on a regional scale.



The initial sampling and subsequent chemical analyses of basalts from the 1998 flow showed that it was very similar in composition to basalts previously recovered from the caldera. In particular, the 1998 flow has a composition like that of young looking flows from the CASM site. All are normal mid-ocean ridge basalts (N-MORB) that have slightly elevated K2O contents (and other incompatible trace elements) compared to other MORB from the Juan de Fuca Ridge. Of a total of 16 samples analyzed from the 1998 flow, fourteen have nearly identical compositions; two from the southernmost part of the flow (Recovered on dive R465) are slightly more evolved. This suggests some chemical differentiation may have occurred during the diking event or that the magma source is chemically heterogeneous.



The detailed sampling completed this year will allow us to place more rigorous constraints on the chemical composition (and variability) of the 1998 flow as well as to determine the composition of older flows in contact with the new flow. Assuming there are significant compositional differences between the 1998 flow and surrounding older flows, we plan to use the geochemical data in conjunction with the observational data and mapping efforts to generate a geological map of the 1998 flow. In addition, comparison of the major and trace element composition of the new flow to older flows can constraints on magma chamber temperature changes and magmatic evolution. Measurement of U-series isotopes (by K. Rubin) will provide information about the age the magma chamber for the different flow units in the Caldera and possibly allow us to relate young flows from CASM and ASHES to the 1998 flow.



From a regional standpoint, the 1998 flow is very important because it provides us with another 'zero age' flow from the JdF to use in comparing an contrasting the current sources and melting parameters along the ridge. Initial analyses of the 1998 flow confirm that the source of Axial magmas is slightly more enriched than the mantle source of southern JdF lavas and quite a bit more enriched compared the source for CoAxial Segment magmas. This years 55 rock cores now gives us more than 115 rock core localities along Axial's north and south rift zones. These samples, together with the samples recovered in the caldera, will be used to investigate the spatial and temporal variability of magma genesis at the entire volcanic edifice that has been created by the interaction of Axial melt anomaly and the Jdf ridge.



ROPOS Operations

During this cruise, we recovered 52 samples from ROPOS dives. Many of these samples were 'opportunistic' in that they were recovered during biological grabs or suction samples for fluids or biology. Most of these consisted of glass chips and were primarily from well documented vent sites within the 1998 flow. In order to facilitate recovery of very glassy samples which can be nearly impossible to grab with ROPOS manipulators, we developed small rock wax corers that were deployed on ROPOS for the first time. These 'chapstick' cores were quite effective in sampling the glassy surfaces of lavas and proved to be a quick, easy way to get additional rock samples. Deployed rock cores were placed into a small mesh bag ('purse') fashioned to attach to the front of the bioboxes or fluid sampler. This bag was also used to recover temperature probes and additional rock samples. The glassy crusts from all large 'whole rock' lava samples were sub-sampled to be hand-carried with core samples. The highlight of the rock sampling came during the biogeology dive R501 during which two E-W traverses were made across the 98 flow while detailed volcanological and geological observations were made. During the dive a record total of 16 rock samples were recovered using the manipulators, rock cores and suction sampler. Precisely located samples at lava flow contacts should aid us in distinguishing flow units.



Rock Coring Operations:

Rock coring during the NeMO 1999 cruise continues the effort begun during NeMO 1998, when 49 successful rock core attempts were made. An additional 61 core samples were collected this year, for a total of 109 for the two cruises. Samples were collected principally in the south rift zone of Axial Seamount (48 samples), from just off the flank of the seamount south to approximately 45 degrees 39 minutes latitude. The remainder of the samples were collected along the Vance segment of the Juan de Fuca Ridge (3 samples), the southwest rift zone (4 samples), and the north rift zone (6 samples). The glass from this effort will be analyzed for major elements using electron microprobe and for trace elements using laser ablation inductively coupled plasma mass spectrometry (ICPMS) and X-ray fluorescence (XRF) techniques. It is hoped this study will lead to a better understanding of the relationships between the seamount, rift zones, and Juan de Fuca ridge, as well as insight into the magmatic, volcanic, and tectonic processes that have created the rift zones.



Rock coring activities took place in periods between ROPOS dives, and were undertaken on the CTD wire on the starboard side of the ship. The corer has 7 wax-tipped cups which collect small shards of basaltic glass from the ocean floor. The corer was sent down on the wire at approximately 60 meters/min. until the winch monitor tension dropped considerably, indicating contact with the floor. An additional 5-10 meters of wire was spooled out prior to reversal of the winch and retrieval of the corer, again at 60 meters/min, after a short period of slow (10 meters/min) retrieval to get the corer off the bottom. This method resulted in a successful retrieval of ocean floor materials in all attempts, although no glass was collected in 2 attempts (sediments only).



3 Non-ROPOS OPERATIONS



3.1 Mooring Deployments /CTD's /XRF Analysis - Jim Gendron

Early in the cruise, one CTD cast was completed South-East of the venting area on the new lava. The water was collected around 1300 meters as a non-plume background sample for many of the groups onboard.



During the cruise, the hot fluid sampler was used to collect 22 filters for XRF analysis. These samples will be analyzed when we return to Seattle. Also collected on the fluid sampler were about 30 samples for SEM and 5 samples for Particulate Organic Carbon (POC). A total of four, good Niskin samples were collected, one each over Castle, Magnesia, Cloud and CASM vents. Another was attempted over Hell vent, but it appears to have pre-tripped sometime during the dive.



Near the end of the cruise, three MTR moorings were deployed. One was at ASHES and two were close to the new lava vent area.

Moorings Deployed Summer of '99

Lat Long UTM X UTM Y

(deployed from Thompson)

99T50 -129.9867 45.9417 423520 5088039

99T49 -129.9867 45.9250 423497 5086188

99T51 -130.0050 45.9300 422083 5086761

(deployed from Wecoma)

99T52 -130.0245 45.9053 420536 5084039

99T53 -130.0142 45.8912 421317 5082455

99V110 -130.0000 460.550 422645 5100644

99V111 -129.9135 45.9493 429201 5088823

99V112 -130.0650 45.9205 417417 5085766

99V113 -130.0017 45.9167 422322 5085276



3.2 NeMO'99 Website and Public Outreach - Nicole Nasby, Greg Pillette, Andra Bobbitt http://www.pmel.noaa.gov/eoi/nemo/

The goal was to create an educational web site that would be for use by students and teachers, primarily at the secondary level. The web site offered daily updates on the cruise and allowed interested individuals to follow the progress of the scientific expedition at Axial Volcano. The site was set up to provide daily updates that included a scientific report on the latest activity. There was also a weekly science summary written by the Chief Scientist. The web site included a daily personal perspective section that highlighted an individual from the scientific party of ship's crew, and a daily log written by the "teacher at sea" Greg Pillette. The last component was an interactive question and answer section so that the public could interact directly with the scientific staff.



Each of these sections was coordinated by Nicole Nasby and relevant digital images were included. They were sent from the ship to HMSC and were added to the web site on a daily bases by Andra Bobbitt in Newport. This material was then presented to the general public at HMSC, each day, by the teacher on shore, Steve Babcock. Feedback and question from the public were sent to the ship to be answered by the scientific staff.



We have received positive feedback from the web site maintained this year as well as the one from NeMO '98. We have heard from teachers who have used last year's site in a classroom situation. Unfortunately classes are not in session at this time, so there is little feedback from teachers and students on the NeMO '99 site at this point, but we did get several questions posted to the web site indicating a positive reaction. As with last year, many of these questions were from family members of science and ROPOS personal.



4 NAVIGATION



4.1 Navigation Overview - Susan Merle and Bill Chadwick

All ROPOS dives were navigated using long-baseline transponder nets with the Seascape navigation software. The navigation computer had three main inputs into Seascape to aid in ROPOS navigation: P-code GPS from the R/V Thomas Thompson, ROV depth data provided by the ROPOS sensor and the PS8000 data from the range meter. Transponders were already in place for the ASHES net and the North Rift Zone (NRZ) net. Deployed last year during NeMO'98, the expendable transponders should have five year lifetimes. The transponders for those two nets only needed to be enabled. A third net, on the South Rift Zone (SRZ), was added with three recoverable transponders. Unfortunately, due to hardware problems with the NOAA PS-8000, the calibration of the SRZ Net was poor and this net was only marginally useful.



Once the cage reached its final depth and ROPOS drove to the seafloor, the cage depth was manually entered into the Seascape program and was held constant, unless the wire out for the cage changed during the dive. The range meter was attached to the top of the cage, was hard-wired to the hydro lab and triggered by Seascape on the navigation computer. Cage and ROV fixes were generally scattered with RMS errors of about 30 meters. Navigation fixes are recorded in latitude/longitude and UTM x/y (in meters) in the log files and were processed by Susan Merle in the IDL programs navedit2 and navedit3 (written by Bill Chadwick).



In 1999 navigation was somewhat less reliable than in 1998. During several dives the 11.0 transponder in the ASHES net was accidentally disabled and this caused the 1999 navigation fixes to be shifted to the west or south of the 1998. Since '98 navigation was more robust, no old vent or marker positions were changed, even if the '99 positions did not agree with those of '98. After the '99 cruise a few of the mysterious navigation problems we had were identified. The NOAA PS-8000 on El Guapo was determined to have a bad microprocessor, which created problems with the SRZ net calibration. Another mystery solved concerned the fact that it seemed that transponder 11.0 in the ASHES net would shut itself off. It was discovered on Bill Chadwick's Cleft cruise in September '99 that the 11.0 had a disable code of C, not B like all the other transponders. This would not have been a problem, if the ROPOS transponder we tried to use, unsuccessfully, had not had an enable code of C. So, we would try to enable the ROPOS transponder, unwittingly disabling the 11.0 transponder in the ASHES net.



The dive plots will be a useful guide to determine navigation gaps, bad navigation, etc. When acoustic nav was not available GPS ship nav was used. Plotted positions are for the reference point, about 37 meters forward of the stern at the GPS antennae.



4.2 FINAL CALIBRATED TRANSPONDER POSITIONS



North Rift Net

Transponder UTM-X (m) UTM-Y (m) Latitude Longitude Depth
9.5 420814.65 5098603.9 46 02.1857' 130 01.3988' 1433.9
10.5 422722.92 5097596.31 46 01.6548' 129 59.9096' 1395.43
8.0 420055.52 5095969.44 46 00.7580' 130 01.9608' 1377.93
7.5 422074.85 5094971.24 46 00.2330' 130 00.3862' 1294.46


ASHES Net

Transponder UTM-X (m) UTM-Y (m) Latitude Longitude Depth
11.5 424283.25 5087181.51 45 56.0418' 129 58.6011' 1305.4
10.5 424221.58 5084426.79 45 54.5540' 129 58.6227' 1340.36
9.5 422490.35 5086188.55 45 55.4937' 129 59.9789' 1324.67
11.0 422556.72 5088014.47 45 56.4800' 129 59.9453' 1330.85


South Rift Net

Transponder UTM-X (m) UTM-Y (m) Latitude Longitude Depth
10.0 423771 5084021 4554.33 ' 12958.968 ' 1471.69
8.5 421721 5082432 4553.46 ' 1300.54 ' 1401.68
12.5 422134 5084021 4554.324 ' 1300.234 ' 1492.90


4.3 VENTS AND MARKERS LOCATION TABLE
 Vents/Markers Area Longitude Latitude UTM X UTM Y Depth
 Bag City SRZ -129.98926 45.91622 423284 5085214 1534
Bob NRZ -130.01283 46.03892 421629.2 5098870.2 1590
Castle North SRZ -129.97990 45.92613 424022.7 5086305.8 1520
Circ North SRZ -129.98165 45.92592 423887 5086283
Cloud North SRZ -129.98156 45.93335 423904 5087110 1524
Coquille North SRZ -129.99306 45.91753 422991 5085365
Crack ASHES -130.01355 45.93330 421424 5087135 1547
Crevice SRZ -129.99040 45.91110 423175 5084648 1538
Dave's ASHES -130.01377 45.93352 421408.3 5087158.6
Dying North SRZ -129.99185 45.91685 423083.7 5085286.4
Easy North SRZ -129.98472 45.94533 423676.5 5088443.2 1532
Fe-Hyde ASHES -130.01378 45.93298 421406 5087099.7
Gollum ASHES -130.01358 45.93358 421422 5087166.1 1546
Hairdo ASHES -130.01398 45.93350 421390.7 5087156.8 1546
Hell ASHES -130.01423 45.93330 421372 5087135 1544
Hillock/Phoenix ASHES -130.01398 45.93325 421390.9 5087130.4 1544
Inferno ASHES -130.01390 45.93355 421397.2 5087162.2 1547
Joystick SRZ -129.98856 45.91884 423341.5 5085505
Joystick2 SRZ -129.98851 45.91875 423345 5085495
Magnesia North SRZ -129.98493 45.94623 423660.7 5088544.7 1530
Marshmallow ASHES -130.01362 45.93370 421420.4 5087179 1547
Medusa ASHES -130.01393 45.93335 421394.7 5087141.1 1546
Milky North SRZ -129.98475 45.94514 423673 5088424 1527
Minisnow North SRZ -129.98422 45.94262 423711 5088141 1522
Mkr-I ASHES -130.01367 45.93370 421416 5087180
Mkr-108 Vent North SRZ -129.98303 45.92865 423784 5086589 1520
Mkr-108b Vent North SRZ -129.98319 45.92881 423772 5086608 1520
Mkr-113 Vent North SRZ -129.98827 45.92260 423370 5085922 1524
Mkr-2 ASHES -130.01397 45.93330 421392 5087136
Mkr-21 ASHES -130.01358 45.93360 421422 5087168
Mkr-33 Vent North SRZ -129.98225 45.93327 423850.3 5087101.1 1524
Mkr-36 SRZ -129.98161 45.93437 423284 5085214
Mkr-42 Joystick -129.98856 45.91884 423341.5 5085505
Mkr-D ASHES -130.01393 45.93325 421399 5087129
Mkr-L ASHES -130.01432 45.93333 421365 5087140
Mkr-N1 North SRZ -129.98408 45.93980 423718 5087828 1521
Mkr-N2 North SRZ -129.98475 45.94514 423673 5088424
Mkr-N3 North SRZ -129.98520 45.94380 423637 5088278
Mkr-N4 North SRZ -129.98167 45.93342 423896 5087119 1523
Mkr-N41 North SRZ -129.98138 45.93622 423922.4 5087428.2 1520
Mkr-N44 North SRZ -129.98483 45.93947 423658 5087792
Mkr-N5 (Flattop) North SRZ -129.97960 45.92610 424043 5086306 1516
Mkr-N6 North SRZ -129.98160 45.93340 423901.4 5087116 1524
Mkr-N7 North SRZ -129.98190 45.93930 423886 5087774 1520
Mkr-N8 North SRZ -129.98190 45.93320 423877 5087088
Mkr-N9 North SRZ -129.98423 45.94260 423710 5088141 1522
Mushroom ASHES -130.01380 45.93360 421405.3 5087167.9 1548
Nascent North SRZ -129.98152 45.93577 423911 5087378 1520
Old Flow North SRZ -129.98170 45.93645 423897.5 5087454.7 1534
Old Worms North SRZ -129.98331 45.94510 423785 5088418 1528
Ouzo North SRZ -129.98468 45.94582 423679.6 5088496.8 1526
Oxide North SRZ -129.98508 45.94545 423647.9 5088456.4 1543
Pillar North SRZ -129.98542 45.92270 423591 5085929.1
Porkchop ASHES -130.01422 45.93332 421373 5087136 1547
Roof North SRZ -129.98448 45.94250 423689.8 5088129.1 1523
ROPOS ASHES -130.01405 45.93328 421386.1 5087134.1 1547
Snail North SRZ -129.98188 45.93317 423878.6 5087089.7 1525
Snow North SRZ -129.98245 45.92712 423827 5086417
Snow Blower North SRZ -129.98407 45.93987 423719 5087835 1521
Steve Mound ASHES -130.01342 45.93325 4214348 5087128.6
Styx ASHES -130.01370 45.93328 421412.2 5087132.2
The Pit North SRZ -129.98408 45.93975 423718.2 5087823.2 1521
Tombstone ASHES -130.01133 45.92948 421590 5086597
Tunnicliff ASHES -130.01582 45.93367 421248.7 5087178 1544
Virgins Daughter ASHES -130.01340 45.93375 421436 5087184
Virgin ASHES -130.01348 45.93365 421430 5087174 1547
White ASHES -130.01363 45.93373 421419 5087182.9
91Vent NRZ -130.01242 46.03860 421661.4 5098834.3
Shepherd CASM -130.02668 45.98990 420486.4 5093373.6 1583
South Casm CASM -130.02625 45.98870 420518.6 5093304.6 1583 


5 NeMO OPERATIONS



5.1 1998 Dive Dates and Locations

ROPOS Dives R460 - R480



  Dive # Date Location
R460 JD 240-241

Aug 28-29

SE Caldera SRZ:

Mkrs N3, 33; Milky, The Pit, Cloud Vents

R461 JD 241-243

Aug 29-31

SE Caldera SRZ:

Rumbleometer; Mkrs 108,33,113; Cloud, Sulfide, Castle, Circular Vents

R462 JD 243 - 244

Aug 31 - Sept 1

SE Caldera SRZ:

Mkr-33, Cloud Vent

R463 JD 244 - 245

Sept 1 - 2

SE Caldera SRZ:

Easy, Milky Vents; (+ Imagenex survey)

R464 JD 245

Sept 2

SE Caldera SRZ:

Oxide, MiniSnow, The Pit, Snail, Mkr-108, Mkr-113, Castle Vents

R465 JD 246

Sept 3

South Rift Zone:

reconnaissance survey

R466 JD 247

Sept 4

ASHES:

Hell, ROPOS, Hillock/Phoenix, Hairdo and Inferno Vents

R467 JD 248 - 249

Sept 5 - 6

North Rift Zone:

Extensometers; Bob Vent: (+Imagenex survey)

R468 JD 250

Sept 7

ASHES:

Gollum, Hell, ROPOS, Hillock/Phoenix, Crack Vents

R469 JD 250 - 251

Sept 7 - 8

ASHES:

Medusa, Mushroom, Marshmallow, Gollum, Daves Styx and Fe-Hyde Vents; (+Imagenex survey)

R470 JD 251

Sept 8

North Rift Zone:

Extensometers

R471 JD 252

Sept 9

ASHES:

Gollum, Mushroom, White, Inferno, Hell Vents

R472 JD 252

Sept 9

ASHES:

Steve Mound, Hell, Phoenix, Medusa, Inferno Vents

R473 JD 253 - 254

Sept 10 - 11

SE Caldera SRZ:

Easy, Milky, Roof, The Pit, Snowblower, Mkr-33, Mkr-108, Cloud, Castle Vents; (+Imagenex survey)

R474 JD 255

Sept 12

SE Caldera SRZ:

The Pit, Milky Vents; Rumbleometer; Lava Flow Mapping Traverses

R475 Dive aborted
R476 JD 256 - 257

Sept 13 - 14

SE Caldera SRZ:

Magnesia, Easy, Old Worms, Milky Vents;

Lava flow traverses; (+ Imagenex survey)

R477 JD 258

Sept 15

SE Caldera SRZ:

Rumbleometer; Mkr-33 Vent

R478 JD 258

Sept 15

SE Caldera SRZ:

Mkr-33, Mkr-n4, Cloud, Nascent Vents

R479 JD 259 - 260

Sept 16 - 17

Northern traverse along caldera wall:

ASHES:

Hell, Virgin, Mushroom, Medusa, Inferno Vents; (+Imagenex survey)

R480 JD 261 - 262

Sept 18 - 19

North Rift Zone and Northern Caldera Wall:

Extensometers; CASM (Shepherd?) Vent







5.2 1999 Dive Dates, Locations and Tasks

ROPOS Dives R482 - R503

 Dive # Launch time Recovery time Dive Time (in the water) Bottom Time  NeMO 1999 Dive Locations and Tasks
R482 (6-22) JD173 - JD173 - 1530 1 hr 13 min 0 hr Didn't get to the bottom. Aborted at 400 m - fluid sampler not communicating
R483 (6-23) JD174 - 1000 JD174 - 1925 9 hr 25 min 7 hr 42 min Mkr-33 - 10 Samples: Suction, bio, gastights. Recovered Bactraps MTR and Osmo deployed in '98. Deployed 2 Fe OsmoAnalyzers. (Dive aborted, kink in tether).
R484 (6-24) JD175 - JD175 - 2023 10 hr 8 min 6 hr 30 min Mkr-33, Cloud Vent - 5 Samples: Suction, bio. Recovered MTR at Cloud and Time Lapse Camera at Mkr-33, both deployed in '98. (Dive, lost hydraulic fluid)
R485 (6-25) JD176 - 0700 JD176 - 1330 6 hr 30 min 1 hr Mkr-33 - 6 HFS Samples. This was the 2-part dive that we didn't change the Dive number for, even though Ropos came to the surface, therefore only 1 hour of bottom time. (Dive aborted, oil resevoir empty)
R486 (6-26) JD177 - JD176 - 0535 2 hr 32 min 0 hr Didn't get to bottom. (Dive aborted, hydraulic failure)
R487 (6-26) JD177 - 1531 JD177 - 1935 4 hr 4 min 2 hr 30 min Mkr-33, Cloud Vent - 15 Samples: HFS, gastight. Hydraulic fluid levels low, ROPOS went back to the cage and was pulled from Mkr-33 to Cloud by the ship. Did several HFS fluid samples of the plume in this fashion.
R488 (6-27) JD178 - 0428 JD179 - 0145 21 hr 17min 18 hr 30 min Mkr-33, Cloud, Snail, Nascent, MiniSnow, Mkr-N3,Milky, Magnesia, Imagenex - 33 Samples: HFS, suction, gastights. (Dive aborted, hydraulic oil pressure lost)
R489 (6-29) JD180 - 0620 JD180 - 0756 1hr 36 min 0 Didn't get to bottom
R490 (6-29) JD180 - 1945 JD180 - 2016 31 min 0 Didn't get to bottom. El Guapo was still plugged in after calibration, therefore transponders were not responding. (Dive aborted, navigation problem)
R491 (6-30) JD181 - 0145 JD182 - 1914 41 hr 29 min 38 hr 55 min Mkr-33, Cloud, Nascent, Old Flow, Old Worms, Magnesia, Imagenex - 28 Samples: Suction, bio, gastight, rocks. Video/photo survey of Cloud, Deployed BioOsmo Mkr-33, MTR & Vemco Mkr-N6, MTR Mkr-N41. Recovered MTR Nascent, bactraps Mkr-N4. (Dive finished)
R492 (7-2) JD183 - 0855 JD183 - 2324 14 hr 29 min 11 hr 18 min Joystick, Coquilles, Bag City - 14 Samples: Suction, Rocks, gastight, bio. Geology reconnaissance dive of '98 lava flow, south of area surveyed in '98. Eastern and western extents surveyed. Elevator recovered. (Dive ended to recover the elevator)
R493 (7-3) JD184 - 0635 -on btm JD184 - 1158 5 hr 23 min 2 hr 13 min Extension of Dive R492 - 2 Samples: rocks. Didn't go to any vents. (Dive aborted, lost telemetry, snap-loading on umbilical)
R494 (7-4) JD185 - 0227 JD185 - 1930 17 hr 2 min 14 hr 11 min Started south of SeaBeam anomaly, moving north to map limits of ,98 lava flow. Samples: 7 rock samples, bio, suction. No acoustic nav this dive - south net not working with north net. Fissure discovered! (Dive aborted, lost telemetry)
R495 (7-5) JD186 - 0330 JD187 - 0158 22 hr 28 min 21 hr Mkr-33, Mkr-109, Mkr-113, Castle, Joystick, Coquilles, Bag City, Fissure on SRZ - 37 Samples: HFS, Suction, gastight, sulfides, nisken.
     Recovered Vemco and deployed Osmo at Mkr-33. Geological traverse of SRZ. (Dive aborted, telemetry problems)
R496 (7-6) JD187 - 1820 JD188 - 1705 22 hr 45 min 19 hr 18 min Mkr-113, Rumbleometer freed from lava, Imagenex - 4 Samples: Bactraps, biology, rumbleometer. Recovered bactraps at Mkr-113, bio sample at Mkr-113, rock chips from rumbleometer. (Dive ended normally)
R497 (7-8) JD189 - 0010 JD189 - 1939 19 hr 29 min 10 hr 9 min Extensometers at NRZ, CASM - 21 Samples: suction, sulfide, bio, gtb. Elevator was recovered. At CASM collected 17 samples, VEMCO, hobo probes. One extensometer placed in elevator. (Dive aborted, sub lost power)
R498 (7-8) JD189 - 2215 JD190 - 0349 5 hr 39 min 3 hr 34 min Extensometer Recovery on NRZ - 3 remaining extensometers deployed in 98 were placed in elevator and all four were recovered. (Normal end of dive)
R499 (7-9) JD190 - 1509 JD190 - 2100 5 hr 51 min 4 hr 4 min Magnesia, Old Worms - 6 Samples: HFS, niskin. Also deployed osmosampler at Magnesia. (Dive aborted, Leak in HFS followed by hydraulic leak in ROPOS)
R500 (7-9) JD190 - 2314 JD191 - 1440 15 hr 26 min 13 hr 20 min Old Worms, FeO, Mkr-N41, Cloud, Mkr-33, Imagenex - 25 samples: HFS, suction gtb, niskin. Video mapping of Cloud. (Normal end of dive)
R501 (7-10) JD191 - 2140 JD192 - 1711 19 hr 31 min 14 hr 14 min Mkr-33, Crevice Vent ,and geologic traverse - 17 Samples: 13 rocks, suction, bio, Niskin. Deployed TLC at Mkr-33. (Normal end of dive)
R502 (7-12) JD193 - 0030 JD193 - 1520 14 hr 50 min 12 hr 17 min ASHES - 38 Samples: HFS, suction, gtb, niskin. Recovered HOBO at Hell and deployed one. Retrieved MTR at Gollum. Also recovered HOBO at Inferno. (Normal end of dive)
R503 (7-12) JD193 - 2015 JD194 - 0150 5 hr 35 min 3 hr ASHES - 3 Samples: bactraps and 2 hobos. Osmosampler deployed and melted at Inferno, Hobo deployed at Inferno and Virgin. Bactraps recovered and deployed at Gollum. Hobo recovered at Inferno. (Dive ended due to weather)




5.3 Experiments Deployed and Recovered - 1998 and 1999

Vent/Marker Experiment/Instrument Dive Deployed Dive Recovered
Mkr-33 MTR 4130 R461 R483
Mkr-33 BacTrps#9&12 R462 R483
Mkr-33 OsmoSampler R477 R483
Mkr-33 Time Lapse Camera R478 R484
Mkr-33 BacTrps#40,41,42,43 R484 R491 (40, 43)
Mkr-33 Grn&Cu OsmoSampler R483 n/a *
Mkr-33 2 Fe OsmoAnalyzers R483 R491
Mkr-33 BioOsmoSampler R491 n/a *
Mkr-33 OsmoAnalyzer R495 n/a *
Mkr-33 Time Lapse Camera R501 n/a *
Mkr-113 VEMCO Alvin 3245(deployed) R461(repositioned) R495
Mkr-113 BacTrap#4 R461 n/a *
Mkr-113 BacTraps#19,20,21 R464 R496 (20,21)
Mkr-N2 Milky OsmoSampler R474 99 dive#?
Mkr-N2 Milky BacTrap#35 R476 n/a *
Mkr-N4 Cloud BacTraps#14,15 R462 R484
Mkr-N4 Cloud BacTraps#36,37,38,39 R484 R491
Mkr-N4 Cloud MTR 0942 R461 R484
Mkr-N6 Cloud MTR 3157 & VEMCO 1108 (taped together) R491 n/a *
Mkr-N41 MTR 4126 R474 R491
Mkr-N41 MTR 3041 ? R491 n/a *
Bag City Vent MTR R495 n/a *
Castle Vent BacTraps#22,23,24 R464 n/a *
Crack Vent VEMCO Alvin 3245(deployed) R461(repositioned) n/a *
Easy Vent BacTrap#17 R463 n/a *
Gollum Vent MTR 4128 R471 R502
Gollum Vent BacTrap #45 R503 n/a *
Gollum Vent BacTraps#32,33,34 R471 R503 (2 -#'s?)
Hell Vent Hobo R502 n/a *
Hell Vent Hobo 98 dive #? on osmo 502
Hillock/Phoenix BacTraps#25,26 R466 n/a *
Inferno Vent Hobo 128 R503 n/a *
Inferno Vent Osmosampler (didn't work) R503 n/a *
Inferno Vent HOBO 137 Alvin 3245 R502
Inferno Vent VEMCO 98-223 Alvin 3245 n/a *
Magnesia OsmoSampler R499 n/a *
Milky Vent BacTrap#35 98 dive #? n/a *
Mushroom Vent BacTraps#29,30,31 R471 n/a *
Nascent Vent MTR 4108 R478 R491
Nascent Vent MTR 3175 R491 n/a *
ROPOS Vent BacTrps#27,28 R466 n/a *
Shepherd Vent CASM VEMCO 98-1113-214 R480 R497
Virgin Vent Hobo 130 R503 n/a *
Virgin Vent HOBO 129 Alvin3246 R503
T&S Spires CASM HOBO 130 R480 R497
T&S Spires CASM MTR 3017 R497 n/a *
* non applicable - experiment not recovered yet.












5.4 1999 ROPOS Samples - Dives R483 - R503

   Dive R483 SAMPLES
 Sample Vent Mkr Long Lat UTM X UTM Y UTCTime Description Investigator Subsamp
SS-j1_dfl_R483-1 Mkr-33 V 33 -129.9822 45.9333 423850 5087101 1257 Fluid over tube worms - for about 1h Tunnicliffe/Marcus
  SS-j2_bio_R483-2 Mkr-33 V 33 -129.9822 45.9333 423850 5087101 1423 polynoids - gastros - rock Juniper/Levesque Levesque/ Perfit
Bactrp_R483-3 Mkr-33 V 33 -129.9822 45.9333 423850 5087101 1515 Bacteria trap #9 retrieved - rock Moyer Tunnicliffe/Marcus/
SS-j3_bio-R483-4 Mkr-33 V 33 -129.9822 45.9333 423850 5087101 15:43 suction stopped - maybe nothing in sample - rock Juniper/Levesque Perfit
Bactrp_R483-5 Mkr-33 V 33 -129.9822 45.9333 423850 5087101 15:48 Bacteria trap #12 retrieved - rock Moyer Tunnicliffe/Marcus/
Bio_R483-6 Mkr-33 V 33 -129.9822 45.9333 423850 5087101 16:10 Tube worm grab Marcus McHugh /Levesque
Gtb-6_R483-7 Mkr-33 V 33 -129.9822 45.9333 423850 5087101 16:38 Gas tight bottles #6 fired Evans Butterfield /Lilley
Gtb-7_R483-8 Mkr-33 V 33 -129.9822 45.9333 423850 5087101 16:38 GTB #7 fired Evans Butterfield/ Lilley
MTR_R483-9 Mkr-33 V 33 -129.9822 45.9333 423850 5087101 16:50 MTR#4130 recovered Embley Butterfield/ Lilley
Osmo_R483-10 Mkr-33 V 33 -129.9822 45.9333 423850 5087101 18:03 1998 osmosampler recovered Chapin
   Mkr-33 V 33 -129.9822 45.9333 423850 5087101 Deployed Green & Copper OsmoSampler Chapin
   Mkr-33 V 33 -129.9822 45.9333 423850 5087101 Deployed 2 Fe OsmoAnalyzers Chapin
        Dive R484 SAMPLES
 Sample Vent Mkr Long Lat UTM X UTM Y Time Description Investigator Subsamples
SS_mat_R484-1 Mkr-33 V 33 -129.9822 45.9333 423850 5087101 1155 Bacterial mat - rock Moyer Feely/Huber/ Perfit
Bio_R484-2 Mkr-33 V 33 -129.9822 45.9333 423850 5087101 1234 Crab Juniper
Bactrp_R484-3 Cloud N4 -129.9817 45.9334 423896 5087119 1414 Bacterial trap #14, deployed in 1998 Moyer
MTR_R484-4 Cloud N4 -129.9817 45.9334 423896 5087119 1425 MTR #0942, deployed in 1998 Embley
Bactrp_R484-5 Cloud N4 -129.9817 45.9334 423896 5087119 1430 Bacterial trap #15, deployed in 1998 Moyer
TLC_R484-6 Mkr-33 V 33 -129.9822 45.9333 423850 5087101 1155 Time Lapse Camera, deployed in 1998 Tunnicliffe
   Mkr-33 V 33 -129.9822 45.9333 423850 5087101 Deployed bactraps # 40 - 43 Moyer
   Cloud N4 -129.9817 45.9334 423896 5087119 Deployed bactraps # 36 - 39 Moyer
       Dive R485 SAMPLES
   Sample Vent Mkr Long Lat UTM X UTM Y Time Description Investigator Subsamples
HFS2_dfl_R485-1 Mkr 33 33 -129.9822 45.9333 423850 5087101 11:17 Filtered diffuse flow for XRF -T1=55 T2=37 to 49 - 566ml Butterfield Gendron
HFS23_dfl_R485-2 Mkr 33 33 -129.9822 45.9333 423850 5087101 11:23 Gas piston of diffuse flow for gases - T1=53 T2=42-50 - 113ml Evans Butterfield/ Lilley
HFS12_dfl_R485-3 Mkr 33 33 -129.9822 45.9333 423850 5087101 11:26 Filtered diffuse flow for DNA - T1=49 T2=37 Tmax=54.6 -1223ml Huber
  HFS14_dfl_R485-4 Mkr 33 33 -129.9822 45.9333 423850 5087101 11:35 Bag with filter of diffuse flow. Tmax=48.8 1110ml Butterfield Graham
HFS17_dfl_R485-5 Mkr 33 33 -129.9822 45.9333 423850 5087101 11:43 Bag sample without filter. Tmax=48.8 Butterfield Huber/ Graham/ Levesque
HFS_dfl_R485-6 Mkr 33 33 -129.9822 45.9333 423850 5087101 Sample interrupted. Dive aborted
       Dive R487 SAMPLES
   Sample Vent Mkr Long Lat UTM X UTM Y Time Description Investigator Subsamples
HFS-12_plume

_R487-1

Mkr 33 33 -129.9822 45.9333 423850 5087101 17:18 Depth 1476m - Plume particles visible - DNA - 1502 ml Butterfield/ Huber  
HFS-15_plume

_R487-2

Mkr 33 33 -129.9822 45.9333 423850 5087101 17:34 Depth 1476m - Plume particles visible - Lipids -1050 ml Butterfield/ Huber
  HFS-17_plume

_R487-3

Mkr 33 33 -129.9822 45.9333 423850 5087101 17:45 Bag sample - no filter - plume particles visible - mirco & chem - 650 ml Butterfield /Huber/
  HFS-19_plume

_R487-4

Mkr 33 33 -129.9822 45.9333 423850 5087101 17:52 Bag sample-with filter -plume particles visible - mirco & chem - 675 ml Butterfield/ Graham
  HFS-16_plume

_R487-5

Mkr 33 33 -129.9822 45.9333 423850 5087101 17:59 Bag sample without filter - 598 ml Butterfield/ Huber/
  HFS-3_plume

_R487-6

Mkr 33 33 -129.9822 45.9333 423850 5087101 18:07 Filter for Sulfur - 1300 ml Butterfield
  GTB_plume

_R487-7

Cloud   -129.9816 45.9334 423904 5087109 18:47 port gastight #6 - 160 ml Evans Butterfield/ Lilley
HFS-23_plume

_R487-8

Cloud   -129.9816 45.9334 423904 5087109 18:49 gas piston - 170 ml Evans Butterfield/ Lilley
HFS-20_plume

_R487-9

Cloud   -129.9816 45.9334 423904 5087109 18:53 piston - 730 ml Butterfield/ Huber/
  HFS-21_plume

_R487-10

Cloud   -129.9816 45.9334 423904 5087109 19:01 filter sulfur - 730 ml Butterfield/ Huber
  HFS-22_plume

_R487-11

Cloud   -129.9816 45.9334 423904 5087109 19:11 piston - 730ml Butterfield/ Huber/
  HFS-18_plume

_R487-12

Cloud   -129.9816 45.9334 423904 5087109 19:18 Bag sample with filter - 600ml Butterfield
  HFS-7_plume

_R487-13

Cloud   -129.9816 45.9334 423904 5087109 19:25 Filter for XRF - 411 ml Butterfield/ Gendron
  HFS-2_plume

_R487-14

Cloud   -129.9816 45.9334 423904 5087109 19:29 Filter for XRF - 423ml Butterfield/ Gendron
  HFS-11_plume

_R487-15

Cloud   -129.9816 45.9334 423904 5087109 19:37 Bag with filter - integrated sample - stop at 1338 ml Gendron/ Graham
       Dive R488 SAMPLES
   Sample Vent Mkr Long Lat UTM X UTM Y Time Description Investigator Subsamples
HFS-14_dfl_R488-1 Mkr 33 33 -129.9822 45.9333 423850 5087101 6:28 Bag with filter T=65-70 Tmax=70.4. 655 ml Butterfield Graham
HFS-6_dfl_R488-2 Mkr 33 33 -129.9822 45.9333 423850 5087101 6:35 Filter#6 - Tmax=73.2 - 1 liter Huber
  HFS-15_dfl_R488-3 Mkr 33 33 -129.9822 45.9333 423850 5087101 6:50 Lipid filter#15 at 0.45um - Tmax=78 1.5 liters Huber
  GTB-7_dfl_R488-4 Mkr 33 33 -129.9822 45.9333 423850 5087101 7:06 Stbd gas Tight - T=68.8 Evans Butterfield/ Lilley
SS-J1_bio_R488-5 Mkr33 33 -129.9822 45.9333 423850 5087101 7:15 Palm worms - rock Juniper/ Levesque Perfit/ Marcus
SS-3_mat_R488-6 Mkr33 33 -129.9822 45.9333 423850 5087101 7:53 Jar#3 - bacterial mat in the crack - rock Moyer Perfit/ Marcus/
SS-J4_bio_R488-7 Snail N8 -129.9819 45.9332 423879 5087090 8:39 Jar#4 - snails - Tmax=13 Marcus Levesque
HFS-7_dfl_R488-8 Cloud N4 -129.9816 45.9334 423901 5087111 9:44 Bag#17 no filter - Tmax=18.7 700 ml - DIDN'T WORK Butterfield/ Huber/
  HFS-22_dfl_R488-9 Cloud N6 -129.9816 45.9334 423901 5087111 10:19 Piston#22 - Tmax=20.3 615 ml Butterfield/ Huber/ Levesque
HFS-24_dfl

_R488-10

Cloud N6 -129.9816 45.9334 423901 5087111 10:25 Gas piston#24 - Tmax=20.3 130 ml Evans Butterfield/ Lilley
HFS-3_dfl

_R488-11

Cloud N6 -129.9816 45.9334 423901 5087111 10:28 Bottle#3 filter - Tmax=20.2 672 ml Gendron
  GTB-2_dfl

_R488-12

Cloud N6 -129.9816 45.9334 423901 5087111 10:39 Port gas tight - T=20.3 Evans Butterfield/ Lilley
HFS-7_dfl

_R488-13

Cloud N6 -129.9816 45.9334 423901 5087111 10:42 Filter#7 for sulfur - Tmax=20.4 Butterfield
  SS-J3_dfl

_R488-14

Cloud N4 -129.9818 45.9334 423888 5087111 11:22 Jar#3 - fauna - snails mat and scale worms Levesque/Marcus
  SS-1_dfl

_R488-15

Cloud N4 -129.9818 45.9334 423888 5087111 11:32 Jar#1 - bacterial mat and rock Moyer Perfit/ Marcus
SS-4_dfl

_R488-16

Cloud N4 -129.9818 45.9334 423888 5087111 12:25 Flushing bottle - worms and other fauna Marcus/

Levesque/ Tsurumi

Perfit
SS-2_dfl_

R488-17

Cloud N6 -129.9816 45.9334 423901 5087111 12:48 Jar#2 - particulates Levesque
  HFS-11_dfl

_R488-18

Nascent   -129.9815 45.9358 423911 5087378 13:55 Bag with filter - at MTR 4108. Tave=10-11 Tmax=15.5 Butterfield
  HFS-18_dfl

_R488-19

MiniSnow N9 -129.9842 45.9426 423710 5088141 15:43 Bottle#18 - varying temp Tmax=4 720 ml Butterfield
  HFS-20_dfl

_R488-20

n/a N3 -129.9852 45.9438 423637 5088278 16:16 Piston#20 - Tmax=16 650 ml Butterfield/ Huber/
  HFS-2_dfl

_R488-21

n/a N3 -129.9852 45.9438 423637 5088278 16:26 Filter#2 for XRF - Tmax=17.3 800 ml Gendron
  HFS-5_dfl

_R488-22

n/a N3 -129.9852 45.9438 423637 5088278 16:37 Gas Piston #5 - Tmax=16.5 112 ml Evans Butterfield/ Lilley
HFS-19_dfl

_R488-23

Milky N2 -129.9847 45.9451 423679 5088420 17:36 Bag#19 with filter Tmax=3.4 700 ml Butterfield/ Graham
  HFS-16_dfl

_R488-24

Ouzo   -129.9847 45.9458 423680 5088497 18:37 Bag without filter - Tmax=3.5 in sample T=3.8 degrees before sample Butterfield
  HFS-23_dfl

_R488-25

Magnesia   -129.9849 45.9461 423665 5088526 19:52 Piston#23 - Tave=4 Tmax=5.6 full piston, oops Butterfield/

Evans/

  HFS-12_dfl

_R488-26

Magnesia   -129.9849 45.9461 423665 5088526 20:02 DNA filter set, Tmax=5.1 1020 ml Huber
  HFS-9_dfl

_R488-27

Magnesia   -129.9849 45.9461 423665 5088526 20:12 Bag#9 with POC filter Tmax=4.3 681 ml Butterfield
  HFS-1_dfl

_R488-28

Magnesia   -129.9849 45.9461 423665 5088526 20:19 Filter#1 for FISH - Tmax=5.1 1005 ml Huber
  HFS-10_dfl

_R488-29

Magnesia   -129.9849 45.9461 423665 5088526 20:30 Filter for lsulfur - T=4.8 820 ml Butterfield
  HFS-21_dfl

_R488-30

Magnesia   -129.9849 45.9461 423665 5088526 Filter#21 0.45 micron lipids filter - Tmax=5.0 1542 ml Huber
  HFS-4_dfl

_R488-31

Magnesia   -129.9849 45.9461 423665 5088526 20:56 Piston#4 - Tmax=4 Butterfield/

Huber

  HFS-13_dfl

_R488-32

Magnesia   -129.9849 45.9461 423665 5088526 21:04 DNA filter set#13 - Tmax=4.9 1789 ml Huber
  SS-j2_dfl

_R488-33

Magnesia   -129.9849 45.9461 423665 5088526 21:26 Jar#2 - floc Juniper/ Levesque
       Dive R491 SAMPLES
   Sample Vent Mkr Long Lat UTM X UTM Y Time Description Investigator Subsamples
Osmo_R491-1 Mkr-33 33 -129.9822 45.9333 423850 5087101 4:43 FeOsmoAnalyzer#1 to elevator - deployed Dive R483 Chapin
  Bactrp-40_R491-2 Mkr-33 33 -129.9822 45.9333 423850 5087101 4:54 Deployed on Dive R484 Moyer
  Bactrp-43_R491-3 Mkr-33 33 -129.9822 45.9333 423850 5087101 4:55 Deployed Dive R484 Moyer
  Osmo-R491-4 Mkr-33 33 -129.9822 45.9333 423850 5087101 5:19 FeOsmoAnalyzer#2 to elevator - deployed Dive R483 Chapin
  SS-j1_bio_R491-5 Mkr-33 33 -129.9822 45.9333 423850 5087101 6:01 3 meters from Mkr-33 crack - limpets snails and mat Levesque Marcus
GTB-6_dfl_R491-6 Cloud N6 -129.9816 45.9334 423901 5087111 6:56 Gastight in hole - very vigorous flow - previously measured T=20.3 Evans Butterfield/ Lilley
SS-1_mat_R491-07 Coud N6 -129.9816 45.9334 423901 5087111 7:17 microbial mat Moyer
  Bactrp-36_R491-08 Cloud N4 -129.9818 45.9334 423888 5087111 7:42 Deployed Dive R484 Moyer
  Bactrp-37_R491-09 Cloud N4 -129.9818 45.9334 423888 5087111 7:48 Deployed Dive R484 Moyer
  Bactrp-39_R491-10 Cloud N4 -129.9818 45.9334 423888 5087111 7:51 Deployed Dive R484 Moyer
  Bactrp-38_R491-11 Cloud N4 -129.9818 45.9334 423888 5087111 7:54 Deployed Dive R484 Moyer
  GTB-7_dfl_R491-12 Nascent   -129.9815 45.9358 423911 5087378 10:08 Gastight fired near MTR Evans Butterfield/ Lilley
SS-j3_bio_R491-13 Nascent   -129.9815 45.9358 423911 5087378 10:13 Over tube worm bush - gastropods and other little animals Tunnicliffe
  SS-j4_bio_R491-14 Nascent   -129.9815 45.9358 423911 5087378 10:48 Gastropods and other small animals near MTR Juniper/

Levesque

Marcus
MTR-4180

_R491-15

Nascent   -129.9815 45.9358 423911 5087378 11:12 MTR deployed 1998 Dive R478 Embley
  Bio_R491-16 Nascent   -129.9815 45.9358 423911 5087378 11:16 Four tube worm grabs in diffuse flow Levesque/

Marcus

  MTR-4126

_dfl_R491-17

  N41 -129.9814 45.9362 423922 5087428 12:46 Deployed 1998 Dive R474 Embley
  Bio_R491-18   N41 -129.9814 45.9362 423922 5087428 12:55 Tube worm grab in diffuse flow Marcus
  SS-j2_FeO

_R491-19a

not at vent   -129.9816 45.9344 423899 5087414 13:41 FeO sample + white tentacled bacteria Scott
  SS-j2-FeO

_R491-19b

not at vent   -129.9816 45.9344 423899 5087414 14:00 FeO and white tentacled bacterial mat - switched from jar2 to jar4 - wasn't Scott
  Bio_R491-20 Old Flow   -129.9817 45.9364 423898 5087455 14:34 Tube worm grab in diffuse flow - worms look dead or at least not very Tsurumi
  Rck_R491-21 Old Flow   -129.9817 45.9364 423898 5087455 15:15 Shell of hollow lobate rock with about 10cm crust - in area of diffuse flow on Perfit Scott
Rck_R491-22 not at vent   -129.9816 45.9344 423902 5087224 16:11 New lava flow sample from roof structure about 100m away from Perfit
  Bio_R491-23 Old Worms   -129.9833 45.9451 423785 5088418 5:10 Tube worm grab in diffuse flow on old lava - also sampled small rock Levesque
  Bio_R491-24 Old Worms   -129.9833 45.9451 423785 5088418 5:30 Tube worm grab in diffuse flow on old lava Tsurumi
  Rck_R491-25 Old Worms   -129.9833 45.9451 423785 5088418 5:40 Rock sample in diffuse flow on sedimented old lava - a couple of rocks Perfit
  SS-2_mat

_R491-26

Magnesia   -129.9849 45.9462 423661 5088545 8:18 White and orange mat (bacterial and sulfur) Moyer Huber
SS-3_mat

_R491-27

Magnesia   -129.9849 45.9462 423661 5088545 8:31 White and orange mat (bacterial and sulfur) Moyer
  SS-4_FeO

_R491-28

n/a   -129.9816 45.9344 423902 5087224 9:21 Sampled earlier at Oxide - in flushing bottle Scott
   Cloud N6 -129.9816 45.9334 423901 5087111 Deployed VEMCO 1108 & MTR3157 (taped together) Embley
   Nascent   -129.9815 45.9358 423911 5087378 12:25 Deployed MTR 3175 Embley
   n/a N41 -129.9814 45.9362 423922 5087428 13:03 Deployed MTR 3041 Embley
   Mkr-33 V 33 -129.9822 45.9333 423850 5087101   Deployed BioOsmoSampler Chapin
       Dive R492 SAMPLES
   Sample Vent Mkr Long Lat UTM X UTM Y Time Description Investigator Subsamples
SS-j1_bio_R492-1 Joystick 42 -129.9886 45.9188 423342 5085505 1351 Tube worms - bag creatures - nemartean? - and other fauna on a rock Marcus
  Rck_492-2 Joystick 42 -129.9886 45.9188 423342 5085505 14:50 Orange oxidized rock - few meters southwest of Mkr-42 - rock corer Perfit
  SS-6_mat_R492-3 Joystick 42 -129.9886 45.9188 423342 5085505 15:04 Bacterial mat and bag creatures Moyer
  SS-8_mat_R492-4 Joystick 42 -129.9886 45.9188 423342 5085505 15:24 Suction of white filamentous material off rock Moyer
  SS-7_bio_R492-5 Coquilles   -129.9931 45.9175 422991 5085365 16:48 Clams - holothurians - background fauna Levesque
  Bio_R492-6 Coquilles   -129.9931 45.9175 422991 5085365 17:28 Tube worm bush containing lots of other organisms Tsurumi Levesque
SS-j4_bio_R492-7 Coquilles   -129.9931 45.9175 422991 5085365 17:46 Sample in middle of tube worm bush for small animals Levesque
  Rck_R492-8 near Dying   -129.9919 45.9168 423080 5085277 18:36 New lobate flow sample at contact - 7 function arm Perfit
  Rck_R492-9 near Dying   -129.9919 45.9168 423080 5085277 18:43 Old flow sample at same contact point - rock corer Perfit
  Bio_R492-10 Bag City   -129.9816 45.9344 423284 5085214 19:49 Tube worm grab Marcus McHugh
GTB-2_dfl

_R492-11

Bag City   -129.9816 45.9344 423284 5085214 20:13 Gastight at site where tube worms were sampled Evans Butterfield/ Lilley
SS-j3_dfl_R492-12 Bag City   -129.9816 45.9344 423284 5085214 20:17 Suction diffuse flow - same site - T=22 Butterfield Huber
SS-4_FeO_R492-13 Bag City   -129.9816 45.9344 423284 5085214 20:30 Iron oxides (red) sample in flush bottle Scott
  Rck_R492-14 near Bag City   -129.9893 45.9162 423281 5085209 21:28 Lava rubble on old flow in area of old and new lava - over 200m southeast of Perfit
      Dive R493 SAMPLES
   Sample Vent Mkr Long Lat UTM X UTM Y Time Description Investigator Subsamp
Rck_R493-1 n/a n/a -129.9944 45.9134 422880 5084908 7:38 Few pieces of old lava at contact - rock corer -ContactW5 Perfit
  Rck_R493-2 n/a n/a -129.9942 45.9135 422899 5084917 7:57 piece of drip structure - new lava flow - 7 function arm Perfit
      Dive R494 SAMPLES (position at GPS antennae on ship - no acoustic
   Sample Vent Mkr Long Lat UTM X UTM Y Time Description Investigator Subsamples
Rck_R494-1 n/a n/a -130.0042 45.8583 422042 5078796 4:30 Old folded sheets - depth 1723 - rock corer (purple) Perfit
  SS-j2_R494-2 n/a n/a -130.0042 45.8583 422042 5078796 5:00 Thin yellowish pelagic or hydrothermal sediment on top of lobate - brittle stars - Juniper
  Rck_R494-3 n/a n/a -130.0037 45.8621 422087 5079221 5:50 Piece of thick sheet flow slab at fissure - about 10cm wide - reddish staining Perfit
  Rck_R494-4 n/a n/a -130.0041 45.8624 422056 5079244 6:15 Rock from bottom of fissure - glassy new flow - end of lava tube - 7 function Perfit
  Rck_R494-5 n/a n/a -130.0041 45.8624 422056 5079244 6:25 Second rock sample from same spot - glassy new flow - 7 function arm Embley Perfit
Rck_R494-6 n/a n/a -130.0041 45.8624 422056 5079244 6:30 Third rock - old slab - only got small piece of glassy top of sheet flow - depth Perfit
  Bio_R494-7 n/a n/a -129.9992 45.8772 422455 5080891 10:51 Crab leg - depth 1702m Juniper
  Rck_R494-8 n/a n/a -129.9958 45.8782 422727 5080994 11:29 Small piece of new flow at east contact Perfit
   Rck_R494-9 n/a n/a -129.9929 45.8894 422965 5082238 14:20 Sample from jumbled new lava - rock corer (black stripes) Perfit
      Dive R495 SAMPLES
   Sample Vent Mkr Long Lat UTM X UTM Y Time Description Investigator Sub-samps
HFS-17_dfl_R495-1   108b -129.9832 45.9288 423772 5086608 8:09 Unfiltered bag #17 - diffuse venting at base of pillar - Tmax=13.6 - 500 ml Butterfield/ Huber/
  HFS-5_dfl_R495-2   108b -129.9832 45.9288 423772 5086608 8:16 Gas piston - Tave=13.2 - 125 ml Evans Butterfield/ Lilley
HFS-10-dfl-R495-3   108b -129.9832 45.9288 423772 5086608 8:18 filtered bag for xrf - 0.45 um - Tmax=13.3 - 600 ml Gendron
  SS-j1_sed_R495-4 near Mkr-108   -129.9834 45.9283 423759 5086555 8:50 Sediment sample of diatom mat near Mkr 108 Tunnicliffe
  SS-j4_sed_R495-5 near Mkr-108   -129.9834 45.9283 423759 5086555 8:58 Sediment sample of diatom mat near Mkr 108 Juniper
  HFS-4_htfl

_R495-6

Castle   -129.9799 45.9261 424023 5086306 10:16 Piston - T1=274 - T2=175 - Tmax=275 - 460 ml Butterfield/ Huber/
  HFS-2_htfl

_R495-7

Castle   -129.9799 45.9261 424023 5086306 10:22 Filtered bag for xrf - 0.45 um - Tmax=255 - 507 ml Gendron
  HFS-18_htfl

_R495-8

Castle   -129.9799 45.9261 424023 5086306 10:31 Filtered bag -0.45 um - Tmax=262 - about 375 ml Butterfield/ Graham
  GTB_htfl_R495-9 Castle   -129.9799 45.9261 424023 5086306 10:40 Gastight - Tmax=245 Evans Butterfield/ Lilley
Niskin_htfl

_R495-10

Castle   -129.9799 45.9261 424023 5086306 10:52 Niskin bottle for xrf Gendron
  SF_R495-11 Castle   -129.9799 45.9261 424023 5086306 11:00 Sulfide chimney spire placed in purse - SAMPLE LOST (purse tore) Scott
  FeO_R495-12 FlatTop N5 -129.9796 45.9261 424043 5086306 11:26 FeO from top of Flattop placed in purse - SAMPLE LOST (purse tore) Scott
  HFS-11_dfl

_R495-13

Mkr-113 113 -129.9883 45.9226 423370 5085922 13:03 Filter bag 11 - T2=11.8 Tmax=22,6 Butterfield/ Graham
  GTB_dfl_R495-14 Mkr-113 113 -129.9883 45.9226 423370 5085922 13:13 Gastight Evans Butterfield/ Lilley
HFS-16_dfl

_R495-15

Mkr-113 113 -129.9883 45.9226 423370 5085922 13:15 Unfiltered bag - T=27 Butterfield/ Huber
  SS-01_mat

_R495-16

Mkr-113 113 -129.9883 45.9226 423370 5085922 13:29 Sucking mat and bag creatures? Moyer
  SS-02_mat_

R495-17

Mkr-113 113 -129.9883 45.9226 423370 5085922 13:29 Sucking mat bag creatures? Hose broke on sampler Moyer
  VEMCO&bio

_R495-18

Mkr-113 113 -129.9883 45.9226 423370 5085922 6:52 Recovered VEMCO-MTR and tube worms attached Embley/ Marcus
  HFS-22_dfl

_R495-19

Joystick   -129.9886 45.9188 423342 5085505 15:03 Piston - Tmax=9.6 - 630 ml Huber/ Butterfield/
  HFS-14_dfl

_R495-20

Joystick   -129.9886 45.9188 423342 5085505 15:11 Bag with filter .45um - collecting sulfur - Tmax=8.7 degrees - 643 ml Butterfield/ Graham
  HFS-12_dfl

_R495-21

Joystick   -129.9886 45.9188 423342 5085505 15:18 3 micron/sterivex filter for DNA - T=8.5 - 1500 ml Huber
  HFS-6_dfl

_R495-22

Joystick   -129.9886 45.9188 423342 5085505 15:35 0.2 micron filter/FISH - Tmax=9 - 1000 ml Huber
  HFS-24_dfl

_R495-23

Joystick   -129.9886 45.9188 423342 5085505 15:46 Gas piston - Tmax=9.9 - 100 ml Evans Butterfield/ Lilley
HFS-15_dfl

_R495-24

Joystick   -129.9886 45.9188 423342 5085505 15:49 0.45 micron filter for lipids - Tmax=9.8 - 820 ml Huber
  HFS-3_dfl

_R495-25

Joystick   -129.9886 45.9188 423342 5085505 16:01 0.45 micron filter of XRF - Tmax=10.1 - 570 ml Gendron
  HFS-19_dfl

_R495-26

Coquilles   -129.9931 45.9175 422991 5085365 17:12 Bag with filter 0.45 microns - Tmax=20.7 - 650 ml Butterfield/ Graham
  HFS-20-_dfl

_R495-27

Bag City 36 -129.9816 45.9344 423284 5085214 18:17 Piston - T1=22 Tave=22.6 Tmax=23.2 - 683 ml Butterfield/ Graham/
  HFS-23_dfl

_R495-28

Bag City 36 -129.9816 45.9344 423284 5085214 18:23 Gas piston - Tmax=23.4 - 128 ml Evans Butterfield/ Lilley
HFS-21_dfl

_R495-29

Bag City 36 -129.9816 45.9344 423284 5085214 18:28 0.45 micron filter for XRF - Tmax=23.3 - 500 ml Gendron
  HFS-13_dfl

_R495-30

Bag City 36 -129.9816 45.9344 423284 5085214 18:30 3 micon/sterivex filter for DNA - Tmax=23.4 - 1000 ml Huber
  HFS-9_dfl

_R495-31

Bag City 36 -129.9816 45.9344 423284 5085214 18:40 Bag with POC filter - Tmax=23.5 - 533 ml Butterfield/ Gendron
  HFS-7_dfl

_R495-32

Bag City 36 -129.9816 45.9344 423284 5085214 18:45 0.45 micron filter for lipids - Tmax=23.2 - 1000 ml Huber
  HFS-1_dfl

_R495-33

Bag City 36 -129.9816 45.9344 423284 5085214 18:56 0.2 um FISH filter - Tmax=23.4 - 1000 ml Huber
  Rck_R495-34 Crevice   -129.9904 45.9111 423175 5084648 21:35 New flow lobate lava from little collapse - SAMPLE LOST (purse tore) Perfit
  SS-j3_FeO

_R495-35

Crevice   -129.9904 45.9111 423175 5084648 22:00 Sediment which is partly composed of some sort of iron oxide - probably basalt Scott Perfit
SS-j2_bio

_R495-36

Crevice   -129.9904 45.9111 423175 5084648 22:16 Small patch of dead tube worms shaken for any other small organisms - FeO Tsurumi Perfit
SS-04_fl_R495-37 n/a   -129.9913 45.9052 423107 5083994   Background water collected while ROPOS coming up - depth 1300m - Juniper/ Levesque/
   BagCity   -129.9816 45.9344 423284 5085214 19:35 Deployed MTR Embley
   Mkr-33 V 33 -129.9822 45.9333 423850 5087101   Deployed OsmoAnalyzer Chapin
      Dive R496 SAMPLES
   Sample Vent Mkr Long Lat UTM X UTM Y Time Description Investigator Subsamps
Bactrp-20_R496-1 Mkr-113 113 -129.9883 45.9226 423370 5085922 7:59 Bacteria trap #20 recovered (deployed in 98 - Dive R464) Moyer
  Bactrp-21_R496-2 Mkr-113 113 -129.9883 45.9226 423370 5085922 7:59 Bacteria Trap #21 recovered (deployed in 98 - Dive R464) Moyer
  Bio_R496-3 Mkr-113 113 -129.9883 45.9226 423370 5085922 8:08 Tube worm grab - (total of 3 grabs) Marcus Perfit
Rumble_R496-4 n/a   -129.9840 45.9302 423711 5086763 16:05 Recovered rumbleometer stuck in the 98 lava flow - Perfit grabbed rock chips Fox Perfit
      Dive R497 SAMPLES (No acoustic nav at CASM - positions at ship GPS antennae)
   Sample Vent Mkr Long Lat UTM X UTM Y Time Description Investigator Subsamp
SS-j3_dfl_R497-1 n/a n/a -130.0264 45.9890 420506 5093338 7:44 Water sample over palm worms, about 15 m east of fissure - log file 2 - Butterfield/ Huber/
  Sf_R497-2 T&S n/a -130.0272 45.9890 420447 5093343 9:32 Little piece of sulfide spire Scott Levesque
Sf&Bio_R497-3 T&S n/a -130.0272 45.9890 420447 5093343 9:39 Top of sulfide chimney and crab Scott Juniper
SS-j4_dfl_R497-4 T&S n/a -130.0272 45.9889 420446 5093322 10:09 Water sample - log file 3 at Vemco and Maya's tube worm sample site - Butterfield/ Huber/
  GTB _R497-5 T&S n/a -130.0272 45.9889 420446 5093322 10:22 Gastight (portside) Evans
  Sf_R497-6 T&S n/a -130.0272 45.9889 420446 5093322 10:25 Sulfide sample where port gas tight was taken (broke up a lot) Scott
  VEMCO_R497-7 T&S n/a -130.0272 45.9889 420446 5093322 10:32 VEMCO 98-1113-214 placed in stbd side biobox Embley
  SS-j1_bio_R497-8 T&S n/a -130.0272 45.9889 420446 5093322 10:40 Particles over tube worms - log file 4;- resumed at 10:58 inside tube worm Juniper/ Levesque
  Bio_R497-9 T&S n/a -130.0272 45.9889 420446 5093322 10:50 Tube worm grab Tsurumi
  SS-1_mat_R497-10 T&S n/a -130.0272 45.9889 420446 5093322 11:22 Suction of mat - Tmax=6.1 Moyer
  GTB_R497-11 T&S n/a -130.0272 45.9889 420446 5093322 11:44 Gas tight bottle near top of chimney - Tspike=28 Evans
  Sf_R497-12 T&S n/a -130.0272 45.9889 420446 5093322 11:48 Sulfide sample - lots of tube worms attached Scott/ Perfit/
  SS-2_dfl_R497-13 T&S n/a -130.0272 45.9889 420446 5093322 12:13 Fluid sample over lush tubes worms - Tave=26 Tmax=38,5 Butterfield/ Huber/
  SS-j2_bio_R497-14 T&S n/a -130.0272 45.9889 420446 5093322 12:30 100 micron mesh over lush tube worms for larvae (suspended particles) - 16 Levesque/ Juniper
  SS-3_bio_R497-15 T&S n/a -130.0272 45.9889 420446 5093322 12:53 Small animals and other material from within community of lush tube worms Levesque/ Juniper/
  Hobo_R497-16 T&S n/a -130.0272 45.9889 420446 5093322 13:12 Recovery of hobo deployed in 98 on DiveR480 Embley
  Niskin_dfl_R497-17 T&S n/a -130.0272 45.9889 420446 5093322 13:36 Niskin about 6 inches from top of spire - diffuse flow Gendron
  Sf_R497-18 T&S ? n/a -130.0272 45.9889 420446 5093322 ? Decimeter piece of sulfide chimney caught in ROPOS - found on recovery. Scott
  Sf_R497-19 T&S n/a -130.0272 45.9889 420446 5093322   Sulfide (chalcopyrite with minor ?anhydrite) scraped off HOBO probe -
   Rck_R497-20 NRZ n/a -130.0219 46.0198    18:03 Lobate flow top from NRZ in area of extensometers Perfit
  Sf_dfl_R497-21 T&S n/a -130.0272 45.9889 420446 5093322   Scraping from Vemco near tubeworm grab R497-10 - appears to be mostly Scott
   T&S n/a -130.0272 45.9889 420446 5093322   Deployed MTR 3017 Embley
      Dive R498 SAMPLES
   Sample Vent Mkr Long Lat UTM X UTM Y Time Description Investigator
  Exten_R498-1    -130.0180 46.0194 421203 5096708   Recover '98 extensometer E1 Chadwick
  Exten_R498-2    -130.0203 46.0110 421026 5096739   Recover '98 extensometer E2 Chadwick
  Exten_R498-3    -130.0212 46.0197 420956 5096745   Recover '98 extensometer E3 Chadwick
  Exten_R498-4    -130.0241 46.0202 420732 5096799   Recover '98 extensometer E4 Chadwick
      Dive R499 SAMPLES
   Sample Vent Mkr Long Lat UTM X UTM Y Time Description Investigator
  HFS-3_dfl_R499-1 Magnesia   -129.9849 45.9462 423661 5088545 17:08 Filter #3 for XRF - Tmax=7.3 - 500 ml Gendron
  HFS-4_dfl_R499-2 Magnesia   -129.9849 45.9462 423661 5088545 17:13 Piston #4 - Tmax=7.4 - 600 ml Butterfield/ Huber/ Graham
  HFS-5_dfl_R499-3 Magnesia   -129.9849 45.9462 423661 5088545 17:20 Gas piston #5 - Tmax=7.1 - 125 ml Evans
  HFS-18_dfl_R499-4 Magnesia   -129.9849 45.9462 423661 5088545 17:26 Bag #18 - Tmax=7.3 - 613 ml Butterfield/ Graham
  niskin_dfl_R499-5 Magnesia   -129.9849 45.9462 423661 5088545 18:18 Niskin diffuse flow Gendron
  HFS-24-dfl-R499-6 Old Worms   -129.9833 45.9451 423785 5088418 20:20 Gas piston #24 - area of lush growth and good flow - Tmax=7.8 - about 150 ml Evans
   Magnesia   -129.9849 45.9462 423661 5088545 Deployed OsmoSampler Chapin
      Dive R500 SAMPLES
   Sample Vent Mkr Long Lat UTM X UTM Y Time Description Investigator Subsamp
HFS-99_dfl_R500-1 Old Worms   -129.9833 45.9451 423785 5088418 0:47 Bag #9 with POC filter - isothermal flow under tube worms - T1=9.5 T2=7.5 Butterfield
  HFS-17_dfl_R500-2 Old Worms   -129.9833 45.9451 423785 5088418 0:59 Bag #17 unfiltered - no valve - Tmax=9.8 - 700 ml Butterfield/ Huber/
  HFS-23_dfl_R500-3 Old Worms   -129.9833 45.9451 423785 5088418 1:06 Gas piston #23 - Tmax=9.8 Evans Butterfield/ Lilley
SS-j3_dfl_R500-4 Old Worms   -129.9833 45.9451 423785 5088418 1:28 64 micron mesh - Larval collection over old tube worm bush Tunnicliffe
  SS-j1_bio_R500-5 Old Worms   -129.9833 45.9451 423785 5088418 1:50 100 micron mesh - Small animal and biofilm sample in old tube worm bush Levesque/ Charpentier
  SS-j2_bio_R500-6 Old Worms   -129.9833 45.9451 423785 5088418 2:20 Clams and other fauna Levesque/ Charpentier Perfit
HFS-19_dfl_R500-7 NE of FeO   -129.9816 45.9362 423904 5087423 5:48 Bag #19 with 0.45 um XRF filter (will analyze small particles on filter witth Butterfield/ Gendron/
  GTB-6_dfl_R500-8 NE of FeO   -129.9816 45.9362 423904 5087423 6:07 Port gastight #6 - Tmax=5 Evans Butterfield/ Lilley
HFS-16_dfl_R500-9 15m S of N41   -129.9815 45.9359 423911 5087391 6:22 Bag #16 unfilterted - no valve - Tmax=25.9 - 600 ml Butterfield/ Huber/
  GTB-7_dfl

_R500-10

15m S of N41   -129.9815 45.9359 423911 5087391 6:30 Stbd gastight #7 - Tmax=26 Evans Butterfield/ Lilley
HFS-11_dfl

_R500-11

15m S of N41   -129.9815 45.9359 423911 5087391 6:43 Bag #11 with 0.45 um filter - no valve - Tmax=22.7 Texit=14 - 620 ml Butterfield
  HFS-10_dfl

_R500-12

15m S of N41   -129.9815 45.9359 423911 5087391 6:50 Filter #10 for lipids - Tmax=24.4 - 1 liter Huber
  HFS-13_dfl

_R500-13

15m S of N41   -129.9815 45.9359 423911 5087391 7:00 Filter #13 - 3 um and sterivex for DNA - Tmax=24.2 - 1200ml Huber
  HFS-4_dfl

_R500-14

n/a   -129.9814 45.9352 423918 5087310 8:05 Piston #4 - Tmax=11.3 - 600 ml Butterfield/ Huber/
  HFS-6_dfl

_R500-15

Cloud N6 -129.9816 45.9334 423901 5087116 9:26 Filter #6 - 0.2 um for FISH - Tmax=19.9 - 1 liter Huber
  HFS-12_dfl

_R500-16

Cloud N6 -129.9816 45.9334 423901 5087116 9:36 Filter #12 - 3 um sterivex for DNA, - Tmax=20 - 1005 ml Huber
  HFS-15_dfl

_R500-17

Cloud N6 -129.9816 45.9334 423901 5087116 9:47 Filter #15 for lipids - Tmax=19.9 - 1002 ml Huber
  HFS-20_dfl

_R500-18

Cloud N6 -129.9816 45.9334 423901 5087116 9:58 Piston #20 - Tmax=19.9 - 675 ml Butterfield/ Huber/
  Niskin_dfl

_R500-19

Cloud N6 -129.9816 45.9334 423901 5087116 10:22 Niskin bottle over the hole at Mkr-N6 Gendron
  HFS-1_dfl

_R500-20

MK33 33 -129.9822 45.9333 423850 5087101 12:24 Filter #1 for FISH - Tmax=75.5 T2=43 - 1 liter Huber
  HFS-2_dfl

_R500-21

MK33 33 -129.9822 45.9333 423850 5087101 12:34 Filter #2 for XRF - Tmax=77.5 - 500 ml Gendron
  HFS-7_dfl

_R500-22

MK33 33 -129.9822 45.9333 423850 5087101 12:40 Filter #7 - sterivex for DNA - Tmax=75.7 T1=66 T2=43 Moyer
  HFS-21_dfl

_R500-23

MK33 33 -129.9822 45.9333 423850 5087101 12:58 Filter #21 for lipids - Tmax=73.2 - 1200 ml Huber
  HFS-22_dfl

_R500-24

MK33 33 -129.9822 45.9333 423850 5087101 13:09 piston #22 for water chemistry - Tmax=74.6 - 623 ml Butterfield/ Huber/
  HFS-14_dfl

_R500-25

MK33 33 -129.9822 45.9333 423850 5087101 13:16 Bag #14 with filter - Tmax 73.5 - 525 ml Butterfield/ Graham
      Dive R501 SAMPLES
   Sample Vent Mkr Long Lat UTM X UTM Y Time Description Investigator Subsamp
SS-1_mat_R501-1 Mkr 33 33 -129.9822 45.9333 423850 5087101 23:40 Bacterial mat in venting crack - piece of basalt that jammed in carousel Moyer/ Lynch Scott
SS-j3_rck_R501-2 n/a   -129.9924 45.9206 423050 5085709 5:37 Slurp of glass from intermediate? age top of pillow Perfit Scott
SS-j7_rck_R501-3 n/a   -129.9914 45.9207 423126 5085715 5:55 Younger of two flows at contact - small tube/toe in lobates - couldn't do a grab Perfit
  Rck_R501-4 n/a   -129.9899 45.9203 423238 5085670 6:32 Intermediate flow? - near the collapse in some broken lobes - blocky chunk with Perfit
  Rck_R501-5 n/a   -129.9877 45.9204 423411 5085680 7:01 Sample from crust of hollowed lobe - stubby with thin glass and angular edges Perfit
  SS-j8_rck_R501-6 n/a   -129.9867 45.9206 423492 5085695 7:25 Glass from lineated sheet flow, intermediate age (?) Perfit
  Rck_R501-7 n/a   -129.9818 45.9204 423865 5085668 8:15 Slab of rock from shallow collapse jumbled area Perfit
  Rck_R501-8 n/a   -129.9777 45.9202 424183 5085645 9:48 Chunk of lava from old flow - small, glassy Perfit
  SS-4_rck_R501-9 n/a   -129.9777 45.9201 424188 5085637 10:08 Little pieces of basalt glass from newer flow Perfit
  Rck_R501-10 n/a   -129.9823 45.9240 423831 5086069 11:29 Big plates (2) of mat covered ropey sheets Perfit
  SS-j4_mat_R501-11 n/a   -129.9823 45.9239 423837 5086063 11:35 Mat on rocks - same spot as last rock sample Juniper
  Rck_R501-12 n/a   -129.9876 45.9240 423427 5086072 12:30 Lava from chaotic collapse area- older with sponges Perfit
  Rck_R501-13 n/a   -129.9886 45.9240 423348 5086075 12:45 Contact area - black glassy newer lava - black striped corer Perfit
  Rck_R501-14 n/a   -129.9886 45.9240 423348 5086075 13:05 Contact area - more sediment and brittle stars - yellow striped corer Perfit
  Niskin_R501-15 n/a   -129.9886 45.9240 423348 5086075 13:18 Niskin sample at contact Juniper
  Rck_R501-16 n/a   -129.9922 45.9240 423065 5086078 13:45 Older? lava with sponges and more sediments - purple core Perfit
  Bio_501-17 Crevice   -129.9905 45.9109 423177 5084629 15:57 Tube worm grab within hole near crevice -diffuse venting Tsurumi
   Mkr 33 33 -129.9822 45.9333 423850 5087101 Deployed Time Lapse Camera Tunnicliffe
      Dive R502 SAMPLES
   Sample Vent Mkr Long Lat UTM X UTM Y Time Description Investigator Subsamp
HFS-5_dlf_R502-1 Gollum   -130.0136 45.9336 421422 5087166 2:41 Gas piston #5 - max=22.5 Evans Butterfield/ Lilley
HFS-6_dlf_R502-2 Gollum   -130.0136 45.9336 421422 5087166 3:16 0.2um filter for FISH - T1=9 T2=16 Tmax=22 - pump shut down for 15min Huber
  HFS-7_dlf_502-3 Gollum   -130.0136 45.9336 421422 5087166 3:27 Sterivex DNA filter #7- Tmax=20.2 T2=16 - 1450ml pumped Moyer
  HFS-13_dlf_R502-4 Gollum   -130.0136 45.9336 421422 5087166 3 micron + sterivex filters for DNA - Tmax=23.5 - 1029mL Huber
  HFS-15_dlf_R502-5 Gollum   -130.0136 45.9336 421422 5087166 3:49 0.45 micron lipid filter #15 for DNA - Tmax=24 - 1426 ml Huber
  HFS-22_dlf_R502-6 Gollum   -130.0136 45.9336 421422 5087166 3:56 Gas piston #22 - Tmax=22.3 - 640ml Butterfield/ Huber/
  SS-j4_bio_R502-7 Gollum   -130.0136 45.9336 421422 5087166 4:00 Slurp sample for palm worms et al - also for fluid Juniper/ Levesque
  MTR_dlf_R502-8 Gollum   -130.0136 45.9336 421422 5087166 4:30 Recover MTR deployed on Dive R471 1998 Embley
  HFS-24_htfl

_R502-9

Marsh-mallow   -130.0136 45.9337 421420 5087179 4:35 Gas piston#24 - Tmax=74.3 Evans Butterfield/ Lilley
HFS-21_htfl

_R502-10

Marsh-mallow   -130.0136 45.9337 421420 5087179 4:45 Lipid filter #21 - Tmax=74.4 - 1200 ml Huber
  HFS-12_htfl

_R502-11

Marsh-mallow   -130.0136 45.9337 421420 5087179 4:55 3 micron + sterivex filters for DNA - Tmax+75.8 - 1000 ml Huber
  HFS-9_htfl_

R502-12

Marsh-mallow   -130.0136 45.9337 421420 5087179 5:09 Bag with POC filter #9 - Tmax =73.6 - 599 ml Butterfield
  HFS-4_htfl

_R502-13

Marsh-mallow   -130.0136 45.9337 421420 5087179 5:10 Piston sample #4 - Tmax 71.8 - 601 ml Butterfield/ Huber/
  HFS-1_htfl

_R502-14

Marsh-mallow   -130.0136 45.9337 421420 5087179 5:15 FISH filter #1 - Tmax 80.7 - 1000 ml Huber
  SS-j3_bio

_R502-15

Marsh-mallow I -130.0136 45.9337 421420 5087179 5:27 Suction sample of fauna (tubes and snails) - fluid Marcus
  GTB-7_htfl

_R502-16

Virgin   -130.0135 45.9336 421430 5087174 6:00 Sttbd gastight - Tmax=306 Evans Butterfield/ Lilley
HFS-18_htfl

_R502-17

Virgin   -130.0135 45.9336 421430 5087174 6:02 Bag #18 with filter - Tmax=309 T2=145 Butterfield/ Graham
  HFS-14_htfl

_R502-18

Inferno   -130.0139 45.9335 421397 5087162 6:54 Bag #14 with 0.45 filter - Tmax 289.9 - 406 ml Butterfield/ Graham
  HFS-16_htfl

_R502-19

Inferno   -130.0139 45.9335 421397 5087162 6:59 Unfiltered bag #16 - Tmax 293 - 500 ml Butterfield/ Graham/
  GTB-2_htfl

_R502-20

Inferno   -130.0139 45.9335 421397 5087162 7:08 Port gas tight #2 Evans Butterfield/ Lilley
HFS-2_htfl

_R502-21

Inferno   -130.0139 45.9335 421397 5087162 7:10 0.45 micron filter #2 for XRF - Tmax=289.6 - 476 ml Gendron
  HOBO_htfl

_R502-22

Inferno   -130.0139 45.9335 421397 5087162 7:34 Retrieved HOBO 137 deployed on Alvin dive 3245 in '98 - scrapings for Embley Scott
HFS-20_htfl

_R502-23

Crack   -130.0136 45.9333 421424 5087135 8:09 Piston #20 - Tmax=89.7 - 628 ml Butterfield/ Huber/
  HFS-10_htfl

_R502-24

Crack   -130.0136 45.9333 421424 5087135 8:16 Filter #10 for lipids - Tmax=90 - 1100 ml Huber
  Niskin_htfl

_R502-25

Hell   -130.0142 45.9333 421372 5087135 8:42 Niskin bottle, 1 meter above Hell Gendron
  HFS-3_htfl_

R502-26

Hell   -130.0142 45.9333 421372 5087135 9:04 0.45 micron filter #3 for XRF - Tmax=294.7 -300 ml Gendron
  HFS-17_htfl

_R502-27

Hell   -130.0142 45.9333 421372 5087135 9:07 Unfiltered bag #17 - Tmax=294.3 - 575 ml Huber
  HFS-19_htfl

_R502-28

Hell   -130.0142 45.9333 421372 5087135 9:11 Bag with 0.45 filter, - Tmax 293.1 - 574 ml Butterfield/ Graham
  HFS-23_htfl

_R502-29

Hell   -130.0142 45.9333 421372 5087135 9:16 Gas piston #23 - Tmax=291.9 - 156 ml Evans Butterfield/ Lilley
SS-j1_bio_R502-30 Porkchop   -130.0142 45.9333 421373 5087136 10:00 Sulfide worms - (scale worms and copepods?) Juniper/ Levesque
  SS-j2_bio_R502-31 Phoenix   -130.0140 45.9332 421391 5087130 10:39 Sulfide worms - palm worms - tube worms Juniper/ Levesque
  SS-4_bio_R502-32 Medusa   -130.0139 45.9334 421395 5087141 10:56 Three anemones - rock - clams Juniper/ Levesque
  SS-1_FeO_R502-33 n/a   -130.0139 45.9332 421395 5087128 11:31 Fe Oxide from mounds - (15 m south of Medusa) - BAD NAV Scott
  SS-4_sf_R502-34 n/a   -130.0139 45.9332 421395 5087128 11:55 Greyish white mat - small chunks of anhydrite Fe- (15 m south of Medusa) Scott Perfit
HFS-11_dfl

_R502-35

n/a   -130.0153 45.9334 421287 5087152 13:15 Bag #11 with filter - no valve, - Tmax=6.2 - 601 ml (west wall) Huber
  SS-2_df

l_R502-36

n/a   -130.0153 45.9334 421287 5087152 13:23 Iron oxides and bacteria (west wall) Scott
  SS-3_dfl_

R502-37

n/a   -130.0153 45.9334 421287 5087152 13:36 Same sample as previous because bottle was contaminated by previous sample Scott
  HOBO_R502-38 Hell   -130.0142 45.9333 421372 5087135 14:08 Collected hobo deployed 1998?? Moyer
   Hell   -130.0142 45.9333 421372 5087135 Deployed HOBO Embley
      Dive R503 SAMPLES
   Sample Vent Mkr Long Lat UTM X UTM Y Time Description Investigator Subsamp
Bactrps_R503-1 Gollum   -130.0136 45.9336 421422 5087166 0:15:00 Recovered 2 bactraps deployed in 1998 on Dive R471 - still one there Moyer
  HOBO_R503-2 Inferno   -130.0139 45.9335 421397 5087162 0:32:00 THIS IS A MISTAKE, ACTUALLY IS SAMPLE 502-22. (FOR SOME Embley
  HOBO_R503-3 Virgin   -130.0135 45.9336 421430 5087174 0:02:41 Recovered HOBO 129 deployed on Alvin Dive 3245 '98 Embley
   Inferno   -130.0139 45.9335 421397 5087162   Deployed HOBO 128 - Top of Inferno Embley
   Inferno   -130.0139 45.9335 421397 5087162   Deployed OsmoSampler - melted it Chapin
   Gollum   -130.0136 45.9336 421422 5087166   Deployed Bactrp #45 Moyer





5.5 Dive Map Nomenclature



The dive maps depict all Vents and Markers visited, samples collected on each dive, in addition all instruments deployed and recovered are also cited.



Nomenclature Example: SS-12_dfl_R503-4



The first letters indicate the sample type. The sample type may be followed by a hyphen and a number, which indicates the number for that specific sample type (could be a HFS filter number, suction sample jar number, etc.)



The letters following the underscore give more information about the sample:

_dlf indicates that the sample was in diffuse flow.



The final number, following the hyphen, designates the dive sample number.

-4 indicates that it was sample number 4 for the dive.





Sample type abbreviations More sample information:

SS Suction Sample mat bacterial mat

HFS Hot Fluid Sampler dfl diffuse flow

Niskin Niskin bottle hfl high temp flow

Gtb Gas Tight Bottle flc bacterial flock

Bactrp Bacteria Trap osmo osmo (sampler or analyzer)

Bio Biological sample HOBO temperature probe (152 - 419C)

Sf sulfide MTR temperature probe (2 - 34C)

Rck rock VEMCO temperature probe (0 - 50C)

FeO iron oxide TLC time lapse camera





Red stars represents vent locations.

Smaller white stars represents marker locations.



5.6 ROPOS Dive Logs, Dives R483 - R502



  Dive R483
      Dive Plan: High resolution survey of Mkr-33 using the Sit camera and Imagenex. Elevator will be used to gather experiments from Mkr-33. Suction sampling will be done there and at other active vents to the north.
      Time-lapse camera will be brought up in the arm of ROPOS.
      ROPOS Configuration: Biobox, Slurp Sampler, Gas Tights, Imagenex, and the other cameras.
      Note: Nav fixes not edited in the logs. All fixes were logged, regardless of accuracy. Use the dive plots to determine actual positions.
  UTC Hour Min Depth (m) Sub Hdng (deg) UTM X UTM Y Comments Frmgrbs, Photos, Samples
     Elevator of the deck and into the water.
  10 8 1 46 414094 5095636 ROPOS is in the water!
  10 24 457 231 414094 5095636 Trying out the frame grab R483-7
10 56 1001 23 414094 5095636 Ropos outside of cage
  11 2 1000 232 414094 5095636 Fixes are still bad
  11 2 1000 232 414094 5095636 Ropos back in cage.
  11 11 1105 214 414094 5095636 Working on problem with video color.
  11 31 1472 327 414094 5095636 Ropos is leaving cage. Near the bottom. Still no navigation.
  11 36 1500 305 423890 5087089 Stopping navigation. Switch channels for transponders.
  11 39 1518 336 423890 5087089 Elevator is on the bottom.
  11 41 1495 65 423802 5087074 Good fix.
  11 45 1491 13 423802 5087074 Elevator deployed. About 20m S of Mkr-33
  11 48 1515 42 423802 5087074 At target called elevator. X=3849 Y=7082
  11 57 1520 217 423802 5087074 Near bottom at Mkr-33. Camera in view.
  11 59 1522 203 423802 5087074 Highlights tape on.
  12 1 1523 206 423802 5087074 Highlights off.
  12 14 1523 119 423849 5087081 Sitting on bottom at Mkr-33. No good sub fixes.
  12 20 1523 123 423849 5087081 Navigation off.
  12 23 1523 117 423848 5087086 Good ROV fix.
  12 25 1523 125 423849 5087086 Time lapse camera and ozmo sampler. R483-8
12 34 1523 131 423848 5087086 Good Fix
  12 35 1521 128 423847 5087102 Highlights on. Doing mosaic
  12 43 1522 282 424108 5087314 Highlights off. Doing mosaic.
  12 45 1523 184 424102 5087915 Still photo of time lapse camera.
  12 46 1523 191 423849 5087082 Still photo 2.
  12 46 1523 193 424244 5087755 Time lapse and crack. R483-12
12 49 1523 197 424422 5086745 Highlights on.
  12 50 1523 196 423852 5087097 Tube worms, palms worms, and scale worms. R483-14
12 51 1523 196 423853 5087097 Same as last framegrab R483-15
12 51 1523 194 423853 5087098 Palm worms, gastropods, tube worms, etc. R483-16
12 53 1523 197 423890 5087617 Same as as last FG R483-18
12 57 1523 204 423889 5087617 Hi
  13 5 1523 186 423853 5087097 Highlights on. Beginning sampling.
  13 5 1523 190 424453 5087458 Sampling J1. Mkr-33 R483-19
13 7 1523 187 423855 5087083 Slurp sample J1. Begin sampling.
  13 9 1523 186 424325 5087985 Slurp sampling J1 for particulates at Mkr-33. First sample J1. R483-20
13 10 1523 190 423847 5087099 Highlights off.
  13 10 1523 189 423852 5087096 Slurp probe sampling J1. R483-21
13 17 1523 188 423854 5087095 Possible samples of worms collected during particulate sampling in J1.
  13 26 1523 192 423842 5087055 Still collecting sample J1.
  13 49 1523 196 423854 5087079 Still sampling J1.
  14 0 1524 190 423853 5087096 VCR tapes changed. Still sampling J1.
  14 6 1523 185 423852 5087098 End sampling J1.
  14 10 1522 185 423852 5087099 Sample area. R483-22
14 10 1521 177 423850 5087094 Still photo.
  14 10 1520 180 423850 5087094 Still photo.
  14 11 1521 203 423855 5087099 Still photo 3rd
  14 12 1520 219 423852 5087098 Camera postion. R483-23
14 12 1521 218 423857 5087082 Photo 4.
  14 12 1522 241 423857 5087082 Photo 5.
  14 13 1522 251 423856 5087096 Photo 6.
  14 13 1522 244 423856 5087096 photo 7.
  14 13 1523 246 423855 5087097 camera position. R483-24
14 19 1523 113 423853 5087099 Picking up time lapse camera to move.
  14 19 1524 109 423854 5087081 Picking up time lapse camera. R483-25
14 23 1525 231 423857 5087095 Camera positioned near elevator anchor.
  14 28 1524 173 423854 5087081 Flush slup sampler.
  14 31 1523 175 423849 5087096 Slup sampling J2 (start).
  14 37 1523 170 423853 5087099 Slurping gastropods. J2 R483-27
14 44 1524 164 423853 5087097 bacterial box from 98 R483-28
14 44 1524 165 423842 5087051 ROPOS reaching in to pick up their box R483-29
14 50 1524 174 423854 5087080 view of vent site, Christian's suction J2 site R483-30
15 7 1523 142 423854 5087095 Finished suction sample J2.
  15 15 1524 201 423848 5087093 Started suction sample J3 - suction top of slab for more polynoids and gastropods
  15 17 1524 196 423854 5087080 Frame grab of suction sample J3 location R483-31
15 47 1524 194 423847 5087098 Bacterial trap collected #12 R483-32
15 54 1524 170 423854 5087082   R483-33
15 59 1524 227 423854 5087080 Tube worm grab sample #1(Marcus) at Mkr-33 R483-34
16 29 1524 203 423845 5087059 Fluid sample of vent at Mkr-33(Evans) sample #1 R483-35
16 37 1524 256 423851 5087099 GTB#6-#7 (Evans) R483-36
16 51 1525 264 423850 5087089 MTR from Mkr-33 (Embely) Probe #4130 R483-37
16 59 1519 228 423843 5087099 Elevator in position above Mkr-33 carrying payload of bacterial traps, bioboxes, osmosamplers and analyzers R483-38
17 1 1516 232 423859 5087094   R483-39
17 2 1516 245 423861 5087096 Elevator positioned above Mkr-33 with payload to be deployed - (bacterial traps, bioboxes, osmosampler and analyzer). R483-40
17 5 1520 240 423853 5087092 Bioboxes for Dr. Moyer R483-41
17 8 1524 246 423852 5087084 Osmo (Fe) analyzer deployment (chapin) set of the vent site for deployment within the vent later at Mkr-33 R483-42
17 23 1524 311 423840 5087045 Fe Osmo Analyzer (Chapin) Mkr-33 R483-43
18 11 1524 262 423847 5087099 Green Osmo sampler and Copper osmo sampler (Chapin) deployed at Mkr-33 R483-44
18 36 1524 203 423851 5087076 Sampler and analyzer probes w/in the vent fluid(Chapin) R483-45
18 39 1522 192 423850 5087091 Green osmo sampler with Copper osmo sampler in foreground with the the Fe Osmo Analyzer in the background at Mkr-33 R483-46
      Dive Summary: Went to Mkr-33 where 8 samples were collected (Suction samples, bactraps, bio, and gas-tights). Recovered MTR and Osmo from NeMO '98 cruise. Two Fe OsmoAnalyzers were deployed. Dive was aborted due to a kink in the tether.


      Dive R484
        Objectives: 1) Deploy other osmosampler from elevator at northern end of Mkr-33. 2) Collect basalt sample at Mkr-33. 3) Suction sample bacterial mats at Mkr-33 for Moyer, gastropods and mat for Levesque.
       Suction sample bacterial mats at Cloud for Moyer, vent fauna larvae for Tunnicliffe. Suction sample for diatom mat on new lavas for Tunnicliffe. 4)Tube worm grab at Cloud. 5)Short Imagenex survey at Mkr-33 and vicinity.
      6) Recover Time Lapse Camera . 7) Recover elevator with OsmoSamplers and microbial traps.
      ROPOS Configuration: 7 function arm, 5 function arm, Biobox, Suction samplers (7 jars), Gas tights (2), Imagenex and other cameras.
        Note: Nav fixes not edited in the logs. All fixes were logged, regardless of accuracy. Use the dive plots to determine actual positions.
  UTC Hour Min Depth (m) Sub Hdng (deg) UTM X UTM Y Comments Frmgrbs, Photos, Samples
10 21 0 0 0 0 On the way down, the navigation people are trying to find a bug in the nav program and so the heading, depth, X, and Y will not be updated until we are near the bottom.
  10 27 0 0 0 0 ROPOS went in the water at 10:15.
  11 36 1490 253 423846 5087083 ROPOS leaving the cage.
  11 37 1522 267 423847 5087092 ROPOS is on the bottom
  11 43 1523 272 423843 5087086 Frame grabber locked up.
  11 44 1522 14 424199 5086776 Navigation still not good.
  12 18 1523 270 423858 5087087 Logging off since 11:55. Begin ss1 at 11:55, target is mat material. Video bad. Nav bad. Sampling stopped to flush at 12:11. 12:16 repositioning sub and trying to fix clog in the line. x=423860 y=5087105 SS_mat_R484-1
12 23 1523 256 423869 5087103 still flushing slurp system into bottle 4.
  12 27 1520 181 423860 5087093 ROPOS moving to attempt to capture a crab.
  12 29 1522 165 423471 5086545 Target crab. R484-1
12 29 1524 174 423860 5087091 Target crab. R484-2
12 31 1522 156 423630 5086828 Into the box. R484-4
12 33 1523 123 423865 5087089 Final solution. R484-5
12 34 1524 126 423899 5086698 Crab collected. Bio_R484-2
12 35 1524 121 424126 5087648 Crab escaped. Sample lost.
  12 39 1524 84 423868 5087093 Second attempt to collect same crab.
  12 40 1524 74 423871 5087114 Crab escaped for the second time.
  12 42 1524 61 423865 5087100 Crab capture completed. Crab in port bio-box.
  12 43 1523 359 423864 5087088 ROPOS returning to cage.
  12 51 1520 2 423595 5086465 Retrieving a bacterial trap from the elevator.
  12 52 1522 114 423861 5087095 Bacterial trap was #42. Moving to deploy it.
  12 56 1523 304 423857 5087097 #42 deployed. R484-7
12 56 1522 308 423858 5087097 Bacterial trap 42 deployed at the end of the crack at Mkr-33.
  13 0 1519 78 423860 5087082 Collecting another bacterial trap #41 from the elevator.
  13 4 1522 171 423858 5087086 Deploy #41 at the same site as #42.
  13 9 1520 155 423860 5087098 Collecting bacterial trap #40 from the elevator.
  13 11 1523 267 423856 5087103 Location of bacterial trap #40. R484-8
13 12 1521 14 423853 5087097 Bacterial trap #40 deployed near OsmoSampler, middle of the crack.
  13 17 1519 160 423865 5087100 Opening Moyer bio-box inside elevator.
  13 19 1520 23 423859 5087083 Collected bacterial trap #43 from elevator.
  13 21 1523 280 423856 5087097 Position of bacterial trap 43. R484-9
13 22 1523 243 423856 5087097 Bacterial trap #43 deployed.
  13 23 1523 245 423860 5087103 Attempting to clear the clogged suction sampler.
  13 28 1520 88 423855 5087092 Back at the elevator collecting 4 bacterial traps.
  13 36 1520 310 423861 5087081 Moving bactrp 36 from the elevator to the strb sub bio-box.
  13 38 1520 268 423862 5087091 Transferring bactrp #39 to sub strb bio-box.
  13 40 1520 203 423860 5087095 Transferring bactrp #38 to sub strb bio-box.
  13 40 1519 194 423857 5087099 Begin changing VCR tapes.
  13 42 1519 193 423862 5087094 Finish changing VCR tapes.
  13 43 1520 282 423860 5087093 Transferring bactrp #37 to sub strb bio-box.
  13 45 1517 24 423937 5086713 Moving to Cloud vent.
  13 45 1522 71 423937 5086713 Nav fixes bad.
  13 51 1517 54 423899 5087123 Nav fixes good.
  13 52 1520 125 424740 5086697 Still looking for Cloud vent. Fixes not good. Found area with worms and warm water.
  13 55 1523 137 423901 5087114 Nav fixes good.
  13 56 1523 137 423909 5087120 Worms and warm water just north of Cloud Vent. Fixes bad. R484-10
13 58 1523 134 423901 5087114 Fixes are good. Cloud 5 meters away.
  14 2 1525 92 423644 5086924 Looking for Mkr-N4, trying to find bactrps.
  14 4 1519 92 424532 5088000 Collapsed area. Fixes are bad.
  14 5 1520 99 423924 5087108 Good Fix.
  14 5 1523 267 423914 5087115 Tube worms. Good fix. R484-11
14 6 1522 275 423904 5087108 Cloud vent Mkr-N4.
  14 6 1524 282 423901 5087116 Mkr-N4. Cloud vent. Tube worms and scale worms. R48412
14 8 1523 274 423901 5087115 Deploy bactrp # 37 at Cloud
  14 10 1524 281 424429 5087476 Deploy bactrp #36 at Cloud vent Mkr-N4. x=423896 y=5087119
  14 14 1524 272 423901 5087111 Recover bactrp #14 into strb bio-box bactrp_R484-3
14 21 1524 260 423625 5086916 Remove bactrp #39 from box and deploy next to Mkr-N4.
  14 25 1524 262 423893 5087118 Attempting to recover an MTR sample. MTR_R484-4
14 28 1524 244 423899 5087114 Finish MTR recovery. Into strb bio-box.
  14 30 1524 252 423898 5087116 Recovering bactrp (can't read number) into strb bio-box. bactrp_R484-5
14 36 1524 253 423897 5087102 Deploying bactrp #38 at Mkr-N4.
  14 39 1523 267 424434 5087448 Moving around to survey for worms.
  14 42 1524 262 423970 5086736 Working to clear suction sampler.
  14 45 1524 262 424049 5087913 Mkr- N4 positions of bactraps. R484-13
14 47 1522 307 423997 5087171 Moving around to look for worms.
  14 48 1522 257 424403 5087529 Nav fixes bad.
  14 49 1522 200 423891 5087118 Nav fix good.
  14 53 1486 237 423965 5087100 ROPOS back to cage.
  14 55 1490 87 424082 5087584 ROPOS going back to bottom.
  14 58 1524 102 423844 5087084 Good fix. First good one in awhile. Heading SW to start of Imagenex line.
  15 1 1518 204 424438 5087440 Stopped, trying to get a fix, bad fixes right now.
  15 2 1518 203 423838 5087002 Good fix, 40 m south of beginning of 1st imagenex transect, moving to start point
  15 55 1479 281 423932 5087100 Still maneuvering for start of line. Problem with bow thruster. Imagenex survey aborted. Weather not good enough to recover elevator. Will pick up time-lapse camera and come to surface.
  16 34 1384 284 423875 5087107 Decided not to reattempt imagenex survey; heading to bottom to get time-lapse camera
  16 44 1523 228 424360 5087689 At elevator, about to recover the time-lapse camera (TLC)
  16 46 1523 153 423854 5087090 Quickly flying back over Mkr-33 to make sure osmosampler intakes are correctly positioned
  16 55 1524 16 423856 5087073 Nav has been bad for around 10 minutes
  16 56 1524 337 423852 5087087 Good fix
  17 4 1524 235 424449 5087467 While checking out the osmosamplers, the bacteria traps #42 and # ? got dislodged, so we are now repositioning them. The ship is drifting to the NE, trying to reposition ship.
  17 9 1514 253 424426 5087486 ROPOS going back to cage. Waiting for ship to reposition.
  17 12 1465 114 424376 5087657 Back at cage. Waiting.
  17 18 1453 206 423926 5087108 Ship repositioned. Going to get Time Lapse Camera.
  17 24 1515 285 424482 5087421 Waiting for a good fix.
  17 24 1517 285 423890 5087082 Good fix.
  17 25 1520 315 424491 5087438 At Mkr-23. Going to get TLC at base of elevator.
  17 26 1523 283 423964 5087122 Disregard previous comment - NOT at Mkr-33, but came down at Mkr-N8.
  17 27 1520 308 424711 5086887 At the elevator. SEE TLC!
  17 30 1525 307 423810 5086523 Picking up the Time Lapse Camera - bad fix.
  17 35 1525 291 423846 5087093 FG of grabbing the Time Lapse Camera. R484-14
17 39 1525 283 424442 5087475 The Queen says NAV SUCKS!
  17 40 1526 283 424441 5087473 Trying to jettison the 65 lbs from ROPOS in order to pick up camera.
  17 42 1526 282 423849 5087346 Weight released!
  17 48 1526 284 425147 5086738 Picking up TLC. x=423850 y=5087101 R484-15
17 48 1526 283 424443 5087473 TLC being positioned in claws to come up.
  17 51 1526 283 424445 5087476 TLC ready to come up - positioned in claws properly. The jettison weight is still attached to ROPOS on port side.
  17 57 1525 309 424889 5087523 WE just lost sub power.
  17 58 1525 309 424889 5087523 Sub power on then off again.
  18 0 1525 307 424889 5087523 Power and lights back on - video up. Sitting on the bottom with TLC in claws. Discussing what to do.
  18 1 1525 305 424889 5087523 Power off, back up again.
  18 4 1524 291 424889 5087523 Flying backward trying to dislodge jettison weight.
  18 6 1514 241 423892 5087061 Lost lights again. Coming up to try to dislodge weights.
  18 7 1505 230 423887 5087047 Weights gone. Coming up to the surface.
  18 13 1445 241 423914 5087082 Time lapse camera recovery should have the sample number R484-6. TLC_dfl_R484-6
20 15 108 155 424102 5086984 Puled ROPOS into the cage with the tether winch in first attempt. Great driving by Keith Shepherd. Bringing cage up fast.
  20 27 1 172 424230 5087291 ROPOS on deck at 1323h.
  20 47 1 160 424230 5087291 Time lapse camera appears to be OK. Crab is dead. Stainless steel bolt in a bactrap was almost eaten through.

END OF DIVE R484

      Summary of Dive R484: For about the first half hour of the dive the log entry program was shut down. During the collection of the first suction sample it was rebooted.
      Mkr-33: SS of mat; one nearly whole crab captured, TLC recovered, deployed bactraps. Cloud: Two bactraps and one MTR recovered, deployed bactraps, short worm survey.
      Suction sampler clogged on first sample. Color balance on frame grabs was poor. The sub lost hydraulic fluid which forced the end of



      Dive R485
        Dive Objectives: Primary objective of the dive is to obtain vent fluids/filtered samples along eruptive line. Dive will begin at Mkr-33 and end at the Milky Vent area. Suction samples will also be done at some of the sites.
        Elevator will be recovered (weather permitting) sometime during the dive.
        ROPOS Configuration: Vent Fluid Sampler, Suction Sampler, Gas Tights, Imagenex, and cameras.
        Note: Nav fixes not edited in the logs. All fixes were logged, regardless of accuracy. Use the dive plots to determine actual positions.
  UTC Hour Min Depth (m) Sub Hdng (deg) UTM X UTM Y Comments Frmgrbs, Photos, Samples
     ROPOS in the water
  7 6 88 22 424231 5087291 This is the real launch time, the previous entry was wrong.
  8 10 1315 140 423873 5087133 A line got caught in the cage, preventing ROPOS from leaving the cage, and we're going to start back up.
  8 13 1239 156 423853 5087110 Correction. A line connecting the vehicle to the cage on deck was not removed, preventing ROPOS from leaving its cage. We're still coming up. Slight delay.
  9 13 1 156 423814 5088795 ROPOS came back up in order to remove the rope attaching it to the cage.
  9 36 0 123 423814 5088795 ROPOS in water again.
  10 53 1520 234 423682 5087227 ROPOS is on the bottom. Two fish. Near Mkr-33.
  10 55 1521 110 424415 5087823 Fixes no good.
  10 57 1518 8 423858 5087078 Looking for Mkr-33. ROPOS is about 24 m away from Mkr-33 according to nav.
  10 59 1523 331 423765 5087178 Fixes still no good.
  11 0 1521 17 424488 5086783 Arrive at Mkr-33. Good fix.
  11 2 1522 344 423856 5087103 Good fix. Frame grab color is good.
  11 8 1522 248 423856 5087105 Positioning HFS (Hot Fluid Sampler) probe for sampling. R485-1
11 9 1522 248 424412 5087823 Temp up to 63C during probe positioning for HFS.
  11 12 1522 228 423850 5087572 Repositioning sub. Trying to find probe position with highest Temp reading.
  11 14 1523 211 423853 5087104 Temp at about 58C
  11 15 1523 211 423861 5087105 Position of probe during sampling. R485-2
11 16 1523 210 423859 5087083 Starting sample HFS2 HFS2_dfl

_R485-1

11 17 1523 213 423842 5087728 temp 5 intake 40 on manifold. Good temps.
  11 19 1523 212 424399 5086596 Range on T2 is 37 to 49.
  11 20 1523 210 423888 5087631 Stop HFS2. Volume sampled was 56ml.
  11 22 1523 213 423857 5087108 HFS23 Piston sample for gas tight started. T1 53 HFS23_dfl

_R485-2

11 23 1523 210 423890 5087613 Done T2 42-50. vol 13 ml.
  11 25 1523 213 424443 5087470 HFS12 DNA sample. T1 = 49 T2 is 37. HFS12_dfl

_R485-3

11 34 1523 211 424410 5087494 Sample finished. max T 54.6 vol = 123ml
  11 35 1523 213 423852 5087105 HFS14 started. Bag sample. HFS14_dfl

_R485-4

11 41 1523 212 423574 5087340 Sample done vol = 10ml.
  11 43 1523 213 423860 5087082 HFS17 bag sample, no filter. Max T on last bag sample was 48.8C. HFS17_dfl

_R485-5

11 45 1523 214 424490 5088046 Video has gone bad due to bright input to overlay.
  11 45 1523 214 424287 5086270 Video OK.
  11 46 1523 212 423858 5087093 Video bad.
  11 46 1523 211 423844 5087062 Sample done. Vol = 650ml. Max T 52.5C
  11 48 1522 218 423856 5087104 Lost position. Alarm for oil reservoir empty suddenly. Shut down.
  11 50 1522 230 424104 5087602 ROPOS returning to surface due to loss of oil. Ship moving forward in order to stream ROPOS behind cage.
  11 55 1492 252 425326 5086999 VCR tapes turned off.
  13 56 186 3 422757 5088424 Begin returning ROPOS to cage.
  14 55 107 216 422710 5089068 Still having great difficulty getting the sub back into the cage.
  16 25 89 112 422710 5089068 The cage is being brought to the surface. An attempt will be made to get the sub into the cage just below surface. The earlier attempts to get the sub in the cage at around 20m did not work.
  16 49 128 123 422710 5089068 ROPOS was driven back into the cage at 128 meters, not at the surface.
  16 59 0 115 422710 5089068 ROPOS is back on deck.
      Dive Summary for R485:

ROPOS was in the water at 0706, 25 June 9. When the cage reached 1315m it was determined that ROPOS was still tied to the cage by a rope. ROPOS returned to the surface and the rope was removed. Dive was restarted at 0936.

      ROPOS reached the bottom at 1053 and Mkr-33 was located at 1100. The sampling probe for the fluid sampler was positioned in the crack, based on the highest temperature measurement. This also appeared to be the area of highest flow.
      Five HFS samples were collected. At 1148 the sub was shut down due to a loss of hydraulic oil. ROPOS returned to the surface, after some difficulty getting the sub back into the cage.



      Dive R486
      Dive objectives: Continue the dive plan from dive R485. More fluid sampling, slurp sampling, reposition bacterial traps, recover animals. Imagenex survey at end of dive.
       ROPOS configuration: Vent fluid sampler, Slurp sampler, gas tights, Imagenex, cameras.
       Note: Nav fixes not edited in the logs. All fixes were logged, regardless of accuracy. Use the dive plots to determine actual positions.
  UTC Hour Min Depth (m) Sub Hdng (deg) UTM X UTM Y Comments Frmgrbs, Photos, Samples
3 6 6 50 424231 5087291 ROPOS in the water
3 19 172 317 423702 5087374 Overlay on video turned on. Press "Super" for superimpose.
3 21 212 319 423985 5087583 going down
3 30 378 40 424411 5087396 narf, what am I doing here? R486-1
4 38 1448 332 423897 5087219 Drove ROPOS back in cage. Pulling up. Going through pretty thick plume. Water is cloudy. Lost oil pressure before made it to the bottom.
4 41 1379 327 424376 5086680 Bringing ROPOS back up. Abort dive.
5 36 1 162 421199 5087374 ROPOS on deck
5 37 2 156 421199 5087374 Dive Summary: ROPOS descended within 10 meters of the bottom when it was forced to return to the surface due to a problem with the hydraulics.


      Dive R487
       Dive Plan: Same as for R485 and R486. Primary objective is to obtain vent fluids/filtered samples along eruptive line. Will begin at Mkr-33 and end at Milky Vent area. Suction sampling will also be done at some of these sites.
       Elevator wil be recovered (weather permitting) sometime during the dive.
       Note: Nav fixes not edited in the logs. All fixes were logged, regardless of accuracy. Use the dive plots to determine actual positions.
  UTC Hour Min Depth (m) Sub Hdng (deg) UTM X UTM Y Comments Frmgrbs, Photos, Samples
15 32 1 133 421199 5087374 ROPOS in the water.
15 43 85 256 421199 5087374 test Dive 487-1
17 6 1519 87 424091 5087593 On the bottom. At Mkr-33. Viewing osmosampler. Photos 2 - 5
17 7 1521 39 424091 5087593 Osmo Sampler at Mkr-33 Dive 487-2
17 8 1521 126 424091 5087593 Taking pictures of Mkr-33. Right at the marker - viewing bacteria trap 40 and green/copper osmosampler. Took Photos 7 - 10
17 10 1524 231 423856 5087075 Checking out ROPOS gauges - making sure all is ok. Hydraulic fluid down, talk of going back to the cage.
17 12 1503 227 424345 5087694 Slowly going back to cage.
17 16 1492 176 423852 5087102 Taking a plume sample with the HFS - filter #12.
17 18 1482 206 423852 5087102 ROPOS in the cage. Staying here to take some samples of the plume. Plume filter sample for DNA with HFS - starting filter #12. HFS12_plume _R487-1
17 23 1475 178 424653 5086900 Going to stay here for around 15 minutes to take a few samples. Still pumping HFS filter #12.
17 30 1475 226 424860 5087225 Something in water column - a jellyfish I think. Photo-13
17 32 1474 116 423852 5087100 Stopped pumping filter #12 - 1502 ml pumped. Starting to pump filter #15 on HFS. HFS15_plume _R487-2
17 44 1475 256 423869 5087113 Stopped pumping HFS filter #15 - pumped 1050 ml.
17 45 1475 301 423890 5087611 Starting HFS bag sample #17.
17 51 1476 300 425149 5086724 Stopped pumping #17 - 650 ml pumped. Start pumping HFS #19 - bag with filter. HFS19_plume _R487-4
17 58 1475 207 424342 5087672 Stopped pumping #19 - 675 ml. Note that the flush pump was not on for the first half of the sample. Start to sample HFS bag #16 with no filter. HFS16_plume _R487-5
18 2 1475 226 423937 5087439 Good nav. x=3846, y=7105 - location of plume fluid sampling.
18 5 1474 212 424117 5087620 Finished pumping #16 - 598 ml pumped. Start pumping HFS filter #3, for sulfur. HFS3_plume _R487-6
18 9 1477 290 423859 5087130 Fluid volume was not recording for a few minutes - now working again.
18 19 1475 227 424340 5087684 Stopped pumping filter #3 - 130 ml pumped. Going to move the ship to a position over Cloud Vent to sample particulates in plume.
18 27 1476 217 425095 5087131 The Queen says BAD FIXES!
18 28 1476 248 423872 5087114 Moving towards Cloud - still no good fixes.
18 34 1476 205 423874 5087096 We are about half way to Cloud.
18 39 1476 275 423882 5087105 Ship finished moving - waiting for the position to settle.
18 40 1474 257 423900 5087099 Good fix at x=3898, y=7110. Between Mkr-N4 and Mkr-N6.
18 42 1474 202 424440 5086790 x=3900, y=7110 - good fix on ROPOS in cage. We are stopping here to take fluid and gas tight samples. We are flushing the gas line now.
18 44 1476 205 423907 5087096 Right over Mkr-N6.
18 46 1474 352 423900 5087100 The gas line in flushed. Preparing to trigger port gas tight - triggered port GTB at 18:47. gtb_plume _R487-7
18 48 1475 205 423897 5087098 Going to trigger the gas piston on the HFS, pump position #23. Starting to pump. Good fix x=3901, y=7111. HFS23_plume _R487-8
18 53 1476 280 423894 5087116 Triggering piston #20. HFS20_plume _R487-9
19 0 1474 198 423902 5087123 Sample R487-9 finished. 730 ml.
19 3 1476 222 423936 5087610 X = 3908, Y = 7109
19 5 1474 278 423898 5087100 Filter #21 for sulfur analysis started at 19:01. HFS21_plume _R487-10
19 10 1476 127 424385 5087640 X = 3902, Y = 713. Sample #21 finished. Piston sample #2 started. HFS22_plume _R487-11
19 17 1476 202 424442 5086790 Sample piston #2 finished. Start bag sample #18 with filter. Still at Cloud Vent. HFS_plume _R487-12
19 24 1475 267 423902 5087115 Sampler #18 finished. 60 ml.
19 25 1476 331 424385 5086607 Sampler #7 with filter started. HFS7_plume _R487-13
19 29 1474 136 423907 5087096 #7 finished. 41 ml.
19 29 1475 140 424444 5086797 Sampler #2 with filter started. HFS2_plume _R487-14
19 33 1476 323 423898 5087098 Sampler #2 finished. Ready to come to surface. Will take samples on the way up.
19 35 1474 285 423904 5087099 Starting integrated bag sampler #1 during rise through the plume because it is "fun" (Butterfield) HFS11_plume _R487-15
19 44 1254 274 423910 5087482 Stopped sampler at 19:41. Continuing to surface.
20 32       ROPOS on deck.
       Dive summary: Hydraulic failure shortly after ROPOS reached bottom. ROPOS put back into its cage from where it sampled the plumes at about 1475 m depth above Mkr-33 Vent and Cloud Vent. Returned to surface for repairs.


       Dive R488
        Dive Objectives: Same objectives as dives R485 - R487. To Mkr-33 for HFS and suction sampling, then on to Cloud and vents to the north on the '98 lava flow. Biology samples will also be collected.
       At the end of the dive and Imagenex data will be collected at the western edge of the '98 lava flow.
       Note: Nav fixes not edited in the logs. All fixes were logged, regardless of accuracy. Use the dive plots to determine actual positions.
  UTC Hour Min Depth (m) Sub Hdng (deg) UTM X UTM Y Comments Frmgrbs, Photos, Samples
4 27 1 138 423861 5087236 ROPOS in the water
5 38 1094 22 423822 5087115 Grape Jellyfish
5 59 1485 89 423822 5087132 On the way down, near bottom, in the hydrothermal plume. Water is cloudy.
6 1 1509 90 423812 5087119 On bottom. Glassy looking lobate flows.
6 3 1517 173 423826 5087129 Lava whorl
6 3 1522 262 423830 5087130 Lava whorl in sheets R488-3
6 4 1521 134 424662 5087328 Sheet flow with bacterial mat between cracks. Ropy sheets some lineations. Some fractures and steps in more or less continuous flow surface.Yellowish sed. thinly coats surface of flow but white mats concentrated on some cracks.
6 7 1520 133 423886 5087090 Looks like a possible contact with more pillowed forms in distance.
6 9 1523 276 423876 5087090 Heading west along flat sheets. Elevator in distance. Photo taken. First 4 photos were not logged here. Photo
6 11 1521 248 423867 5087067 wandering around elevator.(3 photos of elevator)
6 15 1525 190 423841 5087113 All bad fixes since we got to the bottom. More photos here and frame grab. (photo of osmoanalyzer) R488-3
6 17 1519 135 423829 5087115
 6 18 1521 212 423831 5087122 Mkr-33 FG R488-5
6 18 1520 213 423831 5087122 Mkr-33 crack in pushup in lava. R488-6
6 18 1523 210 423839 5087114 Crack at Mkr-33 R488-7
6 18 1524 213 423839 5087114 Getting close to vent in crack. Photo of vent, close up photo of venting in fissure.
6 19 1525 217 423573 5087074 Fluid sample site at Mkr-33 R488-8
6 19 1524 215 423839 5087108 Close up of vent. R488-9
6 22 1525 216 423852 5087097 Putting the fluid sampler in the vent. Temp up to 24 C. 32, 34, 35.40 max. White coating on sides of flow, scale worms and tube worms on edges and under flow. Temps going up. 68, 70. Sampling bag 14 HFS-14_dfl

_R488-1

6 27 1525 215 423837 5087112 When sample 1 HFS was taken. 70 C water T. R488-10
6 28 1525 215 423756 5087359 Diffuse flow hot water sampling, temps in upper 60's.
6 34 1525 218 423838 5087110 Sample stopped at 6:34. Max temp 70.4
6 35 1524 217 423838 5087111 Another sample from same site. Filter 6.  



HFS-6_dfl _R488-2
6 38 1525 216 423832 5087121 Temp as high as 73.2 C at this site during sampling. frame grab taken here. Sampling filter 6
6 39 1525 218 423717 5087586 Sampling At Mkr-33 with the old HFS R488-11
6 42 1525 217 424658 5087313 Bad tearing on the video.
6 45 1525 214 423838 5087108 lost video signal
6 48 1525 216 424662 5087314 Stopping sampling, max 73.2 degrees.
6 49 1525 218 423832 5087109 Going to sample filter 15 of fluid sampler starting at 6:50 max temp 73.9 C. HFS-15_dfl _R488-3
6 51 1525 218 423291 5087263 Worms galore at Mkr-33 R488-12
6 52 1525 214 423838 5087112 Sampling temp rising up to 78.
6 56 1525 218 425062 5086927 More worms at Mkr-33 R488-13
7 3 1524 215 423849 5087100 Stopping sampling. 1.5 liters. Taking stb. gas tight. Same diffuse vent site. 68.8 C. gtb-7_dfl_R488-4
7 5 1525 216 423837 5087112 Yet more worms at Mkr-33 R488-14
7 5 1525 216 423836 5087112 Ended gas tight.
7 10 1525 214 423829 5087128 R488-15
7 10 1525 215 423836 5087113 Photo of palm worms at Mkr-33. Now we're looking at the honey bear in the slurp sampler - I'm sorry I gave you bad MOJO. Photo-13
7 13 1525 220 423840 5087111 Slurping up palm worms from the crack at Mkr-33. Started sample at 07:15. SS-j1_bio

_R488-5

7 19 1524 211 423837 5087113 Frame Grab of suction sampler sucking up palm worms at Mkr-33. R488-16
7 19 1524 213 423837 5087113 Suction of worms at Mkr-33 R488-17
7 22 1525 214 423890 5087319 Looking at bottle J4 at the palm worms-suctioning more.
7 27 1525 218 423838 5087114 Still trying to get more worms and bacterial mat in bottle J1.
7 30 1524 216 423838 5087110 Honey Bear in J1 at Mkr-33 R488-18
7 32 1525 222 423838 5087111 Still sucking palm worms into bottle J1.
7 41 1525 218 423838 5087112 Spewing rocks out of the suction sampler hose.
7 42 1524 217 423838 5087113 Done with Jar J1. Flushing the line.
7 43 1524 214 423830 5087140 Oops, honey bear is in jar #4 R488-19
7 44 1525 218 423823 5087105 Checking the suction sampler for a possible problem-trying to flush.
7 48 1525 218 423827 5087120 We are successfully flushing - the bear is dancing.
7 52 1525 215 423837 5087115 We're in bottle #3 to suction bacterial mat in the crack of Mkr-33. SS-3_mat _R488-6
7 55 1525 231 423826 5087135 Suction sample of bacterial mat at Mkr-33 R488-20
7 56 1525 229 423826 5087135 Suctioning up bacterial mat and a few worms at Mkr-33 R488-21
8 2 1524 244 423839 5087106 Because we're not getting our fourth transponder, we're getting good fixes but they're offset from the real position by 20 m to the northwest.
8 2 1524 243 423840 5087113 Jar #3 with bacterial mat from Mkr-33 R488-22
8 3 1524 245 423834 5087113 Still sucking on bottle 3.
8 6 1524 216 423837 5087107 Moving elsewhere along the crack to get more bacterial mat in Jar 3. Moved 1 m to the side.
8 11 1524 226 423820 5087110 Suction bottle 3 full of bacterial mat and worms. R488-23
8 11 1524 225 423820 5087110 Now we're getting ready to go to Snail Vent to collect snails! 30 m at 110 will take us there.
8 13 1524 226 423829 5087137 Saw a scale worm (Lepidonotopodium) covered in bacterial mat- it has no gills.
8 15 1523 89 423838 5087122 Ropy sheet flows.
8 17 1523 91 423718 5087582 Sheet flows sort of like broken up shelf steps, covered in orangey iron rich sediment/bacterial goop. Looks primarily inorganic.
8 19 1523 354 423846 5087122 We can see some more venting in a crack here - looking for Snail still, but we are investigating this crack with diffuse flow coming out. There are lots of gastropods and tube worms.
8 21 1524 328 423852 5087127 Problems with video.
8 21 1525 328 423851 5087128 Photo of snail vent. The sulfide worm is here- sorry, they're not tube worms. Photo-14
8 22 1524 319 423558 5087585 Snail vent w/ worms. R488-24
8 22 1524 319 423558 5087585 Worms at Snail vent R488-25
8 23 1524 313 423855 5087126 Same worms at Snail vent R488-26
8 23 1525 312 423856 5087122 We see Paralvinella sulfincola or pandorae? We're not sure. Photo-16
8 24 1525 307 423845 5087143 Photos of worms. We're pretty sure this isn't Snail, it's just further down the crack from Mkr-33.  



Photo-17
8 26 1525 304 423857 5087124 We're going to take a few minutes to look for the real Snail Vent now since we don't have the right mesh for this.
8 27 1526 350 423855 5087124 Dead tube worms covered with bacteria near Snail vent. R488-27
8 27 1524 64 423852 5087133 Lumpy lavas - looks pretty lobate. We're seeing some dead tube worms overgrown with bacterial white gooey mat. There is orange stuff floating around. Looking at some of these cool lava arches. Very cool. Photos 18 - 20
8 28 1524 79 423857 5087127 Still searching for Snail. We think we've found it.
8 28 1524 79 423857 5087127 Still searching for Snail. We think we've found it.
8 30 1520 109 423856 5087147 Drained lava lake, lava pillars, lots of neat lava structures.
8 31 1523 110 423858 5087142 We are at Snail! Right near Mkr-N8. There are tons of snails and scale worms too. Photo-21

Photo-22

8 32 1525 73 423872 5087105 We can see some pretty intense webby mat, bag creatures, snails, scale worms.
8 33 1524 64 423872 5087112 We will be suction sampling fauna into Jar J4. SS-j4_bio_ R488-7
8 36 1525 43 423776 5087656 Preparing to sample from Snail Vent.
8 37 1525 14 423851 5087145 Suctioning snails from Snail Vent. R488-28
8 39 1525 357 423851 5087153 Commencing suctioning at Snail Vent.
8 40 1525 3 423848 5087148 Suctioning Snails at Snail Vent R488-29
8 43 1525 63 424463 5086862 Sample R488-7 is looking god.
8 45 1525 68 423884 5087067 The lava is quite flimsy - lots of bacterial action. Note: Julie is using sound effects.
8 48 1525 74 423776 5087352 Still suctioning and spewing at Snail Vent.
8 50 1525 77 424160 5088133 Filter on J4 looks clogged. Trying for a few more snails.
8 52 1525 76 423853 5087141 Suctioning snails at Snail Vent R488-30
8 58 1525 76 423852 5087150 Finished sampling Jar J4.
9 2 1525 78 423842 5087049 Taking a temperature measurement at Snail Vent with the Fluid Sampler temperature probe.
9 3 1525 79 423845 5087056 Max temp is 13.0 C
9 7 1525 77 424571 5087286 We are now making our way to Cloud Vent for Fluid Sampling and Suction Sampling.
9 9 1524 13 423842 5087048 Just finished suction sampling at Snail, and are heading to Cloud vent.
9 11 1522 59 423491 5087529 At Cloud vent. Photos of the approach. Photo-23

Photo-24

9 12 1524 331 424902 5086973 Fixes are still NOT good. We're on the west side of Cloud.
9 14 1526 336 423868 5087049 At Cloud. There are tube worms (small clumps and sparse individuals) and very dense aggregations of polynoids. Will probe for temperature at this location to decide whether to sample here or at the marker location (mkr N4)
9 17 1526 330 423844 5087048 Worms at Cloud vent R488-31
9 18 1525 341 424119 5086879 Lots more tube worms sparsely distributed in the cracks. Don't see any good grabbable clumps yet though.
9 21 1522 256 423870 5087053 We found Mkr-N4 - there are traps, worms, mat, polynoids and scale worms.
9 22 1525 246 423904 5087109 Mkr-N4 R488-34
9 22 1525 252 423870 5087053 Bacteria traps at Cloud vent, Mkr-N4 R488-32
9 23 1525 245 423841 5087043 We're at the western edge of this vent.
9 24 1525 246 423710 5087512 Take it back- we're looking west. Temp around 8 and 9 C.
9 25 1525 245 423742 5087112 The temperature near the bacterial traps is 18 C. The traps already have grazers on them and they were put down a mere 3 days ago! We just saw the new species of worm!
9 28 1524 258 423839 5087056 Photo of mkr N4. Photo
9 31 1524 355 423841 5087053 Probing for temperature right in between bacteria traps at Mkr-N4. Polynoids in view, and the 'unknown polynoid'. Temps up to 18 degrees C.
9 36 1523 259 423839 5087058 Cloud vent (N4) with traps and worms R488-35
9 35 1523 271 423839 5087058 Mkr-N4. Swimming polynoid. Took several photos
9 38 1524 243 423844 5087042 Still probing for temperature trying to find a good place to fluid sample. Up to 18.4 degrees C.
9 43 1524 243 423486 5087525 Taking a fluid sample - bag #17 with no filter. Starting to sample at 09:4. Took Photo-29  



HFS-17_dfl

_R488-8

9 50 1524 243 424560 5087266 stopped filter sample
9 58 1524 242 423844 5087051 Scream mask at N4? R488-39
9 59 1524 243 423844 5087051 View of new scale worms. Lots o' flow and heat - 19.1 degrees C.
10 1 1524 243 423824 5087095 New species of worm at Mkr-N4 (Cloud vent) R488-41
10 1 1524 242 423824 5087095 GREAT close up of new scale worm!
10 2 1524 241 423907 5087074 Distant view of new species of worm at Mkr-N4 R488-42
10 5 1523 249 423907 5087073 Going to look for the 'hotter hole' in the vicinity of Mkr-N4. Photos 30 - 33
10 8 1524 242 423774 5087292 Lots of sparse tube worms. Jumbled lava. Photos 32 and 33.
10 11 1525 354 424006 5086794 Cloud abyss R488-43
10 10 1522 70 423795 5087166 At a different hot spot in Cloud area now, can't say exactly where. A big hole with lots of venting. Sparse tube worms. Stopping to take a temperature reading. Photo-34
10 15 1525 326 423846 5087052 The pit at Mkr-N6 R488-44
10 14 1525 327 423841 5087051 Temperature up to 20 degrees C. We are at Mkr-N6. Some kind of lava drainage pit. Photo-35
10 21 1526 332 423827 5087036 Started piston sample #2 at 10:10. Temperature up to 20.3 degrees C. HFS-2_df

l_R488-9

10 23 1526 329 424480 5087601 Navigation has moved position of Mkr-N4 from where it was last year to x=423905.9, y=5087109.6 - based on big Keith's piloting.
10 25 1526 334 423536 5086617 Stopping pumping into piston #2- 615 ml pumped . Starting gas piston #24 in same spot. HFS-24_dfl _R488-10
10 27 1526 330 423712 5087520 Stopped pumping #24 - 130 ml pumped. Starting to pump bottle #3 with a filter on HFS at 10:28. HFS-3_dfl

_R488-11

10 30 1526 332 423868 5087511 Tube worm at N6 R4845
10 34 1526 332 423714 5087518 Sample #3 stopped - 672 ml pumped.
10 35 1526 330 423871 5087062 Flushing the line for a GTB sample - planning to fire the port GTB.
10 38 1525 331 424718 5087025 Firing port side gas tight bottle. Temperature 20.3 degrees C. gtb_dfl_R488-12
10 42 1525 335 424691 5087066 Start pumping filter #7 on HFS - for sulfur. Still in same spot at Mkr-N6. HFS-7_dfl

_R488-13

10 48 1525 331 423956 5087073 Hightlight on
10 48 1525 328 423956 5087073 Scale worms R488-47
10 48 1525 334 423956 5087073 Worms at N6 (Cloud) R488-46
10 49 1525 333 423717 5087515 Scale worms: high density at N6: eranchintogluma (genus) R488-48
10 50 1525 329 423806 5087148 Stopping sample, 80ml volume.
10 51 1525 329 423845 5087046 Still photo of scale worms in high density and tube worms. Photo
10 51 1525 330 424689 5087060 scale worms at N6 R488-49
10 53 1525 331 423859 5087062 Completed fluid sampling, now going to maker N4 to do suction sampling.
10 55 1525 335 423843 5087047 No good nav. It's getting worse.
11 1 1521 92 423862 5086990 Heading back to N4 to do suction sampling.
11 7 1522 44 423453 5086878 Nav still bad.
11 7 1522 163 423453 5086878 Still photo Photo
11 10 1519 354 423453 5086878 Still photo Photo
11 11 1521 306 424409 5087694 Positioning at N4 to begin particle sampling.
11 13 1522 243 423910 5087074 Flushing the suction sampler bottle with honey bear.
11 18 1523 243 423845 5087051 Temp 18.2 from earlier reading. Positioning suction sampling probe.
11 19 1523 245 424586 5087587 Starting suction sample J3
11 21 1523 246 423470 5086849 Repositioning probe.
11 22 1523 244 423470 5086849 Position of suction probe. R488-50
11 23 1523 245 423470 5086849 suction sample SS-j3_dfl

_R488-14

11 29 1523 243 423887 5087133 Stopping J3 but will come back to it later on.
11 35 1523 263 425230 5086788 Begin suction sampling in 1 for mat material for Moyer. Sucked some rocks. SS-1_mat

_R488-15

11 42 1523 288 424695 5087420 Still sampling for mat but adding in a new worm or two as well. Sampler is reversed periodically to back flush rocks.
11 48 1523 285 424958 5087258 Still sampling for mat material.
11 52 1523 287 423885 5087127 Finished sampling on bottle 1. Mats and worms.
11 53 1523 285 423888 5087134 Returning to jar J3 to continue sampling for snails, mat, etc.
11 56 1523 247 423886 5087137 Repositioning to sample for J3 to sample for animals.
11 58 1523 246 423874 5087582 Probe arm dislodged a large rock and caused a debris storm.
11 59 1523 246 424953 5087259 Begin sampling scale worms into J3
12 4 1523 239 423894 5086984 Mkr-N4 R488-51
12 4 1523 239 423894 5086984 Trying to collect snails.
12 10 1523 235 423870 5087118 Still collecting animals in J3.
12 13 1523 234 423889 5087127 ROPOS is repositioning, will continue with J3.
12 16 1523 270 423890 5087126 Resume sampling J3.
12 19 1523 260 423889 5087126 Finished with J3. Going to the bear bottle (J4) for tube worm. Bottle 2 will be particulate.
12 21 1523 260 423888 5087126 Correction the bear is in bottle 4 not J4.
12 21 1523 261 423889 5087123 VCR tapes were changed at 1204
12 22 1523 260 423889 5087123 Honey Bear R488-52
12 22 1523 260 423891 5087127 Still photo. Photo
12 23 1523 268 424953 5087255 ROPOS is repositioning.
12 24 1523 261 423890 5087126 Begin worm sampling in flushing bottle. SS-4_bio

_R488-16

12 25 1523 278 423890 5087126 ROPOS is repositioning.
12 28 1522 281 423890 5087125 Continuing to sample worms.
12 30 1522 300 424334 5086908 ROPOS made a short reposition move. Sampling more worms.
12 35 1522 246 423891 5087125 ROPOS repositioned for last attempt at worms.
12 37 1522 266 423890 5087126 Sample completed. Will go back to N6 to do a particulate sample.
12 40 1522 218 423890 5087127 ROPOS is repositioning to N6. Highlights on.
12 44 1524 339 423888 5087136 Still photo back at N6 Photo
12 45 1524 349 423887 5087142 Highlights off.
12 47 1524 354 423904 5087113 Start sample J2 at Mkr-N6 for particulates. SS-j2_dfl

_488-17

12 56 1524 1 423892 5087121 We have a problem. The fluid sampler probe has been broken off. Suction sampler is stopped. Probe is broken of at support plate.
12 58 1524 4 423894 5087128 Broken probe for fluid sampler. R488-53
13 0 1524 1 423889 5087128 Testing the 7 function arm to see if it has been damaged. So far it looks ok. It may be possible to still sample with the probe. Probe itself still looks functional.
13 4 1524 1 423892 5087122 Highlights on.
13 6 1524 4 423894 5087128 Tube worms and large Pandorae at N6. R488-54
13 6 1524 6 423893 5087129 Finished sampling at N6 and will move back to the cage before moving north.
13 7 1522 24 423894 5087128 Still photo of worms at N6. Photo
13 9 1517 218 423889 5087128 Heading back to the cage.
13 10 1488 256 423889 5087128 Navigation is still not god.
13 11 1492 294 423884 5087043 Back at the cage.
13 13 1521 4 423853 5087096 ROPOS is on the bottom, heading north about 30m to Nascent vent. Hightlights on.
13 15 1521 4 423839 5087117 Still photo of bottom near Mkr-33. Photo
13 16 1521 360 423839 5087117 Photo of bottom near Mkr-33 photo
13 16 1522 3 423839 5087117 Orange stain on bottom. Photo
13 17 1521 4 423839 5087117 same as above photo
13 17 1520 4 424838 5086690 photo of dead tube worm group photo
13 18 1519 3 424838 5086690 No fixes
13 18 1517 6 425154 5087185 bottom Photo
13 19 1518 3 423834 5087138 Lava pillars. photo
13 21 1516 2 423846 5087134 Stopping highlights.
13 23 1519 4 424973 5087277 Photo of collapsed area with pillars. photo
13 24 1518 6 423653 5087813 Pillar Photo
13 24 1518 4 423841 5087120 bacterial mat on pillar photo
13 25 1518 7 423843 5087160 Edge of collapse area. photo
13 25 1517 7 423843 5087160 collapse pit photo
13 26 1518 2 423843 5087160 Highlights on and photo photo
13 27 1518 6 423836 5087175 collapse area Photo
13 29 1518 1 423840 5087203 photo of rattail. photo
13 31 1517 12 423840 5087214 Photo of lobates Photo
13 31 1517 32 423848 5087239 Rat tail R488-55
13 33 1518 33 423850 5087261 Photo of edge of collapse area. Maybe with crab. photo
13 34 1516 32 423851 5087259 Just changed the dive # on sit overlay from 487 to 488.
13 36 1517 29 423872 5087293 Signs of venting.
13 36 1516 34 423872 5087293 photo of area near Nascent vent. photo
13 37 1518 38 424187 5086782 Bottom showing signs of wear, hydrothermal activity. photo
13 38 1516 38 423901 5087309 Bottom. Getting some fixes. photo
13 39 1518 31 423877 5087352 Bacterial stain near edge of collapse area. photo
13 41 1517 3 423910 5087354 Bottom showing staining. photo
13 42 1518 359 423910 5087354 photo of live worms (but stressed?) photo
13 42 1518 6 423905 5087359 Bottom showing circular marks. photo
13 44 1518 357 423903 5087378 MTR line, shimmering water at Nascent photo
13 44 1518 340 423903 5087378 photo photo
13 45 1517 31 423904 5087388 Nascent photo
13 45 1517 31 423904 5087388 Nascent Vent R488-56
13 46 1518 68 423902 5087380 photo
13 46 1518 99 423902 5087380 photo
13 46 1518 107 423905 5087384 I think we just took 4 photos of Nascent. photo
13 46 1518 110 423904 5087387 position of MTR photo
13 46 1518 114 423904 5087387 Nascent Vent R488-57
13 47 1518 113 423905 5087389 photo
13 47 1518 113 423905 5087389 Positioning probe of fluid sampler (detached from arm.)
13 48 1518 113 423894 5087376 Checking water temp = 4C
13 49 1518 115 423910 5087372 Highlights off. Temp up to 10C
13 50 1518 114 424110 5086989 Up to 12C on temp.
13 52 1518 113 423894 5087376 HFS1 bag with a filter on it ready to start.
13 53 1518 115 423894 5087376 start HFS11 488-18 at Nascent vent next to MTR HFS-11_dfl

_488-18

13 55 1519 112 423894 5087492 Highlights of. Temp is 10-1C during sampling.
13 57 1518 115 423875 5087532 Highlights.
13 57 1518 113 423875 5087532 Nascent Vent near MTR R488-58
      Tube worms at Nascent. Scale worms. R488-59
14 0 1519 115 423907 5087395 Sample HFS1 is complete.
14 0 1519 110 423908 5087396 Max T during sampling was 15.5C.
14 4 1519 112 423906 5087381 Position of MTR with critters at Nascent. R488-60
14 6 1516 355 424900 5087534 Photo of bottom before moving photo
14 6 1519 349 425126 5087098 same photo
14 7 1518 356 423904 5087392 photo of bottom photo
14 7 1517 356 423904 5087392 photo photo
14 8 1518 357 423990 5086180 photo
14 5 1519 114 423905 5087383 R488-61
14 8 1519 5 423990 5086180 bags on rope fish photo
14 8 1518 4 423990 5086180 Nascent Vent R488-62
14 9 1519 7 423905 5087407 VCR tapes changed.
14 9 1519 9 423905 5087409 Rat tail at Nascent vent R488-63
14 9 1519 10 423904 5087412 feeding rat tail w/ worms MTR at Mkr-N41 photo
14 10 1519 13 423908 5087409 Highlights off and photo photo
14 13 1517 3 423905 5087413 Getting ready to move to pit. Fe staining photo
14 13 1518 2 424683 5087400 Photo of Fe staining photo
14 14 1519 20 423893 5087418 Highlights on.
14 16 1517 7 423895 5087423 Heading to pit vent.
14 16 1516 358 423895 5087423 Highlights off.
14 21 1518 344 423894 5087427 photo of 1998 lava flow during traverse between Nascent and The Pit. Lobate uncollapsed part of flow. Few vents now, but lots of tan iron staining. 2 photos
14 24 1517 346 423628 5087169 lobate flow. small pockets of white mats between lobes. sparse crabs 3 photos
14 26 1519 345 423880 5087477 lobate roof. some pillows. few vent pockets photo
14 27 1519 345 423872 5087481 lobate lavas - mostly black now (little staining or venting). some collapsed lobes photo
14 29 1518 340 424951 5087500 lobate lava. sea star. ophuriods. 2 photo
14 30 1518 342 424394 5087609 seemed like we went over a contact into newer-looking glassy lava. maybe a kipuka of older lava? photo
14 31 1518 341 423849 5087537 remnants of brown mat on lava
14 33 1519 337 423851 5087538 2 calanoid copapods on clump of flock R488-64
14 34 1518 345 423852 5087552 black shiny lobate lava. minor brown staining.
14 35 1518 348 423845 5087569 photo photo
14 36 1518 347 423823 5087607 Good fix
14 37 1518 344 423823 5087607 photo of collapse area photo
14 37 1518 341 423823 5087607 photo bottom with crab photo
14 39 1518 345 423816 5087608 collapse area photo
14 40 1518 343 423816 5087608 collapse area photo
14 40 1520 344 423816 5087608 collapse area photo
14 41 1517 340 423786 5087698 good fix
14 44 1518 349 423746 5087736 Arrived at pit
14 44 1518 344 423746 5087736 photo of pillars at Pit photo
14 46 1519 345 423737 5087754 Highlights on.
14 47 1518 348 423742 5087774 Lava pillar. photo
14 54 1518 334 424881 5087801 At Mkr-N1 - 'The Pit'. Lobate flows. Lots of orange floc in water and in cracks. Sunken collapsed pit with white mat. Looks like venting has reduced from last year. photo
14 56 1519 287 423707 5087831 Putting the temperature probe in the pit to test for anomalies.
15 1 1517 302 424678 5087832 Only a 0.2 degrees C anomaly. Will not fluid sample here.
15 1 1518 356 424678 5087832 Collapsed area. photo
15 2 1518 330 424678 5087832 Heading north from Pit towards Roof vent.
15 3 1518 360 423692 5087831 Lobate flow with orange iron oxides. photo of crab.  



photo
15 5 1519 2 423680 5087855 Rattail.
15 5 1519 3 423680 5087855 Rat tail fish observed during move from N1 to Roof R488-65
15 13 1519 346 424755 5087872 Lava pillars. photo
15 14 1518 357 423693 5087877 Uncollapsed roof on lobate flow held up by pillars observed on route to the Roof location. R488-66
15 14 1519 356 424755 5087872 Collapsed area - pillars and caverns. Old lava flow supported by pillars.
15 17 1518 2 423703 5087899 Zoarcid fish swimming away from us. Stopped and is sinking in the water column.
15 19 1518 358 423690 5087931 Still heading north towards Roof - about 20 m to go. Another fish in view.
15 20 1518 350 423682 5087933 Arched layered pillar structure.  

Photo
15 20 1518 356 423690 5087942 Pillars in lobate flows observed on route to Roof"" R488-68
15 22 1519 355 423687 5087952 Pillar with hole at top in lobate flow. R488-69
15 22 1519 353 423687 5087952 Small pillar with a sunken depression on top - hollow inside. Photo
15 23 1518 359 423690 5087969 Two crabs. Photo
15 26 1518 358 423683 5088011 Photo of bacterial mat at base of a cavern. Photo
15 28 1518 357 423683 5088054 Lobate lava. Photo
15 30 1520 6 424731 5086196 Ugly rattail.
15 31 1519 357 422818 5083961 Lobate flow with a little bacteria. Photo
15 32 1519 1 423634 5088073 Highlights of oxide mat/deposits on lobate lavas. Photo
15 34 1520 1 423691 5088124 Photo of rattail in a small pit. Photo
15 35 1521 356 423695 5088134 At Mkr-N9 (MiniSnow). No profuse white floc like last year, but just a bit of white mat with mostly orange iron deposits. NO ANIMALS.
15 38 1522 356 423695 5088134 Probing for temperature at mkr N9. Found an anomaly - up to 3.4 of a degree C.
15 42 1522 354 423408 5087354 Start to sample bottle #18 of HFS at mkr N9. A scale worm and a few limpets here. HFS-18_dfl _R488-19
15 48 1522 351 423755 5087941 Close ups of orange stuff - looks like mat? sediment? Looking around while doing the fluid sampling.
15 51 1522 356 423701 5088129 Stopping sample - pumped 720 ml. Highest temperature was 4 degrees C.
15 54 1521 318 423714 5088135 Moving north again towards marker N3. Photo
15 55 1521 348 423689 5088161 Dead bush of tube worms - dead and covered with the orange mat/sediment stuff. Photo
15 55 1522 342 423689 5088161 Dead tube worm bush covered in orange sediment and surrounded by white mat observed on route to N3 R488-72
15 57 1522 338 424602 5088298 Lots of orange and white mats in between lobates.
15 59 1522 338 425156 5087802 Bright red sediment/mat, much darker red then the typical orange mat we've been seeing. Photo

Photo

16 0 1522 345 425156 5087802 Red material found on bottom on route to N3 R488-73
16 2 1523 328 425156 5087802 Collapsed pit. Photo
16 4 1525 336 423672 5088214 Folded sheet flow lava.
16 5 1526 341 423672 5088214 Jumbled lava, we are probably down in the collapsed area.
16 8 1526 336 423672 5088214 Just came up past talus over a wall, out of the collapsed pit. Photo
16 15 1528 334 423632 5088288 Palm worm and scale worm in vent near N3 R488-74
16 9 1526 335 423672 5088214 Lobate lava flows with lots of bacterial mat in cracks. Lots of snails, a palm worm. Going in to probe for temperature. Good fix for this location: 3635,8290. T max at 16 degrees C. Near Mkr-N3. Starting to suck piston #20 of HFS at 16:16. 2 Photos

HFS-20_dfl _R488-20

16 22 1528 333 423633 5088284 Stopped pumping #20 - aprox. 650 ml.
16 26 1528 339 423632 5088286 Starting to pump filter #2 of HFS for XRF. HFS-2_dfl

_R488-21

16 31 1528 334 423632 5088285 The poor bear. he earned this trip down. he sleeps with the worms and stuff. R488-75
16 33 1528 331 423626 5088302 Stopped pumping into #2. Max temp was 17.3 degrees C and 80 ml was pumped.
16 38 1528 336 423626 5088292 starting new sample. gas piston. HFS-5_dfl

_R488-22

16 40 1528 337 423626 5088292 photo at sampling site. pulling sampler out of vent. near marker-N3 photo
16 43 1526 28 423627 5088293 We're on the MOVE! North to Milky vent!
16 46 1528 30 423652 5088318 More dark red sediment/mat interspersed with regular orange mat.
16 50 1526 31 423676 5088347 Lineated sheet flow. Not striated. Why? because it's lineated (Bill Chadwick quote). Photo
16 56 1528 3 423700 5088402 We were in ropy to jumbled sheet flow, but now in a lineated patch. So probably in a collapse. Photo
17 16 1529 313 423672 5088425 Found the Osmosampler at Milky vent!
17 19 1532 57 423664 5088421 At the OSMO - Mkr N2 is at Milky Vent. There is also a bacteria trap here.
17 21 1532 71 423666 5088428 Osmo sampler at N2-Milky R488-76
17 22 1532 54 423671 5088429 Osmo sampler at Milky R488-77
17 23 1532 57 423671 5088429 Osmo sampler intake at Milky w/Bacteria trap next to it. R488-79
17 25 1532 58 423664 5088427 Flushing the HFS and taking temperature measurements
17 35 1532 61 423665 5088437 Starting to sample bag #19 of HFS at 17:36. HFS-19_dfl _R488-23
17 43 1532 64 422984 5088429 Stopped sampling #19, got 70 ml, max. temp. 3.4 degrees C.
17 44 1532 64 422984 5088429 Main gauge shows the low hydraulic pressure for the ROPOS system R488-80
17 46 1532 64 423665 5088428 Heading to Easy Vent. About 20 m north of Milky.
17 52 1532 62 422976 5088456 Before we leave for Easy we are trying to pick up the osmosampler from Milky.
17 53 1527 27 423666 5088426 We are leaving Milky (Mkr N2) with the osmosampler in the port arm. Trying to get a better grip on the osmo.
17 59 1529 337 424763 5088447 Have the osmo. Heading to Easy.
18 1 1531 0 423670 5088428 Crab near Milky R488-81
18 1 1531 359 423670 5088428 Crab in view. Very strange lava morphology.
18 5 1531 293 423671 5088450 Area near Easy Vent. Photo
18 6 1532 306 423661 5088451 Little oxide mounds. Photo
18 7 1527 201 425416 5087550 Tubular lobate lavas.
18 9 1528 222 423676 5088455 Good ROPOS fix. Should be a couple meters to the NW.
18 10 1529 225 423033 5088302 Going to go north laterally along the Milky-Magnesia line. Couldn't find Easy, so we're going to check out if there is other venting along the line.
18 12 1527 3 423033 5088302 Out of the collapsed area into lobate flows.
18 16 1526 343 423678 5088468 Photo
18 16 1526 349 423678 5088468 Only a little white-ish mat along the line of venting from last year. It seems like the venting has subsided significantly in the last year. 2 Photos
18 20 1527 8 423668 5088495 Nearing Ouzo Vent. Some residual bacterial mat in a deep crack. Photo
18 29 1533 350 423689 5088525 Going in with HFS probe to sample temp at Ouzo? vent - max temp is 3.6 degrees C. Difficulty getting the probe all the way into the crack. Fix on this location is x=3668, y=8518.
18 33 1533 356 423689 5088525 Taking temp between lavas in crack. Highest T is not in the deepest part of hole. Temps in mid 3 degrees, higher up in hole. Fluid sample 16 at 18:37, bag without filter HFS-16_dfl _R488-24
18 40 1534 353 423676 5088519 End fluid sampling, highest T 3.8, but 3.5 when sampling
18 42 1530 184 423668 5088515 Going back to cage for tether management.
18 48 1529 48 423656 5088504 Pillar connected to roof or arch near the stern of ship/cage/Ouzo vent. R488-83
18 48 1529 48 423656 5088504 Down by cage, collapse area with nice pillars supporting collapsed roof. Photo of roof Photo
18 50 1530 85 423645 5088519 Traversing back to Ouzo Vent after tether management.
18 51 1528 92 423701 5088502 Tree pillars in sequence near Ouzo vent. R488-84
18 52 1528 112 423701 5088502 pilars photo
19 6 1521 242 423641 5088484 Good fix.
19 14 1528 90 423686 5088500 Photo taken in error photo
19 15 1528 266 423696 5088494 hmm crappy timing R488-85
19 15 1528 241 423696 5088494 Good fix.
19 20 1529 95 423683 5088508 pillar R488-86
19 24 1532 321 423671 5088516 paw print R488-87
19 23 1530 311 423680 5088510 Arrived at Ouzo, finally. Will try to sample water.
19 35 1527 295 423658 5088515 Repeating . Looking for place that sampling started at 18:40 was interrupted by need for tether management.
19 41 1528 58 423655 5088543 At a previously undiscovered small vent about 10 m south of Magnesia Vent. Named Whiteout Vent
19 45 1530 60 423656 5088554 whiteout vent R488-88
19 46 1530 62 423658 5088545 more blind temp probing R488-89
19 46 1530 53 423658 5088546 T max = 4.6 C R488-90
19 47 1530 53 423658 5088546 HFS sampler piston #23 getting ready. X = 3365.9 Y = 8526 with offset.
19 52 1531 344 423658 5088547 Start sampling piston #23. HFS-23_dfl _R488-25
19 57 1531 343 424852 5085576 Sampling #23 piston completed. 60 ml. Max. T 5.6 C.
20 0 1531 344 425382 5088120 Start DNA filter set #12 at same place as #23 at Whiteout Vent at 20:02.
  20 5 1531 353 423651 5088552 Highlights tape on. Panning around Whiteout Vent while sampling. HFS-12_dfl _R488-26
20 11 1531 354 423555 5088719 Stopped sampling at 20:1. 1020 ml. T max. 5.1 C.

Next sample is bag #9 with filter.

20 12 1531 356 423656 5088550 Start sample #9. Bag has a POC (particulate organic mater) filter HFS-9_dfl _R488-27
20 18 1532 356 423449 5088873 Stopped sampling at 20.18 681 ml. Max. T 4.3 C.

Start HFS sampler filter #1 for FISH (fluorescent in situ hybridization). Photo

HFS-1_dfl

_R488-28

20 22 1531 355 423654 5088548 shelf with hotter fluid on other side (video confused) R488-91
20 24 1531 358 423594 5088576 another shot of the shelf or plateau R488-92
20 25 1531 359 423553 5088710 lava bridge covered in snow (top previously referred to as "shelf" R488-93
20 27 1531 1 423656 5088558 polynoid close-up R488-94
20 28 1531 358 423678 5088562 oozing polynoid (maybe it doesn't like the company) R488-95
20 28 1531 358 423678 5088562 Stopped sampling. 105 ml. Max. T 5.1 C.Highlights on.
20 30 1531 357 423678 5088562 Start HFS sample #10 with filter for lipids. HFS-10_dfl _R488-29
20 35 1531 2 429242 5083886 Stopped HFS sampler.
20 41 1531 358 423649 5088552 Start HFS filter #21 at 20:40 HFS-21_dfl _R488-30
20 41 1531 358 424152 5088238 I now have light that doesn't interfere with anyone and I can finally see the keyboard. (Steve)
20 53 1531 358 424600 5087761 smile button in a more reflective mood R488-96
20 53 1531 358 423648 5088554 snow blowing R488-97
20 53 1531 357 423648 5088554 snow blowing II R488-98
20 54 1531 357 423651 5088558 Stopped sampling. 1542 ml. T max 5.0 C Next will be HFS piston #4
20 56 1531 356 424673 5088570 Started HFS piston sample at 20:56. Still at same place at Whiteout Vent. HFS-4_dfl

_R488-31

20 56 1531 358 423673 5088557 lava bridge again (higher flow on other side?) R488-99
21 2 1531 359 423651 5088553 Stopped sampler. 650 ml. T max 4.2 C Next (last) is HFS filter set #13.
21 7 1531 1 423659 5088555 Started sampling DNA filter set #13 at 21:04. HFS-13_dfl _R488-32
21 11 1531 0 423622 5088584 white out R488-100
21 12 1531 1 423588 5088578 plume at white out R488-101
21 14 1531 360 423653 5088558 plume at white out R488-102
21 17 1531 359 423653 5088558 plume at white out R488-103
21 17 1531 359 423653 5088558 Getting suction sample jar J2 ready at Whiteout Vent.
21 20 1531 358 423650 5088554 Stopped DNA filter sample. End of HFS sampling - al l containers are full.
21 26 1532 14 423650 5088554 Start suction sample into jar J2. Intake is dangling (5 function is holding a recovered osmosampler) over the snowblower at Whiteout Vent at same place where all HFS samples were taken. SS-j2_dfl

_R488-3

      Will be a slow process. Considered opinion is that the floc is mucus - bound sulfur produced by bacteria.
21 39 1532 10 423654 5088555 you WORM! R488-104
22 0 1532 15 423710 5088526 End suction sample in jar J2 at Whiteout Vent.
22 5 1531 84 423648 5088552 Leaving Whiteout vent R488-105
22 5 1531 82 423648 5088552 Highlights on. Whiteout Vent.
22 7 1531 72 423648 5088552 Whiteout vent R488-106
22 7 1531 73 423648 5088552 Photo of Whiteout Vent. Counter says #174 or 177. Monitor says #198. Take your pick, or none of these.
22 11 1522 59 423652 5088540 Highlight video stopped.
22 16 1478 274 420625 5088006 All video recording stopped. Getting ready for (yawn) Imagenex single line back to elevator.
22 28 4397 268 423568 5088479 Oil reservoir has only 4% left on sub display.
22 47 1510 322 423418 5088499 Starting Imagenex line W4 running from due west of Whiteout Vent proceeding south along west edge of '98 lava to elevator as a fill-in line for '98 NeMO survey. Auto altimeter not working.
22 57 1508 178 423568 5088168
 23 2 1509 105 423365 5088471 These are good navigation numbers. Apparently we are behind the cage.
23 53 1499 178 423429 5087592 Approximately 450 m to go.
0 19 1496 182 423430 5087202 10 meters away from the corner
0 32 1497 128 423417 5087126 Turning the corner of the line and heading to the elevator. Lost oil pressure on the hydraulics.
0 39 1397 192 423412 5087128 Lost hydraulic pressure right where we were leaving imagenex line w4 to go over to the elevator. ropos back in cage and heading back to the surface
0 47 1197 215 423412 5087128 On our way up.
       Dive Summary: We started at Mkr-33 and took filter samples,

fluid samples, slurp samples, and one Ti gas-tight sampler. We then moved to Snail and Marcus took a slurp sample.

       We also took temperature measurements at Snail with the HFS and got a Tmax = 13. We then moved to Cloud, Mkr-N4 and Mkr-N6 where Ti gas-tight, gas piston, regular piston, bag samples and filtered slurp samples were taken.
      Bag samples were then taken at Nascent and MiniSnow Vents. At Mkr-N3 bag, filter, piston and gas piston samples were taken. Bag samples were then taken at Milky and Ouzo Vents.
      We then came upon Whiteout Vent, a previously unnoticed area and took a gas piston, filters for DNA, POC, sulfur and lipids, piston and bag samples. The remainder of the dive was an Imagenex survey, until oil pressure was lost and the dive aborted.
      Notice: Due to the nav offset, as not all transponders were working, it was later determined that "Whiteout Vent" is actually Magnesia Vent discovered in '98. Much more active in '99.



    Dive R489
      Dive Plan: Suction sampling - Cloud for Moyer, Mkr-33 for Levesque, Mkr-N41 for Scott/Juniper, Nascent for Levesque/Marcus/Tunnicliffe, Whiteout for Moyer/Juniper, Oxide samples for Scott/Juniper (in flushing bottle).
      BioBox collections - Mkr-N41, Nascent, Old Flow, OldWorms, and for traps if previous collections are successful. Imagenex survey of western part of lava flow. Recovery/deployments of MTR/Vemcos, osmos and traps. Collect basalt samples, recover elevator.
      ROPOS Configuration: Suction sampler, Biobox, Gastights, Imagenex, cameras, MTR's carried down in Biobox, BioOsmosampler carried down in 7-function for deployment
      Note: Nav fixes not edited in the logs. All fixes were logged, regardless of accuracy. Use the dive plots to determine actual positions.
  UTC Hour Min Depth (m) Sub Hdng (deg) UTM X UTM Y Comments Frmgrbs, Photos, Samples
3 20 1 310 423413 5087134 ROPOS in water
3 20 1 310 423413 5087134 ROPOS back on deck because have to remove some weight. Only got down to 80 meters, so will continue this dive as R489
7 25 41 316 423412 5087128 Sub in water. Yahoo!
7 36 96 214 423412 5087128 Sub coming back up, didn't get to bottom. Dive aborted. Lost pressure in main reservoir.
7 50 0 43 423412 5087128 on deck
7 56 1 112 423412 5087128 Sub on deck.


   Dive R490
      Dive Plan: Continuation of previously aborted attempts. Primary objective is to complete basic biological collection at Mkr-33/Cloud Vents and to the north (Nascent and old vents )and Whiteout Vent.
       Osmos and traps to be recovered and put in the elevator. Recover elevator (weather permitting) during the dive.
      ROPOS Configuration: Suction sampler, Biobox, Gastights, Imagenex, cameras, MTR's carried down in Biobox, BioOsmosampler carried down in 7-function for deployment
        Note: Nav fixes not edited in the logs. All fixes were logged, regardless of accuracy. Use the dive plots to determine actual positions.
  UTC Hour Min Depth (m) Sub Hdng (deg) UTM X UTM Y Comments Frmgrbs, Photos, Samples
19 42 1 97 423412 5087128 19:45 ROPOS in water! Everything normal. This COULD be the one!
  19 51 56 78 423412 5087128 ROPOS came out of the cage at 50 m.
  20 7 259 242 423412 5087128 ROPOS back in the cage
  20 16 411 13 423906 5086990 ROPOS coming back on deck because electrical compensation gauge has gone to zero. This is a component that had been disassembled during stipdown for hydraulic system repair.
  20 19 394 21 424024 5087248 20:32 ROPOS back on deck.
      Dive Summary: El Guapo was still plugged in after calibration,


    Dive R491
      Dive Plan: Elevator will be used to gather experiments from Mkr-33, weather permitting. Suction sampling will be done there and at other active vents to the north. Biology and rocks will also be sampled Imagenex survey along western edge of '98 lava flow.
      ROPOS Configuration: Biobox, Slurp Sampler, Gas Tights, Imagenex, and the other cameras.
      Note: Nav fixes not edited in the logs. All fixes were logged, regardless of accuracy. Use the dive plots to determine actual positions.
  UTC Hour Min Depth (m) Sub Hdng (deg) UTM X UTM Y Comments Frmgrbs, Photos, Samples
2 0 368 162 424016 5087158 ROPOS in water at approximately 01:45 UTC. ROPOS now at 389 meterss
  2 5 466 155 423896 5086982 ROPOS descending in cage, systems look ok.
  2 20 756 287 423966 5087087 Having problems and making adjustments with video cables. Color back on main console monitor.
  2 26 878 207 423960 5087078 Oil reservoir started at 10%, now at 83% and steady while descending in cage.
  2 47 1268 35 423903 5087058 Synchronizing computer clocks to UTC
  2 53 1371 29 423949 5086993 Oil reservoir at 79%
  3 3 1468 154 423981 5086921 Parking sub about 50 m off bottom to wait for P-Code GPS to come back on.
  3 47 1492 281 423842 5087094 Pcode came back, ship has been maneuvering back into position. Ropos out of cage and going down. ON bottom. sheet flows with flocky coating. no fix now
  3 53 1524 26 423868 5087084 x=3869, y=7084 good fix. Found elevator
  3 55 1523 268 424393 5086607 heading to 75 degree spot. Looking for it.
  3 59 1524 225 424409 5087488 At Mkr-33 site. Fissure in sheets has much shimmering water. Putting out OSMO samplers. Discussing where first.
  4 1 1523 228 424230 5087962 Putting Osmo sampler in vent site Mkr-33. R491-1
4 2 1523 224 423841 5087087 note that frame grabs appear to have subtle red stripes across them but they are not on the video
  4 4 1523 228 424108 5087609 Positioning OSMO R491-2
4 20 1522 236 423854 5087066 BioOsmo sampler deployed in hottest known part of Mkr-33 vent
  4 22 1524 215 424495 5087976 Osmos and Micro traps R491-3
4 23 1524 236 423610 5086988 Osmo and micro traps R491-4
4 29 1524 244 423839 5087050 Repositioning the green copper OSMO sampler at the 80 degree vent next to the other sampler just positioned
  4 30 1522 235 423851 5087073 Moving Osmo sampler (green copper one) to 80 C spot in crack R491-5
4 31 1523 240 423853 5087101 repositioning green osmo next to the BioOsmo R491-6
4 33 1523 258 423842 5087050 Osmo samplers at fissure Mkr-33 R491-7
4 33 1523 258 423842 5087050 3 Osmosamplers at Mkr-33 R491-8
4 33 1522 226 423856 5087066 another picture of osmo set up R491-9
4 34 1521 227 423875 5087038   R491-11
4 34 1522 220 423856 5087066 3 Osmosamplers at Mkr-33- R491-10
4 34 1521 238 423875 5087038   R491-12
4 35 1521 209 423851 5087097 Mkr-33 view R491-13
4 35 1521 209 423851 5087097 More Osmosamplers at Mkr-33- R491-14
4 37 1524 214 423848 5087102 Osmo analyzer R491-16
4 37 1524 212 423848 5087102 Iron osmo sampler being picked up. R491-15
4 38 1524 213 423849 5087104 Iron baby! Osmo analyzer R491-17
4 38 1523 206 423852 5087092 Picking up Fe osmo sampler to bring back up to surface.
  4 39 1524 192 423840 5087051 Traps and osmoanalyzer R491-18
4 40 1524 201 423838 5087048 Osmoanalyzer R491-19
4 43 1523 211 424308 5087903 Picked up Fe OsmoAnalyzer #1 for return to elevator (which is only 2 m away) FeOsmoAn_ R491-1
4 46 1520 211 424481 5087352 Osmo sampler into elevator. R491-20
4 48 1521 210 423863 5087061 Fe osmo sampler placed in elevator.
  4 54 1524 249 424043 5087941 Picked up bacteria trap #40 and took to elevator. Bactrp-40_ R491-2
4 55 1524 251 423851 5087099 Stopped on way to elevator to pick up bacteria trap #43 and took both to elevator. Bactrp-43_ R491-3
4 57 1522 265 423853 5087092 Bacterial traps 40 and 43 put in elevator.
  5 13 1521 340 423860 5087097 Removing OSMO sampler from elevator.
  5 16 1521 173 423864 5087074 Last entry was wrong. Just repositioning OSMO to make more room. Now getting other OSMO to put in elevator
  5 19 1521 166 424412 5087480 Recovery of iron osmotic analyzer 2, putting in elevator FeOsmoAn_ R491-4
5 29 1525 61 423840 5087041 repositioned bacterial traps 42 R491-21
5 31 1523 38 423840 5087046 repositioned #42 bacterial traps
  5 37 1522 42 424674 5087294 Recovered bacterial traps #41.
  5 39 1524 214 423840 5087045 Repositioned two bacterial traps 41 and 42 in crack where the temp. is hot and where the OSMO samplers were deployed.
  5 42 1524 223 423993 5087682   R491-22
5 42 1524 226 424362 5087686 bac traps 42 and 41 in place R491-23
5 46 1524 202 423992 5087679 osmos and bac traps in bed together R491-24
5 50 1524 229 423849 5087098 Position of bacterial traps R491-25
5 50 1523 223 423849 5087098 OK NOW IT'S THE FINAL POSITION OF BACTERIAL TRAPS NEXT TO OSMOS R491-26
5 50 1522 221 423849 5087098 The big view of Mkr-33 R491-27
5 52 1522 180 423852 5087102 Another great grab of Mkr-33 site R491-28
5 53 1524 188 423857 5087092 Going to suck snails. We will take sample 5. Photo
5 56 1524 183 423855 5087090 Limpets and snails close-up. R491-29
6 2 1524 192 423859 5087097 suction sample J1 starting R491-30
6 2 1524 188 423744 5087252 Sampling limpets and snails in J1 SS-j1_bio

_R491-55

6 7 1524 188 423839 5087040 3852 7103 good fix at Sample site.
  6 19 1524 189 423841 5087042 finished suction of snails.
  6 21 1525 183 423853 5087105 Heading to N6 at Cloud, 50 m away.
  6 30 1514 62 423861 5087050 moving the ship.
  6 31 1521 64 424380 5087671 Traversing to NE to next site over lineated sheet flows, some jumbled sheets.
  6 35 1524 137 423898 5087114 Over the plume, lots of cloudy water. 3901 7112 good fix. over the vent with worms on edges, many scale worms on rock faces. Continuing to move around. Much collapse, some pillars.
  6 42 1522 305 423910 5087105 Found Mkr-N6 site
  6 42 1523 285 423903 5087112   R491-31
6 42 1522 334 423933 5087605 View of Cloud Vent, multiple orifices, Mkr N6 R491-32
6 42 1522 295 423903 5087112 3903 7112 good fix at site N6, vigorous venting in hole, many scale worms. some nice tubeworms.
  6 43 1526 352 423773 5087274 Edge of N6, tube worms, venting vigorous R491-33
6 45 1526 345 423887 5087122 Going to deploy Marker 21 near N6. Marker will have MTR temp recorder on it. Big fish, much bacterial coating on basalts. x=3901 y=7109 good fix.
  6 51 1526 357 423903 5087110 MTR probe at MKR N6 R491-34
6 54 1526 355 424287 5087490 MTR deployed at Mkr N6 Cloud. Mkr-21 being deployed R491-35
6 52 1526 0 423901 5087108 3903 710 good fix where Marker 21 is being positioned down the hole. Gas tight taken in whole. very vigorous flow. gtb-6_dfl_R491-6
7 1 1527 0 423902 5087110 Trying to se exactly how big this hole is by scanning it with the red laser things - approximately 5 cm in width.
  7 4 1527 356 424139 5086948 The 5 cm was at the north side of the hole. Photo-1
7 4 1527 358 424139 5086948 Photo of Cloud Marker N6/21.
  7 4 1527 0 423754 5087341 The pit at Cloud vent - marker N6 and 21 R491-36
7 6 1527 358 423890 5087118 Watching the particles go by to estimate a flow rate out of the pit at Cloud Marker N6/21.
  7 8 1527 360 423902 5087109 Looking at the face of the rock right above the rim to get this flux measurement.
  7 14 1527 312 423892 5087118 We are suctioning at N6 for bacterial mat. SS-1_mat_

R491-7

7 14 1527 310 423901 5087108 Photo of the hole and bacterial mat and worms. Photo-2
7 19 1527 8 423903 5087110 x=3903 y=7110 good ROPOS fix
  7 20 1527 6 423905 5087111 Suction sample of microbial mat at Cloud vent (N6) R491-37
7 21 1527 7 423902 5087096 We're actually sucking mat near Mkr N6 and the Pit, but not actually in it. The good fix is for where we're sampling from now.
  7 29 1526 19 423901 5087107 Still suctioning into Bottle 1.
  7 32 1527 22 423900 5087107 good suction sample at Cloud vent of bacterial mat. R491-39
7 34 1527 22 423903 5087109 Flushing. Finished sampling into Bottle 1.
  7 34 1527 19 423903 5087109 We're moving to the other side of Cloud now.
  7 37 1522 9 423898 5087106 Retrieving bacterial traps from Marker N4.
  7 38 1524 9 423894 5087110 3894 710 good fix at Marker N4
  7 38 1525 6 423894 5087110 Bacteria trap at N4 R491-41
7 40 1525 234 423898 5087112 Trying to pick up bacterial trap from N4.
  7 42 1525 233 423897 5087111 Retrieval of bacteria traps at N4 R491-42
7 42 1525 230 423897 5087111 Getting bacterial trap #36 bactrp-36_

R491-8

7 44 1525 229 423897 5087111 Putting bacterial trap #36 in starboard side of BioBox.
  7 46 1525 236 423898 5087111 Now we're retrieving bacterial trap #37 from Mkr N4.
  7 48 1525 221 423897 5087111 Putting bacterial trap #37 in starboard side of BioBox bactrp-37_

R491-9

7 51 1525 244 423896 5087111 Putting bacterial trap #39 into starboard BioBox. bactrp-39_

R491-10

7 54 1526 237 423900 5087115 Retrieving and placing bacterial trap #38 into starboard BioBox. bactrp-38_

R491-11

7 57 1526 240 423897 5087111 Going to the elevator!
  7 58 1526 238 423896 5087110 3897 7110 good fix
  7 58 1526 238 423896 5087110 3897 7110 good fix
  7 58 1526 240 423896 5087110 Going to the cage for tether management first.
  8 7 1520 226 423866 5087085 We found the cage.
  8 8 1522 239 423858 5087076 x=3863 y=7099 good fix on elevator
  8 10 1518 179 424092 5087574 We're now at the elevator to put the bacterial traps into the BioBox in the elevator and take the other bioboxes out of the elevator.
  8 15 1521 47 423860 5087081 Bacterial trap missed the box in the elevator. Still trying.
  8 28 1522 153 423862 5087095 3862 7095 good fix at elevator as we put the trap in the right box in the elevator this time.
  8 29 1521 181 423859 5087081 Transferring bacteria trap to elevator R491-43
8 29 1521 181 423859 5087081 The second bacterial trap is now entering the elevator.
  8 32 1521 171 423860 5087079 The third bacterial trap is now going in the elevator.
  8 33 1521 176 423858 5087095 3858 7095 good fix at elevator.
  8 35 1521 183 423863 5087099 Placing the fourth and final bacterial trap into the elevator.
  8 35 1521 160 423860 5087077 We're having trouble shutting the lid on the box in the elevator with the bacterial traps in it.
  8 45 1522 3 423857 5087073 We have closed the lid on the box in the elevator.
  8 45 1519 101 423858 5087074 Now we will try to remove the portable bioboxes from the elevator. Two will go on seafloor, two will be held in each of the arms.
  8 58 1518 140 423856 5087098 3856 7098 good fix
  9 3 1525 37 423854 5087081 Picking up two of the portable bioboxes - one in each arm - and carrying them to Nascent Vent.
  9 9 1526 64 423848 5086872 The two boxes left behind are close to the anchor of the elevator.
  9 10 1526 59 423852 5087073 Now we're going back to the cage to check the tether and such.
  9 17 1494 174 423935 5087512 We've found the cage, everything's okay, so we're returning to the bottom to head to Nascent.
  9 18 1522 0 423641 5087320 Moving the ship to get ready to move to Nascent.
  9 20 1522 6 423641 5087320 Going north to Nascent vent.
  9 26 1519 3 423873 5087139 On transit to Nascent, seeing lots of diffuse venting with bacterial mat. No sign of animal life.
  9 27 1518 6 423873 5087146 Now we see some tube worms.
  9 27 1520 4 423873 5087153 Very pretty diffuse flow. A little pit with mat and worms and snails.
  9 28 1521 1 423871 5087156 3871 7156 good fix We're still heading north. Saw the new worm, extensive venting.
  9 29 1520 4 423869 5087164 Now we're seeing pillow lavas with that rusty colored sediment/bacterial stuff on it.
  9 30 1521 2 423858 5087179 More rusty colored stuff on the pillow lavas. Some bacterial mat.
  9 31 1520 5 423863 5087180 3864 7180. good fix en route to Nascent.
  9 32 1520 356 423864 5087191 Funny pitted looking lavas - poxey area. They look like grazing spots except that they're single little circles with no animals or fecal trails.
  9 34 1521 4 423870 5087217 Trying to figure out what these little patches are - fecal patches?
  9 37 1518 356 423867 5087262 3867 7263 good fix en route to Nascent - 140 m away from Nascent.
  9 41 1520 10 423886 5087318 Pillow lavas with orangey stuff- no venting or animals visible.
  9 42 1520 7 423886 5087337 3886 3737 About 60 m from Nascent Vent.
  9 44 1520 73 423882 5087386 More poxey sort of areas on the pillow lavas
  9 44 1521 82 423909 5087757 Rattail.
  9 44 1521 78 423909 5087757 Rattail on the way to Nascent. R491-45
9 45 1520 99 424259 5087166 We are very close, or almost at, Nascent Vent. Looking for the MTR.
  9 46 1520 97 424264 5087173 Bag creatures, snow-looking stuff.
  9 48 1518 164 423908 5087390 3908 7391
  9 49 1520 173 423913 5087390 We have located the MTR.
  9 49 1520 171 423910 5087380 3910 7380 good fix at MTR at Nascent Vent.
  9 50 1521 119 423909 5087376 The bag creatures are super thick! We see snails, limpets, tube worms, etc.
  9 51 1521 127 423911 5087380 Nascent Vent with old MTR in it. R491-46
9 52 1521 122 423911 5087380 Putting down the two portable bioboxes to free up ROPOS's arms.
  9 53 1521 132 423914 5087375 The BioBox has slipped and fallen into the vent hole!
  9 56 1521 113 423904 5087389 Moving the BioBox out of the hole with the MTR.
  9 57 1521 135 423904 5087389 Dropping the other BioBox in the 7 function arm.
  10 2 1521 126 423908 5087368 Opening one of the bioboxes to get it ready for sampling- this BioBox has the rubber stopper popped out and floating about the box.
  10 3 1521 126 423908 5087368 Portable BioBox before it was filled with worms and such. R491-47
10 4 1521 126 423910 5087371 Cancelling opening the box because we seem to be kicking things up to much. Now attempting to get a port side gas tight bottle from Nascent Vent.
  10 4 1521 127 423910 5087371 3910 7381 good fix at Nascent. Positioning for this gas tight bottle. gtb-7_dfl_R491-1
10 8 1521 120 424284 5086718 Triggering the port side gas tight right near the MTR at Nascent Vent.
  10 10 1521 122 423909 5087369 Preparing to suction into Jar J3 over the tube worm bush. Flushing the ja r- lots of bag creatures in the flush jar.
  10 12 1521 124 423886 5087346 Beginning to suction right over the tube worms in the same vent as the MTR. SS-j3_bio_

R491-13

10 13 1521 121 423886 5087346 We are suctioning slowly in order to only get little stuff - not the actual worms.
  10 15 1521 125 423900 5087349 Photo
  10 18 1521 128 424287 5086715 Photo
  10 19 1521 123 423911 5087372 3912 7372
  10 26 1521 123 423914 5087379 3914 7380 good fix where we're suctioning.
  10 27 1521 124 423910 5087370 Suction over tube worm bush at Nascent vent. R491-48
10 31 1521 120 423908 5087370 BETA tape was 60 minutes, but it went to 84 minutes, so 24 minutes of the BETA tape are gone. Oops.
  10 35 1521 126 423908 5087376 Another view of the tube worm bush at Nascent vent. R491-49
10 46 1521 121 423965 5087790 Stopping suction sample after 3 1min 25 sec.
  10 48 1521 128 423907 5087370 37907 7370 good fix
  10 48 1521 128 423911 5087381 Begin sampling sJ4_bio_491-14 SS-j4_bio_

R491-14

10 53 1521 126 423911 5087381 Suction of gastropods near MTR at Nascent vent. R491-51
10 54 1521 126 423910 5087378 Still suctioning worms and gastropods at the MTR site at Nascent vent.
  10 58 1521 125 423910 5087379 Testing the Betacam to make sure error code was ok. Betacam still works.
  11 0 1521 125 423677 5087278 Stopping sample 491-14
  11 4 1521 121 423913 5087384 Grabbing external BioBox-box and moving it away from vent somewhat.
  11 6 1521 121 424287 5086716 Attempting to open the portable BioBox-box. Pulling on bungi cord. Pulling top open.
  11 9 1521 123 423956 5087788 Lid is open.
  11 10 1520 122 423770 5087694 Removing MTR4180 from vent and transferring to BioBox-box. MTR-4180_R491
11 11 1520 122 423901 5087384 Transfer complete.
  11 12 1521 126 423912 5087382 Attempting to grab tube worms.
  11 18 1520 127 423910 5087379 Portable BioBox box tipped over and was repositioned. MTR is still inside.
  11 20 1520 124 423910 5087379 Attempting to reopen the BioBox-box lid which closed on repositioning.
  11 22 1520 123 423912 5087381 Grabbing the BioBox-box with the 5 function arm. Attempting to open the lid once more.
  11 24 1520 122 423897 5087362 Still trying to open portable BioBox-box. Holding top open with 7 function and grabbing handle with 5 function. Box is open but pointed the wrong direction.
  11 26 1520 124 424912 5087076 Attempting to grab tube worm sample w/ 7 function.
  11 29 1520 116 423897 5087362 Worms grabbed. Bio_R491-16
11 29 1520 121 423897 5087362 Worms being transferred to portable BioBox-box. Held by 5 function arm. Transfer complete.
  11 32 1521 119 423897 5087362 Worm Grab-Nascent R491-52
11 32 1520 115 423897 5087362 Attempting to collect second clump of worms at same site.
  11 32 1520 121 424322 5088029 Nav is currently bad.
  11 36 1520 119 423897 5087362 Still attempting to collect second clump of worms.
  11 37 1520 119 423897 5087362 Restarting Kims VCR tape which was stopped for an unknown amount of time.
  11 43 1520 118 423893 5087358   R491-53
11 43 1520 115 423907 5087370 Second clump has been picked up by the arm and transferred to the portable BioBox-box.
  11 44 1520 121 424642 5086950 VCR tapes have been changed.
  11 52 1520 208 424467 5087140 Collecting a third clump at a nearby site after repositioning. This is a larger clump.
  11 54 1520 206 423911 5087381 Highlights on.
  11 54 1520 208 423911 5087381 Worm Grab-Nascent R491-54
11 54 1520 209 423911 5087381 still photo of worms. Photo
11 55 1520 205 423916 5087389 Collecting 4th grab of tube worms.
  11 56 1520 205 423911 5087372 Transferring a large clump to portable BioBox-box. Transfer is complete.
  11 59 1520 204 423911 5087372 Transfer complete.
  12 0 1520 209 424032 5087183 Attempting to grab handle with 7 function.
  12 4 1520 205 423909 5087365 Portable BioBox-box transferred between 5 and 7 function arms. Now attempting to close the lid with the 5 function arm.
  12 7 1520 208 424029 5087194 Closing the lid on the BioBox-box.
  12 7 1520 205 424940 5086954 Tube worm samples and MTR in the portable BioBox-box. R491-55
12 8 1520 208 423910 5087392 Transferring the BioBox-box from the 7 function back to the 5 function arm. Attempting to reattach the bungi w/ the 7 function arm. Bio-box is tipped 90 degrees.
  12 12 1519 207 423910 5087370 Grabbing the handle with the 7 function arm and returning the box to an upright position. The 5 function has let go of the handle. Setting the box down on the bottom.
  12 15 1519 210 424018 5087016 Positioning BioBox-box on the bottom.
  12 17 1520 220 423657 5086835 Nav is still bad.
  12 17 1520 229 425112 5087232 Holding the bio-box lid closed with the 5 function arm while using the 7 function to attempt to close the bungi latch. Latching complete but needs a small nudge to position the bungi completely on the latch. Bungi is now completely secure.
  12 23 1519 231 423910 5087385 Picking up the BioBox-box with the 5 function arm. Repositioning BioBox-box.
  12 25 1520 231 423910 5087383 Checking gauges.
  12 25 1520 232 423912 5087372 Opening the port side BioBox-box. Picking up MTR w/o black tape (# 3175m). Positioning for deployment.
  12 31 1519 90 423821 5087742 Deploying MTR.
  12 32 1519 59 423892 5087353 Positioning to pick up 2nd portable BioBox-box with 7 function arm. Box has been grabbed. Lifting and rotating box.
  12 35 1519 58 423898 5087375 Heading north to Mkr-N41.
  12 35 1519 49 423902 5087384 Leaving Nascent transiting to Mkr-N41.
  12 37 1518 6 423869 5087310 Nav is bad.
  12 41 1520 355 423905 5087427 Nav is showing us NW about 15 meters.
  12 42 1518 70 423911 5087422 good nav fix.
  12 42 1518 198 423910 5087420 Arriving at Mkr-N41.
  12 46 1520 103 424335 5086890 Bio-box on bottom.
  12 46 1520 101 423912 5087405 Pick up BioBox-box in 7 function arm and reposition nearer the marker.
  12 47 1520 107 423905 5087397 Bio-box slide down a pillow lava into a small depression. Attempting to open and reposition. No good nav. Positioning box. Lid is open.
  12 50 1520 104 423904 5087406 good fix.
  12 50 1520 103 423906 5087407 Grabbing MTR 4126. Into box. MTR-4126_dfl _R491-17
12 52 1520 101 423903 5087402 Beginning to collect worm sample at Mkr-N41.
  12 53 1520 92 423903 5087403 Large clump of worms collected and put into box. Bio_R491-18
12 54 1520 95 423904 5087408 Attempting to grab second clump of worms. Worms grabbed and put into the portable BioBox-box.
  12 55 1520 90 423904 5087409 good fix.
  12 55 1520 91 423904 5087409 2nd worm clump is in the box.
  12 56 1520 89 423906 5087411 good fix. Attempting to collect 3rd clump of worms at the same area.
  13 1 1520 90 423820 5087757 Third clump collected and transferred to the portable BioBox-box.
  13 2 1520 90 423904 5087409 good fix.
  13 3 1520 83 423899 5087409 Finished collecting sample. Will deploy the MTR-3041 in the same hole as the other one.
  13 4 1519 77 424085 5086794 MTR deployed and now attempting to close the lid on the portable BioBox-box.
  13 10 1520 173 424289 5086722 Still attempting to close the BioBox-box.
  13 15 1519 217 423907 5087407 good Fix. Still trying to close the lid.
  13 18 1520 23 423760 5087605 Picking up the box and repossessing with the 7 function arm.
  13 20 1520 23 423742 5087073 Repositioning box, still trying to close lid.
  13 22 1520 22 423906 5087405 Positioning 1st BioBox-box on top of the second to secure the lid. Using the 7 function to close the lid.
  13 25 1519 23 423904 5087402 Lid closed on second BioBox-box.
  13 25 1519 22 423904 5087402 Checking MTR position.
  13 25 1519 16 423991 5086488 MTR at N-41 R491-56
13 28 1519 15 423907 5087407 Still photo. good fix. Photo
13 31 1518 347 423906 5087405 Leaving Mkr-N41 transiting to oxide site 15m to the N.
  13 34 1519 10 423904 5087419 good fix.
  13 34 1519 14 423904 5087419 Positioning suction sampler to take oxide sample.
  13 37 1519 353 423899 5087420 Setting down BioBox-box in 5 function arm in order to grab the suction sampler handle. Grab complete.
  13 39 1519 1 423898 5087416 Flushing into J4 getting ready to sample FeO into J2.
  13 40 1519 359 423898 5087416 Steve's tape is running.
  13 42 1519 321 423899 5087414 good fix.
  13 43 1519 317 423900 5087413 Begin sampling FeO SS-j2_FeO
13 43 1519 321 423176 5087152 Still sampling FeO, s .
  13 44 1519 328 423898 5087412 Photo of FeO Photo
13 47 1519 285 423903 5087420 Repositioning to find more oxides.
  13 48 1519 294 423999 5087715 VCR tapes have been changed.
  13 49 1519 312 423903 5087418 good fix.
  13 49 1519 314 423904 5087415 Start sampling more oxides into J2. Mat appears to be very deep at the sampling spot.
  13 51 1519 320 424365 5088172 Possible precursor to red cherts. Need Silicon as well.
  13 52 1519 290 423904 5087418 Photo of white mat. Photo
13 53 1519 289 424878 5087155 Sampled some of the white mat as well.
  13 58 1519 288 424115 5087632 Still attempting to try to clear the clogged suction sampler.
  13 59 1519 292 424519 5087104 Switch to jar 4 and continuing to pump. SS-j4_FeO _R491-19b
14 0 1519 298 423901 5087397 Finished sampling FeO sample.
  14 1 1519 295 423531 5087297 Reposition to pick up the BioBox-box.
  14 2 1518 305 423531 5087297 photo Photo
14 2 1518 309 423531 5087297   Photo
14 3 1519 296 423902 5087419 Releasing s hose with the 5 function arm. Picking up the BioBox-box.
  14 4 1519 294 423884 5087775 Picking up BioBox-box.
  14 4 1519 291 423899 5087415 Returning to cage. Will then reposition ship.
  14 5 1519 302 423899 5087415 photo of FeO area. Photo
14 6 1518 179 423901 5087420 photo Photo
14 6 1517 188 423901 5087420   Photo
14 6 1517 201 423901 5087420   Photo
14 6 1516 223 423901 5087420   Photo
14 8 1492 24 423901 5087420 Heading back to the bottom.
  14 11 1519 359 423889 5087344 ROPOS is ready at the cage. Ship will move N about 81 m.
  14 12 1518 357 423889 5087344 Ship is starting to move.
  14 13 1518 356 423897 5087354 Transit from N41 North to old flow site. Approximately 81 m.
  14 15 1518 356 424614 5086896 Photos of mysterious black circles on lobate lavas - they look like grazing 2 photos
14 17 1518 357 423884 5087384 More photos of the 'pox' marks. Photo
14 19 1518 1 423891 5087397 'Pox' marks photo
14 22 1519 5 423890 5087419 photo of oxide Photo
14 24 1518 42 423894 5087444 good fix
  14 25 1517 27 423869 5087382 Contact with old lava.
  14 26 1518 41 423868 5087380 Crabs and clams.
  14 27 1518 30 423868 5087380 crabs and worms. Photo
14 30 1519 35 423835 5087342 No good fixes photo of BioBox-box Photo
14 32 1520 35 423834 5087362 old worms R491-57
14 34 1520 32 423847 5087359 photo of old worm site to be sampled Photo
14 36 1519 34 423868 5087463 sample of old tube worms. Bio_R491-20
14 38 1520 36 423840 5087337 Still collecting old tube worms sample. Sample placed in port BioBox-box on sub.
  14 42 1519 38 423907 5087448 Putting second clump of old worms into port BioBox-box on sub.
  14 44 1519 35 423855 5087404 Still sampling for old tube worms.
  14 53 1519 99 423835 5087390 Finishing stuffing the 'Old Flow' grab into the port BioBox - doing a little BioBox management.
  15 0 1519 101 423778 5087225 Worms in the BioBox! On to take a rock
  15 5 1519 33 423839 5087406 Going in to sample a rock for the rock guys.
  15 13 1519 83 423949 5087536 Putting the honking rock into the port BioBox. R491-58
15 16 1519 83 424134 5087869 good view of the spider crab.
  15 19 1519 85 423877 5087474 Hollow pillow(or lobate )where rock sample taken near old flow vent in association with the tube worm grab R491-59
15 20 1519 87 424110 5087884 Close-up view of hollow pillow (or lobate) near tube worm grab R491-60
15 21 1519 87 423700 5087141 Another shot of the LOBATE! flow that was sampled above. Cool crab and Nascent sample (in box) in view. R491-61
15 23 1519 86 423814 5087346 Crust of the lobate flows is about 10 cm.
  15 25 1517 104 423834 5087378 Completed sampling at 'Old Flow'. Now preparing to transit back to the elevator to drop of bioboxes.
  15 45 1518 222 423954 5087486 Looking for a basalt sample of new flow southwest of Nascent vent about 10 m.
  15 48 1516 142 423955 5087254 Lobate flows and some collapsed features. Bacterial mats in cracks between lobes. Beautiful features.
  15 52 1515 185 423906 5087295 Lava buttress.
  15 54 1508 170 423881 5087282 Still waiting for the ship to stop in order to take a rock sample.
  15 56 1518 184 423930 5087301 Lava pillar and collapsed lobates.
  15 57 1518 188 423886 5087184 Jumbled sheets on top of collapsed structure. Field of Pillars staring Mike Perfit and Kevin Costner.
  16 1 1517 356 424451 5086748 Lava pillar and arch.
  16 1 1518 360 424451 5086748 Lava pillar connected to arch Unknown fix/near Mkr-33 R491-64
16 7 1518 322 424152 5087643 collection of new lava flow near Mkr-33(10m) R491-65
16 10 1518 317 424647 5087044 Rock sample near Mkr-33 (10m) R491-67
16 7 1518 330 424027 5087484 Going in for a rock sample of new flow on a roof structure. Putting rock in starboard side of BioBox Rck_R491-22
16 10 1518 317 423903 5087223 x=42391, y=508728 - lat/long estimation for the rock sample (R491-22).
  16 12 1518 316 423898 5087246 Better lat/long for the rock sample (R491-22): 3902, 7224.
  16 14 1518 320 423936 5087307 Ship heading back to elevator, rock sampling finished.
  16 32 1492 266 423864 5087170 Ship moved. Moving back to the elevator to drop of the two portable bioboxes.
  16 46 1523 62 423866 5087100 At the elevator. Beginning to position the two portable bioboxes into the elevator.
  16 50 1519 43 423858 5087087 Tubeworm grab from N41 in portable BioBox is being transferred to the elevator.
  16 53 1519 131 423828 5087145 N41 tube worm sample going into elevator. R491-68
16 59 1520 286 423846 5087055 Nascent BioBox deposited in elevator. R491-69
17 0 1519 24 423844 5087054 TW sample of Nascent (R491-16) in left side of elevator.
  17 2 1519 91 423858 5087092 Closing the lid on the BioBox compartment of the elevator.
  17 6 1521 332 424401 5087492 Going to pick up another portable BioBox (on the seaflor at the base of the elevator) to take north to Milky area for sample of 'Old Worms'. First we are heading to Cloud for video/still camera mosaicing work. Still at the elevator doing rope managment.
  17 10 1516 246 423863 5087092 Save for ROPOS guys. Tangled launch line with the anchor. Top of the bridle. R491-70
17 10 1523 148 423551 5087340 The elevator launch line to tangled around the ropes and under the anchor. Working on untangling it.
  17 19 1521 114 424414 5086752 Removed the launch line from under the anchor. Heading to Cloud vent (N4).
  17 30 1524 317 423845 5087046 Just finished repositioning tubeworms in port side of BioBox. On our way to Cloud.
  17 43 1522 92 423885 5087118 Waiting for the ship to reposition so we can transit to Cloud.
  17 45 1519 91 423889 5087115 Collapsed area just before we reach Cloud - a ridge with a bit of venting, some gastropods and a dense aggregation of the unknown polynoid. Photo
17 47 1522 93 423886 5087066 We are at Mk-N6 (Cloud). Stirred up a lot of floc while landing ROPOS. Waiting for the floc to clear before we start taking pictures and survey the area. Kim wants video and still photo images to mosaic Cloud vent area. good fix for N6 at 3900, 7115.
  17 51 1523 268 423901 5087113 Trying to go to Mkr-N4 to begin photo imagery work - to the west.
  17 52 1523 306 423896 5087114 At Mkr-N4, good fix.
  17 55 1522 71 423897 5087118 Stopped to try and get a bunch of good fixes to locate precisely where we are.
  18 1 1522 68 423879 5087069 Starting at N4 we are going to survey all of Cloud with video, 35 mm still camera (mounted vertically) and imagenex from a depth of 1521m. We are going to take overlapping photos and since there is a 5 sec. lag between each photo, the pace will be slow.
  18 6 1522 78 423893 5087117 Lift off! Camera survey starting. Heading east. 3 photos
18 8 1523 93 423898 5087117 Aborting survey because we stirred up to much debris in the water column. Starting again.
  18 11 1522 86 423906 5087135 Back at Mkr N4, positioning to start survey again.
  18 20 1522 105 423775 5086584 Attempt #2 - starting survey from Mkr-N4 at a heading of 105 degrees for approximately 6.5 m to Mkr-N6.
  18 22 1522 104 423880 5087070 Photo survey. Imagenex head is about 2.5-3 m of the bottom. 5 photos
18 24 1522 112 423882 5087064 Still transiting to mkr N6 taking pictures. Imagenex now at 4 m. 3 photos
18 25 1521 101 423859 5087018 Imagenex up to 5 m - need to come back down.
  18 26 1521 120 423859 5087018 Imagenex still at 5 meters. Now down to 2.5 m on right and 5 m on left. Turning ROPOS around, heading back towards N4, heading 285.
  18 27 1522 277 423900 5087097 Transect from N6 back to N4. Heading staying roughly between 285-295. 12 photos
18 30 1520 294 423952 5087220 Still heading west. 2 photos
18 30 1520 280 423860 5086972 Worm Pillar. Near N4 & N6 R491-71
18 31 1518 323 423889 5087105 Moving 2 meters to the north, turning around to run a reciprocal course.
  18 32 1518 126 423890 5087110 See N4, set down briefly. Heading 105 and making a transect back towards N6. 5 photos
18 34 1520 102 423905 5087118 Got out of the venting area. We are going to fly back to N6, go to the eastern edge of the field and then will hop 2 m north, then transect west again.
  18 38 1520 275 423894 5087117 Trying to find mkr N6.
  18 39 1524 248 423904 5087109 good fix. Found N6. Hopping 2 m north.
  18 41 1523 282 423902 5087111 Heading 285 back towards N4. 9 photos
18 42 1520 294 423896 5087114 Swinging around, back to N6 (mkr 21). Hopped south and are going to do another east-west transect (N6 is just slightly to the northwest of where we are starting this transect).
  18 47 1525 283 423902 5087108 Beginning of the last line of transect. good fix
  18 48 1523 287 423948 5087533 Last transect heading west at 285. Crossing over worm pillar. Survey work 13 photos
19 3 1520 271 423742 5086943 worm moves while scientists wait R491-72
19 6 1520 272 423511 5086604 worm eats bag lunch R491-73
19 10 1510 271 424190 5087650 Moving ship to west.
  19 17 1462 188 424422 5087902 At cage. Preparing to go to Old Worm Site. Too rough to recover elevator. Will try later in the day.
  19 56 1502 357 423379 5087115 Start of Imagenex line W5. Starting at middle of line west of Mkr-33, proceeding north.
  22 47 1506 176 422412 5089376 22:16 finish north end of Imagenex line W5. 22:30 start north end of Imagenex line W6 heading south to latitude of the elevator for release
  1 19 1499 186 423306 5087087 No elevator recovery as previously planned. Moving ship towards

Mkr-N4.

  4 24 1530 93 423672 5088415 ON bottom again, very collapsed area, with large blocks of lava strewn around chaotically over rubble. A bit less floc material between flows here.
  4 27 1531 74 423662 5088434 further N, there was more continuous ropy sheets. Turning E it is more scrambled with collapse and rubble. Many of the walls seem to have fallen over or collapsed onto the floor.
  4 31 1529 59 423676 5088439 Typical view of chaotic terrain here. R491-74
4 31 1528 352 423675 5088443 good fix 3671 8442
  4 33 1530 322 423671 5088442 more to north and west hummocky terrain in sheet flows.
  4 35 1531 41 423657 5088465 featureless sheet flow in hummocky terrain, lots of floc along cracks R491-75
4 38 1527 149 423667 5088456 good fixes around this time. On boundary between sheets and collapse
  4 41 1528 122 423668 5088440 Into areas of collapse, good fixes. Very jumbled flows. spider crab.
  4 43 1528 1 423592 5088310 Folded sheets over a fairly large area. Giving up heading to old worms.
  4 51 1526 91 423711 5088440 On bottom again to E looking for old worms. just near contact. Old lavas are more rubble and covered in sed and sponges
  4 53 1527 134 423725 5088442 Older lavas R491-76
4 53 1526 132 423733 5088430 good fix, lavas are less coherent here, no floc, more sed and sponges. Some lobates as we go SE.
  4 57 1524 86 423744 5088369 rat tail in old lobates
  4 59 1525 89 423746 5088406 old lobates R491-77
5 4 1524 89 423758 5088411 Mostly lobates with some large collapse in lobates. Not much relief. some folded sheets.
  5 6 1524 96 423768 5088408 found site with some tube worms, a little floc along lavas. good fixes.
  5 7 1523 149 423785 5088410 old worm site R491-78
5 8 1526 130 423784 5088418 stopped to pick up worms and small rock. Little vent fish. Clams at base of wall in sed.
  5 10 1526 137 423788 5088412 putting box down 3787 8414 good fix for bio grab
  5 11 1526 147 423786 5088415 grabbing worms R491-79
5 13 1526 129 423784 5088419 Going in handheld BioBox. Sample of tubeworm grab. Bio_R491-23
5 16 1526 130 423786 5088417 OOPS tube worms went into stb. BioBox.
  5 22 1526 142 423786 5088417 Opening handheld BioBox now for sampling.
  5 26 1526 141 423788 5088410 second sample at old worm site. R491-80
5 26 1526 131 423785 5088421 sample was a tube worm grab at old worm site, put in handheld BioBox. Bio_R491-24
5 39 1526 90 423785 5088417 sampling the sheetflow substrate from the same locality as the tube worms. Fairly high sed cover here. very thin brittle glass hard to recover with arm. Rck_R491-25
5 44 1526 71 423784 5088419 sampling rock sheet flow R491-25 R491-81
5 45 1526 75 423784 5088417 two small pieces of the glass top of sheet flow are in with the tube worms.
  6 9 1526 119 423787 5088416 Picked up another large piece of old flow but BioBox fell over.
  6 32 1526 120 424627 5088470 still at same spot, having trouble closing lid on BioBox box. There is a bent hinge keeping it open.
  7 6 1527 101 423788 5088416 finally got the lid closed. Shrimp here too.
  7 2 1527 105 423786 5088418 Latched - finally. R491-82
7 9 1527 102 423002 5086798 Heading 31 to Whiteout. Whiteout is 160 km north of Old Worms.
  7 17 1527 116 423786 5088422 Picking up the portable BioBox with the 7-function arm. Heading to Whiteout.
  7 22 1528 309 423724 5088450 Ship is repositioning towards Whiteout.
  7 29 1528 312 423695 5088464 good fix.
  7 38 1526 1 423657 5088492 3657 8492 good fix
  7 42 1527 2 423533 5088229 Still looking for Whiteout Vent
  7 57 1528 73 423660 5088542 We're at Whiteout - which Maia says is equal to Magnesia from last year.
  8 2 1532 9 423663 5088542 Okay we are at Whiteout Vent. We can see the lava bridge.
  8 4 1532 18 423663 5088539 Flushing into bottle #4
  8 6 1532 16 423663 5088541 3662 8541 good fix at Whiteout/Magnesia Vent
  8 7 1532 15 423642 5088605 good fix 3661, 8548 at Magnesia/Whiteout sitting right in front of where we sampled
  8 10 1532 18 423051 5089138 Preparing to sample white bacterial mat/sulfur into suction bottle #2. SS-2_mat_

R491-26

8 14 1531 8 423668 5088554 We are suctioning the white stuff/orange stuff right near the vent.
  8 22 1532 58 423668 5088557 Still trying to suction at Whiteout/Magnesia
  8 25 1532 55 423668 5088558 Suction sample of bacterial mat at Whiteout/Magnesia R491-83
8 26 1532 57 423668 5088561 Still suctioning Bottle #2.
  8 30 1532 56 423666 5088555 Moving to bottle #3 to fill it with the same sort of mat/orange stuff.
  8 31 1532 58 423663 5088546 Beginning to suction Bottle #3 Sample R491-27 SS-3_mat_R491-
8 35 1532 58 423660 5088542 Still suctioning into Bottle #3.
  8 37 1532 57 423671 5088549 Done suctioning into Bottle #3. Flushing and moving on.
  8 40 1532 56 423667 5088557 We're going to suction oxide at 10 m south of here.
  8 46 1527 185 423655 5088543 Steve is waxing poetic about oxides.
  8 46 1527 201 423642 5088551 We are still heading south to look for Steve's oxides.
  8 47 1527 245 423637 5088550 These are pretty bad fixes.
  8 49 1533 157 423647 5088543 Beautiful curtain drape feature of the lavas.
  8 51 1526 28 423648 5088547 Going to do some tether management.
  8 54 1493 155 423660 5088525 Slight electrical fault - still looking for the tether.
  8 57 1490 245 423659 5088522 We can see the cage now.
  8 57 1490 197 423659 5088522 Heading back down to the bottom to go to Oxide.
  8 59 1527 174 423662 5088519 Collapsed lava lake with pillars.
  9 0 1529 176 423654 5088533 Decent ROPOS fix.
  9 1 1528 178 423660 5088548 Proceeding south to Oxide.
  9 4 1529 177 423662 5088505 Relatively good fixes.
  9 7 1534 178 423686 5088443 We se an ophiuroid amidst the oxide stuff.
  9 7 1534 195 423678 5088470 Oxide site R491-84
9 8 1534 186 423678 5088470 Sitting on some sheet flow right near the fluffy iron hydroxide orange stuff along fractures of the lava surface. Very thin glassy surface - it would break up very easily according to Steve. More little brittle stars.
  9 9 1534 184 423677 5088470 We are considering ourselves at Oxide.
  9 10 1533 169 423676 5088470 We are now looking for a place with shimmering water in order to do metal analysis on the hydroxy stuff and the water.
  9 17 1533 190 423667 5088460 iron oxide coating basalt mound
  9 17 1532 176 423667 5088460 small mound on covered with FeO R491-85
9 19 1533 153 423680 5088459 FeO sampling site R491-86
9 20 1533 158 423680 5088459 sampling iron oxide in jar 4 - contaminated sample since stuff from other flushes in the jar SS-4_FeO_

R491-28

9 22 1533 173 423679 5088458 Suction of FeO into flushing bottle R491-87
9 22 1533 170 423679 5088458 3679.1 8458 - OK fix
  9 24 1533 173 423687 5088432 Taking a look at Jar #4 to see what we got.
  9 25 1534 153 423713 5088570 Going to suction more oxides.
  9 26 1531 178 423666 5088465 x=3666 y=8465 good fix.
  9 29 1534 65 423022 5089140 Starting to suction again into Bottle #4.
  9 31 1534 64 423677 5088469 3677 8470 Okay fix
  9 32 1534 89 423671 5088461 3671 8461 good fix where we're trying to sit down and sample from.
  9 38 1534 129 423673 5088460 Still trying to suction the oxide stuff.
  9 48 1535 132 423667 5088444 We're trying to grab the suction hose.
  9 50 1535 134 423063 5089090 Still sampling. Checking out the wind velocity and gusts.
  9 57 1535 131 423055 5089095 Finishing up this suction of the oxide stuff.
  10 6 1490 184 423644 5088509 moving ship to west from Oxide (near Milky) to rejoin Imagenex line W4. We will do the northern end of W4 (going north), then move west and do line W7 (going south).
  10 28 1506 4 423429 5088546 Start of Imagenex line W4 (northern end). Ship heading north.
  11 11 1506 270 423428 5089203 End of Imagenex line W4 (northern end). Moving ship west to northern end of line W7.
  11 27 1506 179 423234 5089205 Start of Imagenex line W7 (northern end). Ship going south.
  14 45 1500 88 423247 5085989 End of line W7 (south end). Ship moving to south end of Imagenex line W6.
  15 5 1500 358 423311 5086038 Start Imagenex line W6 at 15:0 (at south end going north).
  16 22 1501 6 423305 5087098 End line W6 (south). Ship moving 60 m east to line W5 and will drive south.
  16 33 1498 189 423380 5087045 Start Imagenex line W5 (south end) ~16:27. Ship will be moving south along line.
  17 54 1498 191 423364 5086010 End of Imagenex line W5 (south end). End of dive (due to ground fault in gauge).
  18 9 1475 188 423054 5085770 ROV in the cage and coming up.
      Dive Summary: ROPOS set a new record of 41 hrs 34 min total wet time, 38.5 hrs on bottom. Dive was terminated by choice (science done, ground fault pending).
      Activities at Mkr-33, Cloud Vent (Mkr N6/21), Nascent Vent, Mkr-N41, Old Worm Site, Whiteout/Magnesia. Video/photo/Imagenex survey in area of Mkr-N4/Mkr-N6.
      Deployments: BioOsmo at Mkr-33, MTR at Cloud, MTR at Nascent, MTR at Mkr-N41. Recoveries: FeOsmo and bactraps at Mkr-33 taken to elevator, Bactraps at Cloud taken to elevator, MTR's at Nascent and Mkr-N41 taken to elevator.
      Sampling: Limpets and snails at Mkr-33, gastight at Cloud, Bacterial mat at Cloud, gastight at Nascent, suction sample over tubeworm bush, suction sample of bio at Nascent, tube worm grabs at Nascent
      Basalt sample (pillow crust) at Old Worms, basalt sample (sheet surface) near Mkr-33. Imagenex line north to area of Whiteout/Magnesia. Suction sample at Oxide area, then on with long Imagenex survey on western edge of '98 flow.



    Dive R492
      Dive Plan: To map and sample '98 lava flow, SRZ. Sample Dying Vent, if still in existence, and any new vents we might see. Also possible suction sample for Tunnicliffe of diatom/sediment on new lavas.
      ROPOS Configuration :Suction Sampler, 5 and 7 function arms with claws, cameras, and BioBox with a marker in each side and 4 rock samplers in a bag
      Note: Nav fixes not edited in the logs. All fixes were logged, regardless of accuracy. Use the dive plots to determine actual positions.
  UTC Hour Min Depth (m) Sub Hdng (deg) UTM X UTM Y Comments Frmgrbs, Photos, Samples
8 55 0 128 422710 5089068 ROPOS launched and in water. Moderate seas. Keith nervous. Snap load on down umbilical.
  8 59 32 148 422710 5089068 ROPOS out of cage.
10 37 1499 329 425223 5083490 ROPOS is at 150 m.
10 39 1496 311 422938 5085206 We are going to the bottom.
10 40 1527 171 424442 5085776 Bottom in sight.
10 41 1533 198 423629 5085815 We can se ophiouroids on pillow lavas.
10 45 1533 178 424126 5085844 took photo of bottom, pillows
10 49 1531 236 424210 5085780 Making traverses trying to find contacts between old and new lava. Initial bottom was old lava.
10 50 1531 307 424203 5085782 A number of holothorians heading NW back to line. Lobate flow. photo 3
10 52 1533 310 424186 5085799 Lots of cucumbers. A drain-out pit.
10 53 1530 255 423823 5086232 Collapsed lava pit
10 53 1529 251 423823 5086232 Lobate flow with holes drain out holes.
10 54 1529 252 424168 5085788 Going SW in more extensive drainout area. R492-1
10 55 1530 250 424168 5085788 Speed is only .5 knot. More holothorian.
10 55 1530 253 425264 5084278 Collapsed lava pit
10 55 1530 250 425264 5084278 photo of drain area. 2x photo counter advanced but no flash was seen. fish.
10 56 1530 252 424159 5085780 back to more solid lobate flows. R492-2
10 58 1530 250 424165 5085776 Ophiuroids. More solid lobate flow. Some pillows mixed with lobates. photo 6
10 59 1531 270 424142 5085771 Lobate lavas. Two flashes. Counter only advanced one
11 0 1531 273 424136 5085771 Fish and two camera flashes again.
11 1 1531 270 424135 5085769 Holothorians and fecal trails. photo 7
11 1 1531 273 424129 5085775 Holothorian. Start Highlight tape. Lobate flow with ophioids. photo 8
11 2 1530 271 424111 5085774 Double flash
11 3 1530 270 424111 5085774 photo 12
11 3 1530 271 424104 5085772 Young lava flows photo 13
11 3 1530 258 424104 5085772 contact with new flow . Clean looking lava. mats not present. less material photo 14
11 4 1530 257 424091 5085768 Pillow lavas with some brittle stars R492-5
11 3 1530 259 424098 5085771 younger lavas. Ophiods on with brittle stars. Density down from older flow. Lobate flow, flatter with small pockets of sediments. photo 15
11 4 1530 255 424078 5085762 Ophiods. R492-6
11 5 1529 257 424078 5085762 Drainout pit photo 16
11 5 1529 257 424078 5085762 Drain out area
11 5 1529 240 424070 5085759 Skate or ray swimming away R492-7
11 5 1529 244 424070 5085759 a ray going away. photo 17
11 5 1529 259 424062 5085755 young flow glassy especially on the fronts. speed .5 knot R492-8
11 6 1529 257 424044 5085744 young lobate flows. coming at us. flow direction was towards us. photo 18
11 6 1529 257 424034 5085739 Solid lobates close up
11 7 1529 256 424030 5085740 close up of lavas.
11 7 1529 260 424030 5085740 Close up of some lavas
11 8 1529 258 424013 5085734 thin lava about 1 meter thick. photo 19
11 9 1529 259 423996 5085728 so far we have come about 20 m along the 1.2 km line still in young lava, pillowy lobate, probably thin R492-9
11 10 1528 262 423980 5085723 edge of a drain out area
11 10 1528 255 423980 5085723 Edge of drained out pit
11 10 1528 255 423980 5085723 drain out
11 11 1529 260 423540 5085878 shrimp. a little yellow sed at the bottom of lobates R492-10
11 11 1528 257 423968 5085718 back to more solid flow. photo 20
11 11 1529 256 423963 5085716 small pockets of sediment but not as much as last year.
11 13 1529 258 423936 5085703 close up of lavas.
11 13 1529 256 424490 5086026 Broken up lavas
11 13 1529 258 424490 5086026 becoming more broken into drain photo 21
11 14 1528 256 424490 5086026 Skylight of drained out area R492-12
11 14 1528 256 424490 5086026 skylight couple meter thick
11 14 1529 256 424262 5085771 in a major collapse area R492-13
11 15 1530 255 423946 5085716 broken roof material photo 2
11 15 1528 257 423916 5085690 A fish
11 15 1528 258 423916 5085690 in drain out, broken material w/ fish
11 15 1529 256 423933 5085712 broken pohoi hoi R492-14
11 15 1530 257 424306 5085995 broken curtain folded photo 23
11 16 1530 258 423891 5085682 drain out area
11 16 1529 256 423886 5085686 drop of into a pit. shrimp. photo 24
11 17 1530 257 423870 5085678 drain out area photo 25
11 17 1530 253 423853 5085670 collapse sheets. More broken, sharp changes in lava morphology . complex flow
11 18 1530 256 423844 5085667 Jumbled ropy lava photo 26
11 18 1529 252 423829 5085789 jumbled flow with lots of collapse area
11 19 1529 257 423826 5085660 fish and complex lavas in R492-17
11 19 1530 256 423832 5085670 Rat tail fish
11 19 1530 248 423832 5085670 lava spire w/ glass photo 27
11 19 1529 265 423832 5085670 broken up area, jumbled flow. R492-18
11 20 1529 260 423826 5085667 sometimes coherent ropes, other times more broken up. going over ridges. photo 28
11 20 1530 252 423806 5085663 close up of jumbled flow
11 21 1529 256 423797 5085662 over a ridge into a deeper area.
11 21 1530 243 423740 5085939 we are in line with the deepest photo 29
11 22 1530 257 423777 5085660 complex terrain jumbled broken
11 23 1532 257 423761 5085658 mostly random chunks of angular lava, piled up
11 24 1532 259 423753 5085654 hydrothermal sediments
11 24 1530 257 423753 5085654 Scary looking rat tail
11 24 1529 260 423750 5085652 yellow sediment, over top of ridge, lineated flows photo 30
11 25 1531 258 423728 5085644 ridges are several meters high, last down drop was 4-5 m R492-19
11 25 1532 258 423727 5085642 lobate surface
11 26 1531 258 423721 5085638 jumbled again, don't see many animals in here
11 26 1530 257 423710 5085634 blocks of lava on ridge, on the other side drained out lobates with circular ?
11 27 1529 257 423696 5085629 lobates
11 27 1529 257 423696 5085629 Collapsed lava pits photo 31
11 27 1529 256 423696 5085629 back into lobate flows again photo 32
11 27 1530 257 423682 5085623 very glassy lobate R492-20
11 28 1529 261 423674 5085618 Window of old lava surrounded by new lava
11 28 1529 258 423682 5085623 window of older lava indicating thin ?
11 28 1528 262 423674 5085618 same R492-21
11 28 1527 264 423674 5085617 lobate photo 3
11 29 1524 245 423689 5085615 highlights off, lost controls, joystick is not functioning. photo 34
11 29 1522 253 423711 5085634 stopping the ship photo 35
11 45      Video off.
12 25 1496 298 422940 5084868 testing repairs that were made to the joystick
12 27 1525 260 422940 5084868 back on the bottom. Video back on.
12 28 1527 260 422407 5084350 back to the lobate flow
12 28 1528 264 423052 5084936 solid glassy lobate lavas.
12 29 1529 261 423532 5085556 good sub fix, to SW of ship
12 30 1529 259 423565 5085596 shallow drained out area, collapse material
12 31 1529 249 423023 5085025 collapse material
12 31 1529 261 422830 5084834 small spire, back on the solid lobate surface of the flow.
12 33 1524 259 423496 5085535 good fix photo 36
12 33 1529 259 423496 5085533 lobate lavas w/ some sediment , on glassy lobes. fish - lots of big ones
12 33 1529 259 423496 5085533 Pillow lavas
12 34 1530 260 423583 5085692 close up of lobate surface
12 34 1528 252 423481 5085530 somewhat more sediment, but not many organisms, shallow drain out area, back to lobate area R492-22
12 35 1529 258 423468 5085528 lobate surface photo 37
12 36 1530 260 423466 5085528 Lobate lavas
12 36 1530 258 423482 5085547 lot of small pockets of sediment, holothorian, more sediment photo 38
12 37 1529 261 423444 5085531 another holothorian R492-23
12 37 1529 261 423444 5085531 another holothorian
12 37 1530 261 423438 5085531 large pillow lavas
12 37 1530 256 423429 5085531 lobates w/ some pillow lavas
12 38 1531 261 423422 5085523 Cool lavas
12 39 1530 258 423413 5085521 appears to be an older surface underneath, back to solid flow photo 39
12 39 1531 255 423405 5085520 surface R492-24
12 41 1532 258 423408 5084952 more sediment, local pocket, lavas glassy, not a lot of biology
12 41 1533 254 423377 5085517 Lavas with sediment in cracks photo 40
12 41 1533 264 423344 5085738 surface
12 42 1532 257 423344 5085738 flatter with more sediment pockets, looking older, holothorian R492-26
12 42 1532 257 423536 5085638 sediment pocket photo 41
12 42 1532 260 423365 5085513 quick look at gauges
12 43 1532 268 423359 5085508 possible hydrothermal area, a call for stopping the ROV while looking at gauges, so as not to miss hydrothermal activity photo 42
12 44 1533 260 423355 5085508 possible bacterial mat, pockets
12 44 1533 258 423358 5085498 Little white patches of mat on lavas
12 44 1533 257 423349 5085504 A vent we've just found! photo 43
12 44 1532 257 423358 5085498 bacterial mat, swimming worm. stop the ship, warm water. R492-29
12 45 1534 244 423343 5085503 vent R492-30
12 45 1534 245 423343 5085503 Vent fish photo 43
12 45 1534 245 423343 5085503 photo of vent fish photo 45
12 45 1534 245 423343 5085503 Bacterial mat and flow R492-31
12 46 1534 245 423341 5085504 fixes good photo 46
12 46 1534 244 423342 5085503 bacterial mat on the edge of pit, warm water and worms R492-32
12 47 1534 247 423338 5085501 40m s of Mkr-113
12 48 1534 247 423340 5085504 good fix, highlights
12 48 1534 272 423275 5085430 bag creatures
12 49 1534 10 423337 5085506 highlights off for repair of joystick
12 50 1534 36 422850 5084915 surface near the venting area photo
12 51 1534 56 423341 5085504 worm dance
13 1 1534 54 423346 5085505 still repairing joystick photo 50
13 9 1534 226 423338 5085502 still fixing joystick,
13 10 1533 204 423343 5085504 video on, joystick fixed, good nav
13 10 1534 15 423344 5085505 Joystick Vent
13 10 1533 49 423344 5085505 vent has been named , 'joystick' vent
13 11 1535 53 423342 5085504 hole with lots of venting. R492-34
13 11 1535 50 423342 5085507 tube worms just below ROPOS, bag creature, scale worms, tube worms
13 13 1535 47 423338 5085502 vent is similar to those at the north end, lots of flock, temps are about 1.5 degrees too cold. photo 51
13 14 1535 48 423340 5085506 3 degree area, lots of white flock coming out
13 14 1534 45 423341 5085506 highlights on, going to do a short survey here
13 15 1535 46 423340 5085509 vent appears to be in new flow
13 15 1535 48 423341 5085505 looking down into deep pit, looks somewhat like whiteout vent, but not as large particles, lots of water
13 16 1534 49 423339 5085502 inside vent
13 17 1532 74 423341 5085508 Joystick Vent
13 16 1532 90 423337 5085499 trying heading 10 deg east of north to look for photo 53
13 17 1534 92 423341 5085508 worms and hole R492-37
13 18 1533 80 423339 5085508 Joystick Vent
13 17 1533 74 423341 5085508 roof collapsed into pit, ROPOS is facing NE, line of venting, grayish milky water photo 54
13 18 1533 112 423339 5085508 of bottom R492-38
13 19 1533 141 423336 5085512 Joystick Vent
13 19 1533 141 423336 5085512 lots of scale worms swimming in the water, flock from a small hole photo 5
13 20 1533 135 423337 5085512 Joystick Vent R492-39
13 20 1533 143 423337 5085512 continuing to follow pit to NE , we are looking to the SE photo 56
13 21 1534 120 423339 5085518 jumbled flow, R49240
13 21 1534 120 423339 5085518 Jumbled flow with bacterial mat near Joystick Vent
13 22 1531 74 424174 5085502 heading back to the main vent and maybe a little south photo 57
13 23 1530 168 423347 5085523 looks like a little fracture, NE trend, that has collapsed R492-41
13 25 1534 92 423339 5085508 picture of the hole
13 25 1534 98 423339 5085508 fixes good
13 25 1534 95 423340 5085504 edge of pit photo 58
13 28 1534 136 423327 5085489 Stirred up floc near Joystick Vent
13 28 1533 167 423327 5085489 still moving SW, lots of flock venting photo 59
13 29 1535 157 423337 5085495 crab, vent fish, material in water may be stirred up by ROPOS R492-44
13 30 1535 174 423336 5085492 highlights off
13 30 1534 172 423336 5085492 driving a little SW along the line
13 32 1531 72 423337 5085486 at the edge of the depression. looking around for anything else
13 34 1532 133 423339 5085477 venting appears to have run out, not the best fixes at this point.
13 35 1532 127 423342 5085467 moving a little south and west, then return to joystick vent
13 36 1534 196 423339 5085463 pillars in lobate flows, some with bacterial mat
13 37 1533 195 423324 5085455 Top of a pillar
13 37 1533 197 423324 5085455 photo of top of large pillar, staining on top, possible tube worms, confirmed tube worms.
13 38 1533 183 423320 5085462 Some sort of worm/mat stuff on top of a new lava pillar R492-45
13 39 1533 170 423322 5085455 view of top of pillar, probably to hard to sample on top of the pillar. decided to wait for a better sample photo 60
13 39 1533 169 423321 5085456 More gunk on top of this pillar R492-46
13 40 1532 189 423321 5085456 pillars conduits for vent fluids since it had tube worms on it
13 40 1530 185 423321 5085454 SW 5 m of joystick vent R492-47
13 42 1535 355 423308 5085463 may have stop soon due to weather, heading back to pit to attempt a sample
13 44 1534 18 422832 5085037 large spires near the pit
13 44 1530 54 422673 5084771 no good fixes now
13 47 1532 62 423327 5085506 back at the vent area
13 50 1534 40 423341 5085507 highlight on beta
13 50 1534 44 423341 5085507 Rock with tube worms at Joystick
13 50 1534 38 423340 5085507 Rock with tube worms at Joystick Vent
13 53 1534 45 423338 5085504 decided to suction worm first , into J1 64 micron R492-48
13 55 1535 40 423342 5085507 bag creatures at the joystick vent R492-49
13 56 1535 38 423339 5085505 suction sampling worms
13 58 1534 34 423271 5085640 suctioning tube worm photo 62
13 59 1535 39 423338 5085505 continuing to suction sample SS-j1_bio _R492-1
14 1 1535 29 423341 5085507 continuing to sample worms and bag creatures
14 3 1534 15 423337 5085501 ROPOS lifted off, ship is losing heading stop sampling
14 6 1533 255 423332 5085497 looking for a good place to put the marker
14 11 1527 92 423332 5085499 still looking around for the vent
14 14 1531 85 423345 5085515 Located sampling site.
14 15 1533 326 423342 5085510 sitting on the bottom, deploying the Mkr-42
14 17 1533 356 423321 5085488 marker 42deployed.
14 18 1532 360 423344 5085506 good fix at joystick vent , marker 42
14 19 1532 358 423343 5085506 Mkr-42 at Joystick Vent
14 20 1532 14 423343 5085505 going back to sample area to continue suction sampling in J1 R492-51
14 23 1534 104 423340 5085508 Sitting at the venting orifice where we got the odd-looking worms (nemerteans?) and getting ready to suction more for fauna.
14 25 1534 103 423340 5085509 good fix: 3340x 5508y. Suctioning Part Deux. R492-1.
14 26 1534 97 422935 5085145 More red worms on the rock below where we were suctioning. Trying to get gastropods, but failing, so going in for more of these worms.
14 37 1534 77 423339 5085503 scanning area, short camera survey R492-52
14 37 1535 68 422621 5084739 Joystick Vent
14 38 1535 68 423291 5085462 Joystick Vent
14 38 1534 59 423291 5085462 getting highlights of joystick vent R492-55
14 39 1535 83 423337 5085508 moving up on top to look for a rock sampling area R492-56
14 41 1534 72 423341 5085509 stopping highlights
14 43 1533 24 423344 5085510 VCR tape change
14 43 1532 20 423284 5085119 going to attempt a modified rock corer, hand rock corer, ROPOS-core. up on top, near marker
14 45 1533 25 423344 5085507 picking up yellow core sampler
14 47 1533 27 423339 5085503 Unused core with wax. Yellow
14 47 1533 27 423346 5085509 preparing to take core sample, grabbing sampler with 5 function arm, now grabbing with 7 function arm
14 52 1533 26 423341 5085504 Rock core success! Just went in for a second hit - got even more! Took sample a few meters southwest of marker 42 - heading 030. R49258
14 53 1533 31 423285 5085643 Core sample after collection. Yellow corer Joy Stick Vent Rck_core_

R492-2

14 55 1532 17 423346 5085513 x=3342, y=4405 - good fix for rock core R492-2.
14 57 1532 23 423344 5085507 Putting rock core in sample basket on front of BioBox. R492-59
14 58 1532 23 422968 5085238 The rock core R492-2 was orange.
14 58 1532 23 422968 5085238 Going in to suction sample for C. Moyer into jar #6. Note: two of the long, skinny nemertean looking worms got sucked into the flushing jar from the first suction sample (Jar J1, R492-1).
15 4 1534 51 423278 5085451 Suctioning bag creatures and mat for Moyer into Jar #6.
15 14 1534 42 423343 5085515 Still suctioning. A few drops of hydraulic fluid in view.
15 21 1534 35 423233 5085498 Finished suctioning jar #6, moving to jar #8. SS-6_mat_

R492-3

15 23 1534 40 423266 5085838 Positioning to suction the 'fuzzy rocks' into jar #8 for C. Moyer. Stopping for a moment to watch hydraulic fluid leak out of 5-function arm.
15 31 1533 330 423345 5085508 Starting to suction into Jar #8.
15 38 1532 91 423325 5085788 Finished suctioning some white fuzzy stuff of the rock; now looking around for another place to suction.
15 42 1534 125 423131 5085308 Suctioning more bags, mat and worms into bottle #8. SS-8_mat_

R492-4

15 47 1534 114 423338 5085511 Suction sample bottle #8 from Joystick vent (Mkr 42)
15 56 1534 124 423332 5085508 End of suctioning into Jar #8.
15 59 1533 149 423354 5085530 Heading back to the cage for tether management. R492-60
16 0 1535 263 423792 5085753 Heading southwest along original transect towards 'Dying Vent'. Lineated sheet flows. Fair amount of sediment in cracks.
16 2 1535 251 423291 5085408 Jumbled, folded flow now, moving into lineated, ropy sheet flows again.
16 5 1534 299 423281 5085722 A collapsed area with lobates on the floor. A cool pillar with square windows. A collapsed wall.
16 7 1533 209 423278 5085481 Wall with Lava Pillar just west of Joystick Vent.
16 7 1533 210 423278 5085481 Same collapsed area, wall with a lava pillar. Top of the pillar has bacterial mats. 63 photo

64 photo

16 11 1532 259 423275 5085479 Still in collapsed area with pillars and ridges. Floor is littered with collapsed plates and folded sheets. Photo of hole in the top of the pillar. R492-61
16 16 1532 260 423226 5085458 Short pillars with top. 65 photo

66 photo

16 16 1532 258 423226 5085458 Jumbled flow. Looks like short pillows on top of lobates. photo 68

photo 69

16 17 1533 262 423214 5085457 Lobate roof supported by a thin pillar - can see hole in top of pillar. R492-62
16 17 1533 259 423210 5085455 Lobate roof supported by thin pillar. SW(250) of Joystick photo 70
16 18 1532 258 423199 5085452 Ophuroids and fecal trails. Lobates on top of pillar. Holothurian and fecal trails. photo 71
16 19 1532 260 423193 5085451 Jumbled flow and lobates next to each other. R492-63
16 20 1533 259 423178 5085443 Well formed glassy lobates. Looks like a contact. We think we were on old lavas for a while and came onto new lavas. The new flow is probably getting thin in this area. Old lavas with lots of sponges, ophuroids, holothurian. photo 72

photo 73

16 22 1534 268 423162 5085445 Mixture of old and new lavas. Sea fan (gorgonian). Photo of lobate flow.
16 25 1533 255 423117 5085401 Lobate flow. Sea star and brittle stars. Old flow underneath and new bulbous, glassy flow on top. Photo of contact.
16 28 1534 256 423083 5085388 Lots of sponges, brittle stars, holothurians on older lobate flows. Sponge city"." Photo 75
16 29 1534 257 423020 5084913 Older flow. Observe sponges (white blobs) on the lava. S. west of Joystick. Photo
16 31 1534 258 423036 5085365 A little bit of bacterial mat in-between lobate flows. Tons of tiny sponges. Photo 77

Photo 78

16 33 1535 258 423039 5085379 Crab and clam site. R492-64
16 32 1533 257 422850 5084756 Tube worms and clams! Stopping to get a suction sample. x=3000, y=5350 approximate location.
16 47 1536 294 422625 5084986 Suction sampling into bottle #7 for clams and background fauna. R492-65
16 50 1536 290 423015 5085362 x=423016, y=5085362 - good fix for this sample location. Photo 80

Photo 81

16 57 1535 274 422970 5085316 Just suctioned another holothurian into bottle #7. SS7_bio_R492-5
17 2 1536 256 423013 5085361 Clams, tubeworms and collapsed lobate.
17 5 1535 270 422923 5085293 Clam sampling site, 'Coquille'. Temperature up by 0.2 degrees.
17 13 1535